### v2

# An accuracy-runtime trade-off comparison of scalable Gaussian process approximations for spatial data

Abstract
Gaussian processes (GPs) are flexible, probabilistic, non-parametric models widely employed in various fields such as spatial statistics and machine learning. A drawback of Gaussian processes is their computational cost having $\mathcal{O}(N^{3})$ time and $\mathcal{O}(N^{2})$ memory complexity which makes them prohibitive for large data sets. Numerous approximation techniques have been proposed to address this limitation. In this work, we systematically compare the accuracy of different Gaussian process approximations concerning likelihood evaluation, parameter estimation, and prediction taking into account the computational time required to perform these tasks. In other words, we analyze the trade-off between accuracy and runtime on multiple simulated and large-scale real-world data sets. We find that Vecchia approximations consistently emerge as the most accurate in almost all experiments.
Keywords: Spatial statistics; Big data; Computational efficiency
1 Introduction
Gaussian processes (GPs) are flexible, probabilistic, non-parametric models which are widely used in spatial statistics (Cressie and Wikle, 2011; Banerjee et al., 2014) and machine learning (Williams and Rasmussen, 2006). Estimation and prediction with Gaussian processes has $\mathcal{O}(N^{3})$ computational complexity and requires $\mathcal{O}(N^{2})$ memory. These costs render Gaussian processes computationally prohibitive, even when $N$ is only moderately large. To mitigate this problem, various approximations have been proposed; see Heaton et al. (2019) and Liu et al. (2020) for reviews. The goal of a Gaussian process approximation is to achieve computational efficiency with as little degradation of quality as possible compared to exact calculations. Most approximations have tuning parameters which can be adjusted to balance a trade-off between accuracy and computational efficiency. In particular, many approximations recover the exact GP calculations in a limit case for a certain choice of tuning parameters. What matters in practice, though, is which approximation has the highest accuracy for a given computational budget. Analyzing the accuracy of Gaussian process approximations without taking computational cost into account thus provides only limited insights. The goal of this article is to compare various popular Gaussian process approximations regarding this accuracy-runtime trade-off. To the best our knowledge, no such comparison has been done before in a systematic manner. Heaton et al. (2019) also compare scalable methods for modeling spatial data. However, the runtimes of the different methods are vastly different, and the methods are run using different numbers of CPU cores. Similarly, Huang et al. (2021), Abdulah et al. (2022), and Hong et al. (2023) compare various methods for analyzing spatial data but do not consider runtimes in their analysis.
We consider the following approximations: Vecchia approximations (Vecchia, 1988; Stein et al., 2004; Datta et al., 2016; Katzfuss and Guinness, 2021), covariance tapering (Furrer et al., 2005), modified predictive process approximations also called fully independent training conditional (FITC) approximations (Snelson and Ghahramani, 2005; Banerjee et al., 2008; Finley et al., 2009), full-scale approximations (FSA) combining predictive processes with tapering (Sang and Huang, 2012), multi-resolution approximations (MRA) (Katzfuss, 2017), the stochastic partial differential equations (SPDE)-based approximation introduced in Lindgren et al. (2011), fixed rank kriging (FRK) (Cressie and Johannesson, 2008), and the Fourier transform-based periodic embedding approach of Guinness (2019). Note that we are mainly interested in understanding how well the different methods approximate a certain exact Gaussian process in a sense specified below. This means that we are not primarily interested in comparing different methods concerning pure prediction accuracy. If this were the goal, we would also have to include other methods such as tree-boosting, neural networks, and other machine learning models. Most of the methods we consider (specifically, Vecchia, FITC, FSA, SPDE, and MRA) do indeed aim to approximate an exact Gaussian process with a given covariance function. FRK and the periodic embedding approach, however, do not aim to approximate a certain Gaussian process. We nonetheless include these two methods in our comparison since they are frequently used in spatial statistics (Heaton et al., 2019).
We compare the accuracy of the different approximations concerning log-likelihood evaluation, parameter estimation, and prediction on several simulated and real world data sets. For prediction accuracy, we distinguish between interpolation and extrapolation. For data sets with a moderate sample size, we also compare the quality of the approximations to exact calculations, i.e., when no approximation is used. For every approximation, we conduct multiple computations with varying complexity by adjusting tuning parameters such as the number of neighbors for Vecchia approximations and the number of knots or inducing points for modified predictive process approximations. This allows us to assess the trade-off between the quality and the computational time of the different approximations. We find that Vecchia approximations consistently emerge as the most accurate in almost all experiments.
1.1 Background on Gaussian processes
A stochastic process $f(\boldsymbol{s}),s\in D\subset\mathbb{R}^{d},$ is a Gaussian process if $(f(\boldsymbol{s}_{1}),\dots,f(\boldsymbol{s}_{N}))$ follows a multivariate Gaussian distribution for any finite set of input locations $\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{N}\in D$. A Gaussian process is defined by a mean function $\mu(\boldsymbol{s})=E[f(\boldsymbol{s})]$ and a positive definite covariance function $\quad k(\boldsymbol{s},\boldsymbol{s}^{\prime})=E\left[(f(\boldsymbol{s})-\mu(\boldsymbol{s}))(f(\boldsymbol{s}^{\prime})-\mu(\boldsymbol{s}^{\prime}))\right],\quad\boldsymbol{s},\boldsymbol{s}^{\prime}\in D$. In regression settings, one usually has a set of $N$ noisy observations $y_{1},\ldots,y_{N}$ at training locations $\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{N}$: $y(\boldsymbol{s}_{i})=f(\boldsymbol{s}_{i})+\varepsilon_{i},\quad\varepsilon_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\sigma^{2}_{n}),\quad i=1,\ldots,N.$ Assuming a zero prior mean and a parametric covariance function $k_{\theta}(\cdot,\cdot)$, where the covariance parameters denoted by $\theta$ also include the error term variance $\sigma^{2}_{n}$, the log-marginal likelihood is given by
$\log p(\mathbf{y}\mid\theta)=-\frac{1}{2}\mathbf{y}^{T}\left(K_{\theta}(S,S)+\sigma_{n}^{2}I\right)^{-1}\mathbf{y}-\frac{1}{2}\log\left|K_{\theta}(X,X)+\sigma_{n}^{2}I\right|-\frac{N}{2}\log(2\pi),$
where $\mathbf{y}=(y(\boldsymbol{s}_{1}),\dots,y(\boldsymbol{s}_{N}))^{T}$, $K_{\theta}(S,S)\in\mathbb{R}^{N\times N}$ indicates the covariance matrix evaluated at training locations $S=(S_{1},\dots,S_{N})$ with entries $K_{\theta}(S,S)_{ij}=k_{\theta}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})$. Estimation of the covariance parameters $\theta$ is usually done by maximizing this log-marginal likelihood. Further, probabilistic predictions for $f$ at a set of locations $S^{*}=(s_{1},\dots,s_{N_{p}})$ are obtained using the posterior predictive distribution given by:
$\begin{split}\mathbf{f}_{*}\mid\mathbf{y}&\sim\mathcal{N}\left(K_{\theta}(S_{*},S)\left[K_{\theta}(S,S)+\sigma_{n}^{2}I\right]^{-1}\mathbf{y},K_{\theta}(S_{*},S_{*})-K_{\theta}(S_{*},S)\left[K_{\theta}(S,S)+\sigma_{n}^{2}I\right]^{-1}K_{\theta}(S,S_{*})\right).\end{split}$
The rest of the article is structured as follows. In Section 2, we document the experimental setup including the data sets, the metrics used to evaluate the quality of the approximations, and the hardware and software resources employed. In Section 3 , we present the results from our simulated and real-world data experiments.
2 Experimental setting
In the following, we present the data sets, the evaluation criteria, the choice of tuning parameters for the approximations, and our computational environment.
2.1 Simulated data
[…]
2.2 Real-world data
2.2.1 House price data
The house data set, available from the R package spData, contains information about 25’357 single family houses sold in Lucas County, Ohio between 1993 and 1998. The response variable is the logarithmic selling price. We use the spatial coordinates as input locations for the Gaussian processes, and we assume that the mean is zero after centering the data using the sample mean of the training set. We do a random $70$/$30$ training-test split allocating $75117^{\prime}$ 751 observations to the training set, while the remaining data points are evenly divided between “interpolation” and “extrapolation” test sets each comprising $8033^{\prime}$ 803 data points. An illustration of the split as well as the distribution of house prices can be found in Figure 12 in Appendix A.1.
2.2.2 Laegern satellite data
The Laegern data set (Schneider et al., 2023) encompasses measurements of plant functional traits of the Laegern temperate mixed forest in Switzerland. We focus on the canopy height which is provided in scaled form between 00 and $1$. Values are recorded on a $1099\times 384$ grid with $6$ m spatial resolution. After removing missing values and excluding non-forest areas, there are $180279^{\prime}$ 180 observations left. Similarly to the housing data set, we allocate 70% of the observations to the training set, 15% to the “interpolation” test set, and 15% to the “extrapolation” test set. A graphical illustration can be found in Figure 13 in Appendix A.1.
2.2.3 MODIS 2016 satellite data
Next, we consider land surface temperatures (LST) recorded in degrees Celsius by the MODIS satellite (https://mrtweb.cr.usgs.gov/) on August 4, 2016. Measurements are made on a $500\times 300$ spatial grid with a resolution of $0.009273987$ degrees. The same data was used by Heaton et al. (2019), and we use the same train-test split as in Heaton et al. (2019) with $569105^{\prime}$ 569 observations in the training set and $74042^{\prime}$ 740 in the test set.Due to the presence of a clearly visible trend, see Figure 14 in Appendix A.1, we include a linear regression term with longitude and latitude coordinates as well as an intercept term in the GP models.
2.2.4 MODIS 2023 satellite data
We additionally use another temperature data coming from the MODIS satellite recorded on August 20, 2023. In this case, the spatial grid has dimensions $1200\times 1200$ with a resolution of $926.6254$ meters. We use the same data split as in Gyger et al. (2024): out of the $000600^{\prime}$ 000 available observations, $000400^{\prime}$ 000 are used for the training set, while the remaining $000200^{\prime}$ 000 are allocated to the test set. Illustrations of the temperature data and the data split are shown in Figure 15 in Appendix A.1. For the same reasons discussed previously, we again make use of the intercept and the coordinates as predictors.
2.3 Evaluation criteria
The following section introduces the criteria used to assess the quality of the different approximations. For each evaluation metric, we measure the wall-clock time needed to obtain the corresponding quantities using the computing environment described in Section 2.5. For instance, when evaluating the accuracy of covariance parameter estimates, we record the time required to fit the models. In the simulation-based experiments, where multiple repetitions are conducted, both times and metrics are averaged over the performed iterations. We use a Matérn $3/2$ covariance function for all models where this applies, i.e., all GP approximations, except FRK and periodic embedding, as well as exact calculations. For the MODIS 2016 data set, we use an exponential covariance function since Heaton et al. (2019) also used this covariance function.
2.3.1 Log-likelihood evaluation
We first analyze the accuracy of log-likelihood evaluations of the different GP approximation. In experiments with $N=10^{\prime}000$, we perform $100$ simulation iterations and consider the absolute difference to the exact log-likelihood since exact calculations can still be run in a reasonable amount of time. For the synthetic data sets with $N=100^{\prime}000$, we inspect the negative log-likelihood value itself by carrying out $20$ repetitions. For the simulated data, the log-likelihood of the training data is evaluated both at the true data-generating parameters and by using another set of parameters obtained by doubling each parameter value. For the real-world data sets, we evaluate the log-likelihood for all methods using the set of covariance parameters estimated by the most accurate method that could be implemented within a reasonable amount of time. This corresponds to using estimates of exact Gaussian processes for the housing data, and estimates of Vecchia approximations with the highest number of neighbors for the three other data sets. Additionally, we evaluate the log-likelihood at a second set of parameters obtained again by doubling the former parameters.
2.3.2 Parameter estimation
Next, we measure the accuracy of parameter estimates on the simulated data considering the bias and the mean square error (MSE). We carry out $100$ repetitions for the experiment with $N=10^{\prime}000$ and $20$ for the case when $N=100^{\prime}000$.
2.3.3 Prediction
We also evaluate the accuracy of predictive distributions. On synthetic data sets, we use the true data-generating covariance parameters for prediction. For the real-world data set, we use each technique’s estimated parameters, and prediction times include the time required to fit the model, whereas for the simulated data set we only measure the time to calculate predictions. In the simulated setting, prediction is done for the latent process, and for the real-world data sets, we predict the response variable.
The first metric used is the RMSE between the predicted and true values given by $\sqrt{\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left(f_{i}-\mu_{p,i}\right)^{2}}$ and $\sqrt{\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left(y_{i}-\mu_{p,i}\right)^{2}},$ for the simulated and real-world data, respectively, where $N_{s}$ is the sample size of the considered test data, and $\mu_{p,i}$ indicates the predictive mean for the i-th observation. To asses the accuracy of probabilistic predictions, we consider the univariate log-score defined as average univariate Gaussian negative log-likelihood (NLL) of the latent and observable processes, $\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left[\frac{\left(f_{i}-\mu_{p,i}\right)^{2}}{2\sigma_{p,i}^{2}}+\frac{1}{2}\log\left(2\pi\sigma_{p,i}^{2}\right)\right]$ and $\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left[\frac{\left(y_{i}-\mu_{p,i}\right)^{2}}{2\left(\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}\right)}+\frac{1}{2}\log\left(2\pi\left(\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}\right)\right)\right],$ where $\sigma_{p,i}^{2}$ and $\hat{\sigma}_{n}^{2}$ indicate the predictive variance of the i-th latent process realization and the estimated noise variance. The former version for the latent process is used for the simulated data and the latter version for the real-world data.
In real-world applications, we also use the continuous ranked probability score (CRPS) (Gneiting and Raftery, 2007) which can be computed as $\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\sqrt{\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}}\left(-\frac{1}{\sqrt{\pi}}+2\phi\left(\frac{y_{i}-\mu_{p,i}}{\sqrt{\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}}}\right)+\frac{y_{i}-\mu_{p,i}}{\sqrt{\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}}}\left(2\Phi\left(\frac{y_{p,i}-\mu_{p,i}}{\sqrt{\sigma_{p,i}^{2}+\hat{\sigma}_{n}^{2}}}\right)-1\right)\right)$ Gaussian distributions, where $\phi$ and $\Phi$ denote the probability density function and the cumulative density function of a standard Gaussian distribution.
On data sets with moderate sample sizes, we additionally assess the quality of the approximations by comparing their predictions to the ones obtained using exact calculations. For this, we consider the RMSE of both approximate univariate predictive means and variances compared to the ones of exact GP calculations. Further, we consider the Kullback–Leibler (KL) divergence between the exact and approximate univariate predictive distributions. In the case of univariate Gaussian distributions, the KL divergence is given by $KL(N\left(\mu_{q},\sigma_{q}\right),N\left(\mu_{p},\sigma_{p}\right))=\log\frac{\sigma_{p}}{\sigma_{q}}+\frac{\sigma_{q}^{2}+\left(\mu_{q}-\mu_{p}\right)^{2}}{2\sigma_{p}^{2}}-\frac{1}{2}.$ The number of simulation runs to evaluate the prediction accuracy for the simulated data is $50$ for $N=10^{\prime}000$ and $10$ for $N=100^{\prime}000$.
2.4 GP approximations and choice of tuning parameters
We consider the following approximation techniques: Vecchia approximations, covariance tapering, modified predictive process aka Fully Independent Training Conditional (FITC) approximations, full-scale approximations (FSA) combining predictive processes with tapering, multi-resolution approximations (MRA), the stochastic partial differential equations (SPDE)-based approximation of Lindgren et al. (2011), fixed rank kriging (FRK), and the periodic embedding approach of Guinness (2019).
The runtime and accuracy of the different approximation methods are determined by one or several tuning parameters such as the number of neighbors for Vecchia approximations and the number of knots for modified predictive process approximations. For every approximation, we use up to five different choices of tuning parameters. The selection of tuning parameter is done such that the different methods have roughly similar run-times. The specific choices of the tuning parameters thus vary across data sets and are reported in Appendix A.2. In the simulated experiments, the choice of the tuning parameters is done based on the time for evaluating the log-likelihood. In real-world applications, we instead use the total time for estimation and prediction on the (“interpolation”) test set. We also set an upper bound on the total runtime for every iteration in the experiments, including likelihood evaluation, estimation, and prediction, and we do not consider tuning parameter choices that exceed this limit. For simulated data sets with sample sizes $N=10^{\prime}000$ and $N=100^{\prime}000$, this limit is $10$ and $30$ minutes per repetition, respectively. For the real-world data, we use the following limits: $30$ minutes for the house price data set, two hours for the MODIS 2016 data set, three hours for the Laegern data set, and four hours for the MODIS 2023 data set.
Vecchia approximations are applied to the observable process $y(\cdot)$ and not the latent process $f(\cdot)$, random orderings are used as this has been found to give accurate approximations (Guinness, 2018), nearest neighbors are determined using the Euclidean distance, and for prediction, the observed locations appear first in the ordering and we condition on the observed locations only. For FITC and FSA approximations, the knots are determined using the kmeans++ algorithm. For covariance tapering, the time for log-likelihood evaluation does not include the time required to perform a symbolic Cholesky decomposition for sparsity pattern since this needs to be done only once.
2.5 Software and hardware resources
We use the following software implementations for conducting our experiments. For Vecchia approximations, covariance tapering, FITC approximations, and full-scale approximations, we use the GPBoost library version 1.5.1 (Sigrist et al., 2021). All these methods are implemented in C++ in a comparable manner which makes their comparison as fair as possible. For the other four techniques, we use the following R packages: GPvecchia version 0.1.5 for the multi-resolution approximation, INLA version 24.6.27 and rSPDE version 2.3.3 for the SPDE approach, FRK version 2.3.0 for fixed rank kriging, and npspec version 0.1.0 for periodic embedding (https://github.com/joeguinness/npspec). Parameter estimation is done with the default optimizers in the corresponding software packages. This means that differences in the estimation times of the methods implemented in GPBoost and of the remaining methods are also influenced by the optimizers used. This is, arguably, not ideal, but it is not possible to use the same optimizers in all packages. However, the runtimes for log-likelihood evaluation for the simulated and real-world data sets and for prediction for the simulated data sets do not suffer from this drawback.
Our experiments are run on a machine with a 2.25 GHz AMD EPYC 7742 processor with eight cores and 512 GB of random-access memory (RAM). Additionally, to analyze the impact of multi-core parallelization, we repeat the analysis for the large simulated data sets and the house price data set using a machine with single-core processor and the same processor model, clock speed, and amount of RAM. The code to reproduce our experiments and the data sets used in this article can be found in the following GitHub repository: https://github.com/Filippo-Rambelli/accuracy_runtime_scalable_GP_approx.
3 Results
3.1 Simulated data with $N=10^{\prime}000$
We first discuss the results for the simulated data with a sample of size $N=10^{\prime}000$ and a practical range of $0.2$. We use figures to present the results in the main article, the results in tabular form can be found in the supplementary material. The tuning parameters used for every method are reported in Table 1 in Appendix A.2.
In Figure 1, we show the absolute mean difference between the exact and approximate log-likelihood versus the evaluation time. The left panel shows the results when the evaluation is done at the true data-generating parameters, while the right panel shows the results using “wrong” parameters obtained by doubling the true parameters. Vecchia approximations clearly achieve the highest accuracy. The difference to the exact log-likelihood decreases quickly with respect to time, approaching values close to zero even for small numbers of neighbors. The second most accurate approximation for a given computational budget is the FITC approximation followed by the MRA and FSA. Tapering is the method with the overall lowest accuracy, achieving a high accuracy only for relatively large taper ranges. The results for larger and smaller practical ranges of $0.5$ and $0.05$ are reported in Appendix B.2 and Appendix B.3, respectively. Overall, we observe similar findings. Vecchia approximations are clearly the most accurate approximations. As expected, FITC approximations are more accurate for large range parameters compared to small ones, and vice versa for covariance tapering. Note that the SPDE-based approach is not included in the log-likelihood comparison as, to the best of our knowledge, the corresponding software implementation does not allow for evaluating the log-likelihood at user-defined covariance parameters. Further, FRK and and periodic embedding do not try to approximate a GP with a Matern covariance function, and it is thus pointless to analyze the accuracy of log-likelihood evaluations.
Figure 1: Absolute mean difference between approximate and exact log-likelihood on simulated data sets with a practical range of $0.2$ of and a sample size $N=10^{\prime}000$. The true data-generating parameters are used on the left and “wrong” parameters obtained by doubling the true ones are used on the right.
We next present the results for parameter estimation. In Figure 2, we show the bias and the MSE for the range, marginal variance, and error variance parameters, respectively, versus the time required to estimate these parameters. We again find that Vecchia approximations are clearly the most accurate with the lowest bias and MSE for most computational times and very small biases and MSEs even for small estimation times. The range and the variance parameters are overestimated by most methods. For tapering and FITC, some results with extreme values are excluded from the plots for better visibility of the other results. The full results in tabular form can be found in Appendix B.1. Note that we were not able to extract estimates for range and marginal variance parameters using the corresponding software implementation for the SPDE approach.
Figure 2: Bias and MSE of the estimates for the range, marginal variance, and error term variance on simulated data sets with a practical range of $0.2$ and a sample size $N=10^{\prime}000$. The dashed lines correspond to the results from the exact calculations.
In Figure 3, we report the prediction accuracy results including RMSEs, log-scores, and KL divergences to exact calculations for predicting the latent process at training locations, the “interpolation” test data, and the “extrapolation” test locations. We observe that Vecchia approximations are clearly the most accurate compared to all other methods regarding all evaluation metrics for most computational times. Overall, the MRA is the second most accurate method in this experiment, followed by the FITC and SPDE approximations. The FITC, FSA, MRA, and SPDE methods eventually achieve similar prediction accuracies as Vecchia approximations but for much larger computational times. We were not able to attain this level of accuracy with tapering in the computational times considered. Finally, FRK results in considerably less accurate predictions. For the “extrapolation” test set, we observe that the variability is much higher for all methods due to the uncertainty involved in making predictions outside of the training domain. Note that for FRK, the metrics calculated on the “extrapolation” set are not included in the figure as they were very large, i.e., very inaccurate. All results in tabular form can be found in Appendix B.1. In Appendix B.1, we additionally report the RMSE comparing the approximate predictive means and variance to the exact values.
Figure 3: Average RMSE, log-score, and KL divergence between exact and approximate predictions on simulated data sets with a practical range of $0.2$ and a sample size $N=10^{\prime}000$. We distinguish between predicting the latent process at (i) training locations (left), (ii) interpolation test locations (center), and (iii) extrapolation locations (right). The dashed lines correspond to the results from the exact calculations.
The results for experiments conducted on simulated data sets with practical ranges of 0.5 and 0.05 are reported in Appendices B.2 and B.3, respectively. In both scenarios, Vecchia approximations consistently rank among the most accurate, particularly for short runtimes. The MRA shows similar behavior across the different ranges, though it tends to be slightly less accurate than Vecchia approximations. As expected, all techniques generally are more accurate for a larger range parameter. An exception is covariance tapering, which, also as expected, is more accurate for small range parameters and less accurate for large ones.
3.2 Simulated data with $N=100^{\prime}000$
Next, we discuss the results for the simulated data with a sample of size $N=100^{\prime}000$. The tuning parameters used for every method are reported in Table 2 in Appendix A.2. Figure 16 in Appendix A.1 displays the average approximate negative log-likelihood versus the evaluation time for a practical range of $0.2$. Recall that, for this large sample size, we do not consider the difference to the exact log-likelihood, as the latter can not be calculated in a reasonable amount of time. The results are very similar to the ones of the smaller sample size. The Vecchia approximation is again the most accurate method, with the maximal accuracy being achieved for relatively small runtimes.
Concerning parameter estimation accuracy, Figure 17 in Appendix A.1 reports the bias and MSE for the range, marginal variance, and error variance parameters, respectively, versus the time required to estimate these parameters. We again find very similar results as for the smaller sample size of $N=10^{\prime}000$. Vecchia approximations are clearly the most accurate with the lowest bias and MSE for most computational times and very small biases and MSEs even for small number of neighbors. Note that for tapering, some results with extreme values are again excluded from the plots for better visibility of the other results. The full results in tabular form can be found in Appendix B.4.
For assessing prediction accuracy, Figure 18 in Appendix A.1 shows the RMSE and log-scores for predicting the latent process at training locations, the “interpolation” test data, and the “extrapolation” test data. We find that Vecchia approximations are the most accurate compared to all other methods regarding all evaluation metrics for most computational times. The SPDE-based approach results in the second most accurate predictions overall followed by the MRA and the FITC approximation. FRK and tapering generally generate the least accurate predictions. Note that the FITC approximation performs relatively well on the extrapolation test data. For FRK, the results have been excluded again from the extrapolation plots due to extreme and very inaccurate values; they can be found in Appendix 49 in tabular form. The results for the experiments using simulated data with $N=100^{\prime}000$ and practical ranges of $0.5$ and $0.05$ are reported in Appendix B.5 and Appendix B.6, respectively. Overall, the results are qualitatively very similar to the above-reported results.
3.3 House price data set
In this subsection, we present the results for the house price data set. Note that we do not use the periodic embedding approach for this data set, as the locations are not arranged on a lattice. The tuning parameters used on this data set are reported in Table 3 in Appendix A.2.
In Figure 4, we show the prediction accuracy results on the “interpolation” and “extrapolation” test sets of the housing data set. Specifically, we plot the RMSE, the log-score and the CRPS versus the prediction time. In addition, we also report the results when doing predictions with exact calculations. To further compare the differences between exact and approximate calculations, Figure 5 shows the KL divergence between the exact and approximate predictive distributions versus the runtime on the two test sets. We observe that Vecchia approximations are again the most accurate approximations having almost equal prediction accuracy as exact calculations on both the interpolation and the extrapolation test sets even for small runtimes. This is confirmed by the very small KL divergences in Figure 5 of predictive distributions of Vecchia approximations compared to exact calculations. For the interpolation test set, full-scale approximations are the second most accurate. For long runtimes, the prediction accuracy measures are similar across all approximations except for FRK. Note that for the extrapolation test data, the FITC and full-scale approximations with a small number of inducing points lead to more accurate predictions than an exact Gaussian process models. As above, the poor FRK scores on the “extrapolation” set had to be removed from the plots, but can be found in Appendix B.7. In Figure 19 in Appendix A.1, we additionally report the negative log-likelihood values versus the wall-clock time. Except for tapering, all approximations yield a similar log-likelihood for large runtimes. Vecchia approximations followed by full-scale approximations are the most accurate for a given computational budget. Additional comparisons to exact calculations are presented in Appendix B.7, where we report the RMSE between exact and approximate predictive means and variances.
Figure 4: RMSE, log-score and CRPS on the “interpolation” and “extrapolation” test sets of the house price data set. The dashed lines correspond to the results from the exact calculations.
Figure 5: KL divergence between exact and approximate predictions on the “interpolation” test (left) and “extrapolation” test (right) sets on the house price data set.
3.4 Laegern data set
In the following, we report the results for the Laegern data set. The tuning parameters used are listed in Table 4 in Appendix A.2. Figure 6 show the RMSE, the log-score, and the CRPS computed on the interpolation and extrapolation test sets. We again find that Vecchia approximations are the most accurate ones even when few neighbors are used. No other approximation reaches the accuracy of Vecchia approximations on the interpolation test set not even for long runtimes. For the extrapolation test data set, the results of the Vecchia approximation, the full-scale approximation, SPDE and the MRA are very close for large runtimes. Overall, FRK and FITC approximations are the least accurate approximations with poor scores even for large runtimes. Note that some very bad results for FITC and FRK are not shown in Figure 6 for better visibility of the other results, but they can be found in Appendix B.8. Unfortunately, it was not possible to consider results from periodic embedding as its run time exceeds our 3-hours limit. In addition, Figure 20 in Appendix A.1 shows the negative log-likelihood versus the runtime. Vecchia approximations clearly converges the fastest and achieves the lowest negative log-likelihood.
Figure 6: RMSE, log-score, and CRPS on the Laegern data set.
3.5 MODIS 2016 data set
Next, we report the results for the MODIS 2016 data set. The choices of the tuning parameters for all approximations are reporte in Table 5 in Appendix A.2. Note that for the SPDE approach, a limited range of approximation complexity can considered due to the algorithm’s failure to converge when the largest triangulation edge is smaller than $0.36$. Figures 7 presents the prediction accuracy results versus the wall-clock-time including the RMSE, the log-score, and the CRPS. Once again, the Vecchia approximation emerges as the most accurate approximation, excelling already for short runtimes, followed by the MRA. Periodic embedding also achieves good results in terms of RMSE and CRPS, but shows a relatively higher log-score due to very small estimated variances. In Figure 8, we plot the negative log-likelihood versus the wall-clock time. Vecchia approximations quickly achieve much lower negative log-likelihoods compared to all other methods. Note that negative log-likelihood values computed with the MRA have not been included in Figure 8 as they are very large, but they can be found in Tables 79 and 80 in Appendix B.9.
Figure 7: Test RMSE, log-score, and CRPS on the MODIS 2016 data set.
Figure 8: Negative log-likelihood on the MODIS 2016 data set. The parameters estimated by the Vecchia approximation with the largest considered number of neighbors are used on the left and two times these values are used on the right.
3.6 MODIS 2023 data set
We next discuss the results for the MODIS 2023 data set. The choices of the tuning parameters for all approximations are reported in Table 6 in Appendix A.2. Due to the large data size, it was only possible to implement one level of complexity, $r=1$, for the MRA. In addition, for FRK we had to leave out the case with 3 resolutions as this exceeds our time limits as well. Figure 9 reports the test set prediction accuracy results. Similarly as for other data sets, the Vecchia approximation quickly converges do a certain value. The full-scale and FITC approximations attain better prediction accuracy measures compared Vecchia approximations. This is likely a consequence of the fact that the test data contains a lot of extrapolation areas. It was already observed in Section 3.3 that low-rank approximations can be more accurate in extrapolation than exact calculations. Figure 10 displays the negative log-likelihoods versus time. Vecchia approximations converge quickly to low negative log-likelihoods while all other methods fail to converge. Note that the results of MRA are absent from the plot as they are very large, but they can be found in Tables 82 and 83.
Figure 9: Test RMSE, log-score, and CRPS on the MODIS 2023 data set.
Figure 10: Negative log-likelihood on the MODIS 2023 data set. The parameters estimated by the Vecchia approximation with 240 neighbors are used on the left, and two times these values are used on the right.
3.7 Results without multi-core parallelization
To analyze the impact of multi-core parallelization, we repeat the analysis for the large simulated data sets with $N=100^{\prime}000$ and the house price data set using a machine with single-core processor and the same processor model, clock speed, and amount of RAM. The results for the simulated data sets with ranges $\rho=0.05$, $\rho=0.2$, and $\rho=0.5$ can be found in Appendices B.11, B.12, and B.13, respectively. In Appendix B.14, we report the results for the house price data sets. The results are overall very similar to the results when using multi-core parallelization. Vecchia approximations are the most accurate approximations for a given computational budget in almost all settings. We note that there are differences among the methods regarding the speed-up of multi-core parallelization. However, given that approximations such as Vecchia, FITC, full-scale, and tapering are implemented in a different library compared to MRA and SPDE approximations, it is difficult to say whether these speed-up differences are method-inherent or due the specific implementations.
4 Conclusion
We conduct an extensive accuracy-runtime trade-off comparison between eight popular Gaussian process approximations on simulated data with different sample sizes and multiple large-scale real-world spatial data sets. We analyze the accuracy of the different methods concerning likelihood evaluation, parameter estimation, and prediction taking into account the time required to achieve a certain accuracy. Overall, Vecchia approximations clearly emerge as the most accurate approximation type. There are several open questions for future research. First, we only considered the regression case. Future research can analyze the quality of Gaussian process approximations using non-Gaussian likelihoods. For this, a challenge is that there are no closed-form expressions for marginal likelihoods and predictive distributions which means that additional approximations have to be made. Other covariance functions besides Matérn covariance functions can also be considered. Furthermore, while we did consider eight of the most prominent Gaussian process approximations in spatial statistics, we obviously did not include all available approximations. Future research can extend this to more methods.
Acknowledgments
This research was partially supported by the Swiss Innovation Agency - Innosuisse (grants number ‘55463.1 IP-ICT’ and ‘57667.1 IP-ICT’).
Appendix
A.1 Additional figures
Figure 11: Illustration of training and test locations for a simulated dataset with a sample size of $N=10^{\prime}000$.
Figure 12: Logarithmic selling price over space (left) and illustration of training and test locations (right) for the house price dataset.
Figure 13: Canopy height over space (left) and illustration of training and test locations (right) for the Laegern dataset.
Figure 14: Land surface temperature over space (left) and illustration of training and test locations (right) for the MODIS 2016 dataset.
Figure 15: Land surface temperature over space (left) and illustration of training and test locations (right) for the MODIS 2023 dataset.
[…]
Figure 17: Bias and MSE of the estimates for the range, marginal variance, end error term variance on simulated datasets with a practical range of $0.2$ and a sample size N=100’000,
Figure 18: Average RMSE and log-score on simulated datasets with a practical range of $0.2$ and a sample size N=100’000.
Figure 19: Negative log-likelihood on the house price dataset. The parameters estimated via exact calculations are used on the left and the pointwise doubling of them on the right. The dashed lines correspond to the exact negative log-likelihood.
Figure 20: Negative log-likelihood on the Laegern data set. The parameters estimated by the Vecchia approximation with the largest considered number of neighbors are used on the left and two times these values are used on the right.
A.2 Tuning parameters
Tuning parameter
Vecchia nb. neighbours 5 10 20 40 80
Tapering avg. nb. non-zeros / row 11 30 60 130 263
FITC nb. inducing points 47 254 500 950 1500
Full-scale nb. inducing points 10 24 120 300 450
avg. nb. non-zeros / row 5 8 28 100 150
FRK nb. resolutions 1 2 exceeds time limit
MRA nb. knots per partition 1 2 4 7 9
SPDE max edge 0.07 0.05 0.03 0.025 0.02
Table 1: Tuning parameters chosen for the comparison on simulated datasets with an practical range of $0.2$ and a sample size of $N=10^{\prime}000$.
Tuning parameter
Vecchia nb. neighbours 5 10 20 40 80
Tapering avg. nb. non-zeros / row 8 17 22 40 111
FITC nb. inducing points 20 200 275 900 2000
Full-scale nb. inducing points 26 91 122 250 650
avg. nb. non-zeros / row 6 11 15 27 76
FRK nb. resolutions 1 2 exceeds time limit
MRA nb. knots per partition 1 2 3 4 5
SPDE max edge 0.04 0.02 0.005 0.004 0.003
Table 2: Tuning parameters chosen for the comparison on simulated datasets with and practical range of $0.2$ and a sample size of $N=100^{\prime}000$.
Tuning parameter
Vecchia nb. neighbours 5 10 20 40 80
Tapering avg. nb. non-zeros / row 32 100 200 512 739
FITC nb. inducing points 220 500 1000 2200 3700
Full-scale nb. inducing points 106 252 444 1050 1900
avg. nb. non-zeros / row 17 36 71 256 420
FRK nb. resolutions 1 2 3 exceeds time limit
MRA nb. knots per partition 1 6 15 22 32
SPDE max edge 3000 1100 650 298 200
Table 3: Tuning parameters chosen for the comparison on the house price dataset.
Tuning parameter
Vecchia nb. neighbours 5 10 20 40 80
Tapering avg. nb. non-zeros / row 7 10 16 23 53
FITC nb. inducing points 400 760 1100 1400 2200
Full-scale nb. inducing points 180 450 775 850 1100
avg. nb. non-zeros / row 1 7 10 16 29
FRK nb. resolutions 1 2 3 exceeds time limit
MRA nb. knots per partition 1 2 3 4 6
SPDE max edge 48 30 20 15 11
NPSPEC m exceeds time limit
Table 4: Tuning parameters chosen for the comparison on the Laegern dataset.
Tuning parameter
Vecchia nb. neighbours 5 10 20 40 80
Tapering avg. nb. non-zeros / row 8 16 34 64 114
FITC nb. inducing points 380 800 1400 2000 3000
Full-scale nb. inducing points 250 400 580 800 1100
avg. nb. non-zeros / row 2 10 17 25 48
FRK nb. resolutions 1 2 3 exceeds time limit
MRA nb. knots per partition 1 3 5 8 17
SPDE max edge 0.042 0.036 convergence failure
NPSPEC m 1.2
Table 5: Tuning parameters chosen for the comparison on the MODIS 2016 dataset.
Tuning parameter
Vecchia nb. neighbours 5 10 20 40 80
Tapering avg. nb. non-zeros / row 4 8 10 14 31
FITC nb. inducing points 338 425 690 1200 2200
Full-scale nb. inducing points 5 180 345 625 1100
avg. nb. non-zeros / row 1 1 2 8 16
FRK nb. resolutions 1 2 exceeds time limit
MRA nb. knots per partition 1 exceeds time limit
SPDE max edge 0.9 0.7 0.58 0.45 0.256
NPSPEC m 1.2
Table 6: Tuning parameters chosen for the comparison on the MODIS 2023 dataset.
Supplementary material
B.1 Simulated data with $N=10^{\prime}000$ and practical range of $0.2$
B.1.1 Log-likelihood evaluation
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 310 2505 2978 3102 488
(SE) 8.49 16.6 24.4 23.7 9.31
Time (s) 5.89e-03 0.0101 9.69e-03 5.74e-03 0.24
2 Value 77.5 1533 789 2366 300
(SE) 8.95 9.28 11.3 16.8 8.91
Time (s) 0.0116 0.0664 0.11 0.0937 0.424
3 Value 19.2 994 261 661 189
(SE) 8.78 6.68 9.73 9.16 8.93
Time (s) 0.0193 0.228 0.228 0.274 0.77
4 Value 7.68 587 59.3 105 132
(SE) 8.82 6.03 9.1 8.54 8.99
Time (s) 0.0422 0.722 0.784 0.8 1.38
5 Value 0.717 353 16.1 40.2 112
(SE) 8.88 6.33 8.86 8.71 8.9
Time (s) 0.121 1.88 1.86 1.88 1.88
Table 7: Absolute mean difference between approximate and exact log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.2$ of and a sample size N=10’000. Log-likelihoods are evaluated at the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 276 3192 2376 3275 367
(SE) 4.79 8.24 18.2 13.8 6.41
Time (s) 6.31e-03 8.79e-03 8.12e-03 5.32e-03 0.246
2 Value 31.2 2361 426 2392 226
(SE) 5.39 4.5 7.83 12 6.07
Time (s) 8.72e-03 0.0651 0.108 0.0889 0.451
3 Value 17.4 1798 116 590 141
(SE) 5.59 3.24 6.49 8.38 6.23
Time (s) 0.0147 0.229 0.235 0.298 0.877
4 Value 3.71 1278 25 115 98.5
(SE) 5.81 3.03 6.07 6.44 6.24
Time (s) 0.0386 0.724 0.786 0.793 1.58
5 Value 6.33 909 7.03 48.3 83.5
(SE) 5.89 3.3 5.99 6.18 6.15
Time (s) 0.122 1.89 1.84 1.96 2.19
Table 8: Absolute mean difference between approximate and exact log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.2$ of and a sample size N=10’000. Two times the true parameters are used.
B.1.2 Parameter estimation
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias 2.95e-04 2.13 0.106 8.95e-03 5.53e-03
(SE 1) 3.08e-04 0.0884 0.0262 1.45e-03 3.66e-04
MSE 9.57e-06 5.32 0.0797 2.89e-04 4.38e-05
(SE 2) 1.21e-06 0.402 0.0536 5.66e-05 5.27e-06
Time (s) 0.17 0.812 0.378 0.356 48.2 27.4 4.53
2 Bias 1.62e-04 3.27 0.0225 9.99e-03 3e-03
(SE 1) 2.86e-04 0.116 8.18e-04 9.96e-04 2.9e-04
MSE 8.21e-06 12 5.71e-04 1.99e-04 1.73e-05
(SE 2) 1.48e-06 0.858 4.4e-05 3.13e-05 2.52e-06
Time (s) 0.316 4.2 5.42 1.72 94.5 46.1 6.54
3 Bias -1.1e-05 5.27 8.48e-03 0.0103 2.54e-03
(SE 1) 2.48e-04 0.179 3.01e-04 4.6e-04 3.45e-04
MSE 6.13e-06 31 8.1e-05 1.28e-04 1.82e-05
(SE 2) 1.04e-06 1.89 5.19e-06 1.02e-05 3.7e-06
Time (s) 0.654 12.6 13.9 13.5 81.6 16.1
4 Bias -1.06e-04 4.95 1.12e-03 2.85e-03 1.54e-03
(SE 1) 2.5e-04 0.144 2.53e-04 2.63e-04 3.11e-04
MSE 6.24e-06 26.6 7.65e-06 1.5e-05 1.19e-05
(SE 2) 9.48e-07 1.45 1.03e-06 1.87e-06 2.61e-06
Time (s) 1.75 34.9 41.4 43.7 152 24.3
5 Bias -1.43e-04 2.09 -1.95e-04 3.46e-04 1.33e-03
(SE 1) 2.4e-04 0.178 2.55e-04 2.39e-04 3.08e-04
MSE 5.8e-06 7.56 6.53e-06 5.81e-06 1.11e-05
(SE 2) 8.72e-07 1.69 9.45e-07 9.17e-07 2.71e-06
Time (s) 6.06 66.2 101 89.2 210 41.1
Table 9: Bias and MSE for the GP range, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.2$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias 0.0136 0.0412 19.1 0.179 0.0755
(SE 1) 0.0102 7.22e-03 15.5 9.93e-03 0.0146
MSE 0.0106 6.91e-03 24467 0.0419 0.0268
(SE 2) 1.47e-03 9.11e-04 23746 4.5e-03 6.08e-03
Time (s) 0.17 0.812 0.378 0.356 48.2 27.4 4.53
2 Bias 0.0121 -0.184 0.529 0.0967 0.0638
(SE 1) 9.91e-03 5.38e-03 0.0427 0.0105 0.0123
MSE 9.97e-03 0.0368 0.462 0.0204 0.0189
(SE 2) 1.87e-03 2.02e-03 0.0806 3.56e-03 3.51e-03
Time (s) 0.316 4.2 5.42 1.72 94.5 46.1 6.54
3 Bias 0.0151 -0.307 0.276 0.193 0.0881
(SE 1) 9.86e-03 4.19e-03 0.0152 0.0169 0.0163
MSE 9.96e-03 0.096 0.0995 0.0661 0.0339
(SE 2) 2.1e-03 2.59e-03 0.0104 9.49e-03 8.07e-03
Time (s) 0.654 12.6 13.9 13.5 81.6 16.1
4 Bias 9.36e-03 -0.373 0.0805 0.133 0.0527
(SE 1) 9.68e-03 3.38e-03 0.0102 0.0118 0.0133
MSE 9.47e-03 0.14 0.0168 0.0317 0.0204
(SE 2) 1.62e-03 2.53e-03 2.55e-03 4.17e-03 6.36e-03
Time (s) 1.75 34.9 41.4 43.7 152 24.3
5 Bias 6.76e-03 -0.33 0.0269 0.0842 0.0466
(SE 1) 9.44e-03 3.44e-03 9.75e-03 0.0103 0.0136
MSE 8.95e-03 0.11 0.0102 0.0177 0.0205
(SE 2) 1.49e-03 2.32e-03 1.69e-03 2.69e-03 6.92e-03
Time (s) 6.06 66.2 101 89.2 210 41.1
Table 10: Bias and MSE for the GP marginal variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.2$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias -1.83e-05 -0.309 0.344 -0.298 -0.0428 5.7e-03 0.132
(SE 1) 9.34e-04 1.92e-03 0.0101 4.12e-03 1.45e-03 1.06e-03 1.52e-03
MSE 8.73e-05 0.0959 0.129 0.0907 2.04e-03 1.44e-04 0.0177
(SE 2) 1.24e-05 1.19e-03 7.48e-03 2.33e-03 1.23e-04 1.93e-05 3.99e-04
Time (s) 0.17 0.812 0.378 0.356 48.2 27.4 4.53
2 Bias -2.33e-04 -0.206 0.0341 -0.231 -0.0428 2.18e-03 0.0647
(SE 1) 9.01e-04 1.48e-03 1.92e-03 2.89e-03 1.45e-03 9.61e-04 1.14e-03
MSE 8.12e-05 0.0428 1.53e-03 0.0544 2.04e-03 9.62e-05 4.32e-03
(SE 2) 1.24e-05 6.1e-04 1.41e-04 1.36e-03 1.23e-04 1.29e-05 1.52e-04
Time (s) 0.316 4.2 5.42 1.72 94.5 46.1 6.54
3 Bias -5.77e-04 -0.147 -6.91e-03 -0.1 1.1e-03 0.0158
(SE 1) 8.56e-04 1.19e-03 1.18e-03 1.12e-03 8.88e-04 9.24e-04
MSE 7.35e-05 0.0216 1.86e-04 0.0101 7.93e-05 3.35e-04
(SE 2) 9.75e-06 3.5e-04 2.39e-05 2.3e-04 1.02e-05 3.11e-05
Time (s) 0.654 12.6 13.9 13.5 81.6 16.1
4 Bias -4.54e-04 -0.0996 -0.0114 -0.0429 6.92e-04 8.88e-03
(SE 1) 8.64e-04 1.03e-03 9.86e-04 8.9e-04 8.81e-04 9.1e-04
MSE 7.48e-05 0.01 2.26e-04 1.92e-03 7.74e-05 1.61e-04
(SE 2) 9.99e-06 2.06e-04 2.83e-05 7.92e-05 9.94e-06 1.91e-05
Time (s) 1.75 34.9 41.4 43.7 152 24.3
5 Bias -3.91e-04 -0.075 -5.98e-03 -0.0268 5.04e-04 3.58e-03
(SE 1) 8.59e-04 1.05e-03 8.54e-04 8.6e-04 8.8e-04 8.75e-04
MSE 7.39e-05 5.74e-03 1.09e-04 7.92e-04 7.68e-05 8.86e-05
(SE 2) 9.87e-06 1.54e-04 1.58e-05 4.97e-05 9.78e-06 1.14e-05
Time (s) 6.06 66.2 101 89.2 210 41.1
Table 11: Bias and MSE for the error term variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.2$ and a sample size of N=10’000.
B.1.3 Prediction
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.27 0.466 0.529 0.509 0.557 0.3 0.382
(SE 1) 6.89e-04 7.01e-04 1.66e-03 1.04e-03 1.54e-03 7.88e-04 1.68e-03
Log-score 0.102 0.668 0.802 0.749 0.834 0.199 2.82
(SE 2) 2.59e-03 1.11e-03 3.85e-03 2.18e-03 2.82e-03 2.71e-03 0.0294
Time (s) 0.028 0.145 0.191 1.57 6.07 0.272 1.78
2 RMSE 0.247 0.402 0.379 0.467 0.487 0.278 0.304
(SE 1) 8.67e-04 5.59e-04 1.2e-03 8.21e-04 1.78e-03 7.89e-04 1.29e-03
Log-score 0.0121 0.542 0.45 0.672 0.699 0.123 0.745
(SE 2) 3.63e-03 8.92e-04 3.17e-03 1.59e-03 3.68e-03 2.93e-03 0.0128
Time (s) 0.0536 0.563 0.698 2.35 13.8 0.492 2.25
3 RMSE 0.244 0.362 0.304 0.336 0.265 0.253
(SE 1) 8.81e-04 5.33e-04 1.12e-03 6.78e-04 7.94e-04 9.17e-04
Log-score 2.49e-04 0.451 0.221 0.359 0.0743 0.0698
(SE 2) 3.74e-03 8.89e-04 3.44e-03 1.53e-03 3.12e-03 4.72e-03
Time (s) 0.111 1.37 1.64 4.93 0.96 4.71
4 RMSE 0.242 0.326 0.259 0.261 0.258 0.248
(SE 1) 8.81e-04 5.43e-04 9.92e-04 7.42e-04 7.84e-04 9.49e-04
Log-score -0.0104 0.356 0.0616 0.0969 0.0493 0.0302
(SE 2) 3.78e-03 9.68e-04 3.56e-03 2.24e-03 3.21e-03 4.62e-03
Time (s) 0.309 3.27 5.17 14.7 1.92 6.35
5 RMSE 0.241 0.3 0.247 0.25 0.255 0.244
(SE 1) 8.73e-04 5.66e-04 9.31e-04 8e-04 7.99e-04 8.97e-04
Log-score -0.0123 0.275 0.0104 0.0371 0.0407 5.2e-03
(SE 2) 3.73e-03 1.09e-03 3.68e-03 2.73e-03 3.31e-03 4.12e-03
Time (s) 1.17 8.8 12.8 25.9 2.81 9.92
Table 12: Average RMSE and log-score on the training set, standard errors and time needed for making predictions on the training set on simulated datasets with an practical range of $0.2$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.286 0.55 0.804 0.715 0.901 0.334 0.395
(SE 1) 8.91e-04 3.05e-03 6.22e-03 4.77e-03 6.6e-03 1.03e-03 1.68e-03
Log-score 0.157 0.935 1.22 1.11 1.31 0.293 2.95
(SE 2) 2.98e-03 2.14e-03 7.42e-03 4.43e-03 7.33e-03 2.82e-03 0.0281
Time (s) 0.0271 0.162 0.114 1.53 6.09 1.67 1.81
2 RMSE 0.264 0.388 0.455 0.573 0.664 0.305 0.319
(SE 1) 1.01e-03 1.28e-03 2.23e-03 3.43e-03 4.28e-03 9.64e-04 1.4e-03
Log-score 0.0739 0.715 0.635 0.948 1.01 0.204 0.833
(SE 2) 3.75e-03 9.14e-04 3.98e-03 2.87e-03 6.5e-03 3.04e-03 0.013
Time (s) 0.0554 0.613 0.469 2.15 13.5 2.23 2.29
3 RMSE 0.261 0.333 0.343 0.34 0.288 0.269
(SE 1) 1.03e-03 8.5e-04 1.59e-03 1.17e-03 1.03e-03 1.09e-03
Log-score 0.0622 0.585 0.347 0.486 0.15 0.135
(SE 2) 3.88e-03 7.59e-04 3.59e-03 1.59e-03 3.45e-03 5.24e-03
Time (s) 0.111 1.45 1.1 4.53 3.69 4.8
4 RMSE 0.257 0.305 0.283 0.274 0.278 0.264
(SE 1) 1.02e-03 7.75e-04 1.26e-03 1.01e-03 9.85e-04 1.06e-03
Log-score 0.0501 0.462 0.152 0.177 0.12 0.0939
(SE 2) 3.89e-03 8.41e-04 3.54e-03 2.2e-03 3.49e-03 4.8e-03
Time (s) 0.318 3.33 3.6 14.7 6.31 6.47
5 RMSE 0.257 0.289 0.266 0.265 0.275 0.26
(SE 1) 1.01e-03 7.99e-04 1.13e-03 1.01e-03 1e-03 1.04e-03
Log-score 0.0483 0.364 0.0855 0.109 0.109 0.0673
(SE 2) 3.87e-03 1.01e-03 3.66e-03 2.75e-03 3.61e-03 4.34e-03
Time (s) 1.19 8.88 8.97 26.1 9.39 10.1
Table 13: Average RMSE and log-score on the test “interpolation” set, standard errors and time needed for making predictions on the test “interpolation” set on simulated datasets with an practical range of $0.2$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.937 0.971 0.971 0.973 1.54 0.952 0.957
(SE 1) 0.0114 0.0112 0.0113 0.0112 0.0798 0.0112 0.0117
Log-score 1.34 1.39 1.39 1.4 2.36 1.36 1.46
(SE 2) 0.0114 0.0108 0.011 0.0109 0.175 0.011 0.0187
Time (s) 0.0944 0.229 0.208 1.99 6.05 1.62 1.8
2 RMSE 0.935 0.964 0.954 0.971 15.4 0.946 0.941
(SE 1) 0.0114 0.0112 0.0112 0.0112 1.9 0.0115 0.0113
Log-score 1.33 1.39 1.37 1.39 6.71 1.35 1.36
(SE 2) 0.0113 0.0108 0.0109 0.0108 0.633 0.0115 0.0126
Time (s) 0.148 0.787 0.908 2.8 13.5 2.15 2.29
3 RMSE 0.934 0.958 0.945 0.958 0.944 0.935
(SE 1) 0.0113 0.0113 0.0114 0.0116 0.0116 0.0113
Log-score 1.33 1.38 1.36 1.38 1.35 1.34
(SE 2) 0.0112 0.0108 0.0112 0.0112 0.0116 0.0112
Time (s) 0.259 1.88 2.2 5.9 3.41 4.8
4 RMSE 0.933 0.95 0.937 0.945 0.942 0.935
(SE 1) 0.0113 0.0113 0.0114 0.0114 0.0115 0.0113
Log-score 1.33 1.37 1.34 1.35 1.35 1.34
(SE 2) 0.0112 0.0107 0.0113 0.0111 0.0114 0.0112
Time (s) 0.66 4.56 7.17 20.5 5.55 6.47
5 RMSE 0.933 0.943 0.936 0.941 0.941 0.934
(SE 1) 0.0113 0.0113 0.0114 0.0115 0.0116 0.0114
Log-score 1.33 1.35 1.34 1.34 1.34 1.34
(SE 2) 0.0112 0.0107 0.0114 0.0112 0.0115 0.0112
Time (s) 2.35 12.8 18 37.5 7.71 10.1
Table 14: Average RMSE and log-score on the test “extrapolation” set, standard errors and time needed for making predictions on the test “extrapolation” set on simulated datasets with an practical range of $0.2$ and a sample size of N=10’000.
B.1.4 Comparison to exact calculations
Figure 21: RMSE between exact and approximate predictive means and variances on simulated datasets with an practical range of $0.2$ and a sample size N=100’000. Every approximation makes predictions using its parameter estimates.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE means 0.121 0.399 0.47 0.448 0.502 0.178 0.293
(SE 1) 2.61e-04 6.36e-04 1.56e-03 1.02e-03 1.61e-03 4.51e-04 1.55e-03
RMSE variances 0.0153 0.232 0.238 0.248 0.239 0.0446 0.0361
(SE 2) 1.16e-05 4.31e-05 5.42e-05 5.1e-05 1.22e-03 6.37e-05 7.08e-05
KL 1148 6802 8130 7617 8458 2101 27792
(SE 3) 4.27 8.38 31.5 18.4 26.3 7.74 187
Time (s) 0.028 0.145 0.191 1.57 6.07 0.272 1.78
2 RMSE means 0.0529 0.323 0.29 0.4 0.421 0.138 0.182
(SE 1) 1.42e-04 4.66e-04 1.03e-03 7.76e-04 1.87e-03 4.14e-04 9.28e-04
RMSE variances 3.06e-03 0.192 0.117 0.225 0.176 0.0294 0.0263
(SE 2) 5.64e-06 3.76e-05 1.16e-04 5.45e-05 1.44e-03 4.47e-05 9.21e-05
KL 237 5546 4593 6833 7086 1337 7421
(SE 3) 1.24 5.71 19.3 12.9 33.6 6.19 59.1
Time (s) 0.0536 0.563 0.698 2.35 13.8 0.492 2.25
3 RMSE means 0.0369 0.271 0.182 0.233 0.109 0.0743
(SE 1) 1.41e-04 4.21e-04 7.33e-04 4.65e-04 4.8e-04 2.29e-04
RMSE variances 1.47e-03 0.161 0.0587 0.125 0.0212 0.0133
(SE 2) 3.42e-07 3.14e-05 1.08e-04 6.25e-05 3.22e-05 7.53e-05
KL 116 4643 2298 3708 860 810
(SE 3) 0.894 5.01 12.1 6.74 5.97 6.15
Time (s) 0.111 1.37 1.64 4.93 0.96 4.71
4 RMSE means 0.0142 0.22 0.0939 0.1 0.0906 0.0552
(SE 1) 4.08e-05 3.7e-04 4.5e-04 2.96e-04 4.51e-04 2.06e-04
RMSE variances 2.38e-04 0.128 0.0229 0.0461 0.0169 0.0109
(SE 2) 4.68e-07 2.5e-05 8.02e-05 5.7e-05 3e-05 4.75e-05
KL 17.3 3689 725 1089 605 419
(SE 3) 0.097 4.21 5.27 3.08 4.49 3.24
Time (s) 0.309 3.27 5.17 14.7 1.92 6.35
5 RMSE means 3.18e-03 0.178 0.0503 0.0639 0.0837 0.0362
(SE 1) 1.41e-05 3.22e-04 2.33e-04 2.09e-04 4.17e-04 1.91e-04
RMSE variances 1.18e-05 0.101 9.67e-03 0.0257 0.0153 8.36e-03
(SE 2) 1.19e-08 1.95e-05 4.86e-05 4.63e-05 2.47e-05 5.52e-05
KL 0.869 2875 222 481 519 177
(SE 3) 7.81e-03 3.5 1.69 1.75 3.93 1.99
Time (s) 1.17 8.8 12.8 25.9 2.81 9.92
Table 15: Average RMSE between means, RMSE between variances, KL divergence between exact and approximate predictions, standard errors and time needed for making predictions on the training set on simulated datasets with an practical range of $0.2$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE means 0.125 0.485 0.76 0.666 0.862 0.212 0.296
(SE 1) 2.99e-04 3.3e-03 6.23e-03 5.05e-03 6.74e-03 7.64e-04 1.59e-03
RMSE variances 0.0165 0.657 0.717 0.76 0.788 0.0694 0.0435
(SE 2) 2.25e-05 2.43e-04 2.34e-04 1.9e-04 0.013 1.87e-04 7.78e-05
KL 1083 8854 11651 10592 12627 2418 28340
(SE 3) 4.3 20.2 68.7 43 71.8 10.6 182
Time (s) 0.0271 0.162 0.114 1.53 6.09 1.67 1.81
2 RMSE means 0.0576 0.29 0.373 0.511 0.609 0.164 0.187
(SE 1) 1.99e-04 1.29e-03 2.2e-03 3.57e-03 4.53e-03 8.03e-04 9.87e-04
RMSE variances 3.66e-03 0.461 0.243 0.635 0.418 0.0447 0.0322
(SE 2) 8.59e-06 2.47e-04 3.65e-04 2.79e-04 6.88e-03 1.36e-04 1.03e-04
KL 244 6654 5828 8964 9582 1548 7677
(SE 3) 1.52 6.37 29 27.9 63.4 10.1 59.8
Time (s) 0.0554 0.613 0.469 2.15 13.5 2.23 2.29
3 RMSE means 0.0414 0.213 0.224 0.221 0.129 0.079
(SE 1) 1.73e-04 5.77e-04 1.42e-03 1.09e-03 8.07e-04 3.33e-04
RMSE variances 1.88e-03 0.341 0.111 0.253 0.0318 0.0169
(SE 2) 2.54e-06 1.94e-04 2.5e-04 2.26e-04 1.09e-04 7.68e-05
KL 128 5353 2928 4353 1007 865
(SE 3) 1.02 3.18 17.1 8.76 8.86 6.15
Time (s) 0.111 1.45 1.1 4.53 3.69 4.8
4 RMSE means 0.0156 0.164 0.117 0.093 0.107 0.059
(SE 1) 4.92e-05 3.6e-04 6.9e-04 4.53e-04 7.06e-04 2.68e-04
RMSE variances 2.94e-04 0.241 0.0444 0.0771 0.0251 0.014
(SE 2) 9.6e-07 1.35e-04 1.67e-04 1.2e-04 8.07e-05 4.51e-05
KL 18.4 4130 1012 1277 708 451
(SE 3) 0.11 2.35 6.82 3.17 6.83 3.26
Time (s) 0.318 3.33 3.6 14.7 6.31 6.47
5 RMSE means 3.59e-03 0.133 0.0663 0.0627 0.0986 0.0383
(SE 1) 1.79e-05 3.01e-04 3.61e-04 2.62e-04 6.78e-04 2.37e-04
RMSE variances 1.53e-05 0.174 0.0209 0.0424 0.0227 0.0109
(SE 2) 3.76e-08 9.35e-05 1e-04 8.98e-05 8.13e-05 7.58e-05
KL 0.971 3156 359 589 606 191
(SE 3) 9.39e-03 2 2.51 1.83 6.1 1.9
Time (s) 1.19 8.88 8.97 26.1 9.39 10.1
Table 16: Average RMSE between means, RMSE between variances, KL divergence between exact and approximate predictions, standard errors and time needed for making predictions on the test “interpolation” set on simulated datasets with an practical range of $0.2$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE means 0.0808 0.275 0.275 0.284 1.18 0.185 0.199
(SE 1) 2.17e-03 7.2e-03 6.98e-03 7.39e-03 0.0917 5.86e-03 6.43e-03
RMSE variances 0.0138 0.227 0.224 0.232 0.225 0.118 0.204
(SE 2) 9.7e-05 5.64e-04 6.25e-04 6.01e-04 3.43e-03 3.64e-03 5.94e-03
KL 67 677 678 705 10321 322 1238
(SE 3) 3.37 21 20.9 22.2 1658 14.9 40.2
Time (s) 0.0944 0.229 0.208 1.99 6.05 1.62 1.8
2 RMSE means 0.0559 0.253 0.203 0.275 15.3 0.156 0.104
(SE 1) 1.41e-03 6.75e-03 4.89e-03 7.31e-03 1.9 5.66e-03 4.84e-03
RMSE variances 6.77e-03 0.214 0.135 0.228 38.4 0.094 0.147
(SE 2) 7.08e-05 4.85e-04 9.26e-04 5.81e-04 3.72 3.01e-03 6.42e-03
KL 31.1 601 433 678 53593 240 291
(SE 3) 1.53 18.4 12.7 21.6 6390 12.9 13.4
Time (s) 0.148 0.787 0.908 2.8 13.5 2.15 2.29
3 RMSE means 0.0274 0.228 0.152 0.225 0.137 0.0474
(SE 1) 6.8e-04 6.25e-03 4.57e-03 5.57e-03 5.23e-03 2.35e-03
RMSE variances 1.95e-03 0.196 0.0879 0.175 0.0788 0.102
(SE 2) 3.22e-05 4.08e-04 9.07e-04 5.34e-04 2.94e-03 8.9e-03
KL 6.88 515 272 499 186 64.8
(SE 3) 0.337 15.6 11 14.5 10.7 6.58
Time (s) 0.259 1.88 2.2 5.9 3.41 4.8
4 RMSE means 0.012 0.19 0.0964 0.152 0.125 0.0356
(SE 1) 3.27e-04 5.5e-03 2.83e-03 4.61e-03 4.58e-03 2.14e-03
RMSE variances 4.86e-04 0.169 0.0488 0.102 0.0683 0.0941
(SE 2) 1.03e-05 3.18e-04 6.16e-04 7.48e-04 2.03e-03 8.96e-03
KL 1.25 393 119 247 160 46.1
(SE 3) 0.0732 11.7 4.87 9.08 8.68 6.12
Time (s) 0.66 4.56 7.17 20.5 5.55 6.47
5 RMSE means 4.4e-03 0.151 0.0672 0.12 0.117 0.0265
(SE 1) 1.36e-04 4.6e-03 2.18e-03 4.21e-03 4.5e-03 1.78e-03
RMSE variances 8e-05 0.14 0.0286 0.0741 0.0635 0.086
(SE 2) 1.91e-06 2.45e-04 5.82e-04 6.91e-04 1.82e-03 8.88e-03
KL 0.16 278 58.7 156 143 34.5
(SE 3) 0.0104 8.08 2.72 7.18 8.05 5.6
Time (s) 2.35 12.8 18 37.5 7.71 10.1
Table 17: Average RMSE between means, RMSE between variances, KL divergence between exact and approximate predictions, standard errors and time needed for making predictions on the test “extrapolation” set on simulated datasets with an practical range of $0.2$ and a sample size of N=10’000.
B.2 Simulated data with $N=10^{\prime}000$ and practical range of $0.5$
Tuning parameter
Vecchia nb. neighbours 5 10 20 40 80
Tapering num. non-zero entries 11 30 60 130 263
FITC num. inducing points 47 254 500 950 1500
Full-scale num. inducing points 10 24 120 300 450
num. non-zero entries 5 8 28 100 150
FRK num. resolutions 1 2 exceeds time limit
MRA num. knots per partition 1 2 4 7 9
SPDE max edge 0.07 $0.05$ 0.03 0.025 0.02
Table 18: Tuning parameters chosen for the comparison on simulated datasets with an practical range of $0.5$ and a sample size of $N=10^{\prime}000$.
B.2.1 Log-ikelihood evaluation
Figure 22: Absolute mean difference between approximate and exact log-likelihood on simulated datasets with an practical range $0.5$ of and a sample size $N=10^{\prime}000$. The true data-generating parameters are used on the left, and two times these values are used on the right.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 545 3195 1086 2788 242
(SE) 8.6 41.2 17 38.3 8.11
Time (s) 5.68e-03 0.0111 8.01e-03 5.29e-03 0.241
2 Value 208 2132 53 1523 142
(SE) 7.83 21.7 7.47 18.1 7.66
Time (s) 0.011 0.0649 0.083 0.0109 0.449
3 Value 56.5 1515 8.1 153 81.8
(SE) 7.7 13.1 7.52 7.73 7.62
Time (s) 0.019 0.224 0.268 0.244 0.908
4 Value 15.4 1009 0.969 12.5 55.3
(SE) 7.56 8.38 7.54 7.45 7.56
Time (s) 0.0422 0.816 0.774 0.983 1.66
5 Value 7.25 685 0.178 3.48 46.2
(SE) 7.5 6.91 7.53 7.51 7.6
Time (s) 0.122 1.93 1.88 1.57 2.29
Table 19: Absolute mean difference between approximate and exact log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.5$ of and a sample size N=10’000. Log-likelihoods are evaluated at the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 528 3808 648 2345 182
(SE) 4.44 20.5 11.2 30.2 5.18
Time (s) 6.17e-03 9.74e-03 6.69e-03 5.18e-03 0.253
2 Value 185 2938 23 1189 108
(SE) 4.07 10.8 4.27 15.1 4.67
Time (s) 7.91e-03 0.0634 0.0758 0.0107 0.49
3 Value 27.9 2337 4.03 99.3 62.3
(SE) 4.07 6.44 4.24 4.78 4.59
Time (s) 0.0144 0.225 0.267 0.269 0.971
4 Value 10.5 1762 0.657 8.33 42.3
(SE) 4.15 4.09 4.23 4.24 4.43
Time (s) 0.0385 0.816 0.775 1.01 1.73
5 Value 1.51 1335 0.155 2.26 35.5
(SE) 4.23 3.39 4.23 4.23 4.43
Time (s) 0.124 1.93 1.89 1.57 2.42
Table 20: Absolute mean difference between approximate and exact log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.5$ of and a sample size N=10’000. Log-likelihoods are evaluated at two-times the true data-generating parameters.
B.2.2 Parameter estimation
Figure 23: Bias and MSE of the estimates for the GP range, GP variance end error term variance on simulated datasets with an practical range of $0.5$ and a sample size N=100’000,
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias 2.31e-03 2.23 0.21 0.0524 0.0261
(SE 1) 1.95e-03 0.0865 0.0464 7.23e-03 1.52e-03
MSE 3.84e-04 5.74 0.259 7.97e-03 9.1e-04
(SE 2) 6.48e-05 0.429 0.125 2.16e-03 1.13e-04
Time (s) 0.153 0.925 0.557 0.432 47.2 21.7 4.08
2 Bias 3.44e-03 3.82 8.02e-03 0.0388 0.0235
(SE 1) 1.68e-03 0.146 1.35e-03 4.86e-03 1.32e-03
MSE 2.95e-04 16.7 2.46e-04 3.87e-03 7.24e-04
(SE 2) 5.22e-05 1.31 3.75e-05 9.02e-04 7.26e-05
Time (s) 0.357 4.16 4.55 0.742 108 36.9 7.05
3 Bias 2.23e-03 5.14 8.14e-04 0.0104 0.0217
(SE 1) 1.32e-03 0.189 1.16e-03 1.42e-03 1.26e-03
MSE 1.79e-04 30 1.36e-04 3.09e-04 6.29e-04
(SE 2) 3.26e-05 2.07 2.39e-05 5.64e-05 6.8e-05
Time (s) 0.851 11.9 16.1 14.1 70.4 20.9
4 Bias 2.16e-03 6.6 1.15e-03 -1.19e-03 0.0193
(SE 1) 1.16e-03 0.167 1.13e-03 1.12e-03 1.34e-03
MSE 1.39e-04 46.3 1.29e-04 1.27e-04 5.51e-04
(SE 2) 2.22e-05 2.24 2.48e-05 2.18e-05 6.55e-05
Time (s) 2.55 43.4 52 60.7 133 29.3
5 Bias 2.2e-03 6.99 1.67e-03 -2.07e-04 0.0207
(SE 1) 1.19e-03 0.181 1.13e-03 1.12e-03 1.2e-03
MSE 1.46e-04 52.1 1.3e-04 1.26e-04 5.69e-04
(SE 2) 3.17e-05 2.78 2.46e-05 2.3e-05 6.12e-05
Time (s) 9.26 95.7 133 101 175 47.5
Table 21: Bias and MSE for the GP range, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.5$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias 0.0347 0.0317 75.1 0.614 0.423
(SE 1) 0.027 0.0175 43.9 0.145 0.0355
MSE 0.0739 0.0315 198314 2.49 0.304
(SE 2) 0.013 4.29e-03 134023 1.62 0.0604
Time (s) 0.153 0.925 0.557 0.432 47.2 21.7 4.08
2 Bias 0.0532 -0.215 0.219 0.767 0.484
(SE 1) 0.0295 0.0129 0.0299 0.143 0.033
MSE 0.0898 0.0629 0.137 2.63 0.342
(SE 2) 0.0214 5.63e-03 0.0273 1.14 0.0431
Time (s) 0.357 4.16 4.55 0.742 108 36.9 7.05
3 Bias 0.0455 -0.37 0.0593 0.358 0.493
(SE 1) 0.0259 9.96e-03 0.0224 0.0419 0.0321
MSE 0.0691 0.147 0.0537 0.304 0.345
(SE 2) 0.0164 7.29e-03 0.0103 0.0917 0.0367
Time (s) 0.851 11.9 16.1 14.1 70.4 20.9
4 Bias 0.0442 -0.506 0.0391 0.0619 0.458
(SE 1) 0.0222 7.26e-03 0.0219 0.0226 0.0356
MSE 0.0514 0.261 0.0495 0.0551 0.336
(SE 2) 8.44e-03 7.31e-03 8.86e-03 0.011 0.0462
Time (s) 2.55 43.4 52 60.7 133 29.3
5 Bias 0.0472 -0.592 0.0404 0.0387 0.501
(SE 1) 0.0236 5.41e-03 0.022 0.0218 0.0324
MSE 0.0579 0.353 0.0498 0.049 0.355
(SE 2) 0.0107 6.38e-03 8.92e-03 9e-03 0.0432
Time (s) 9.26 95.7 133 101 175 47.5
Table 22: Bias and MSE for the GP marginal variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.5$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias -1.58e-04 -0.307 0.0533 -0.138 -9.7e-03 1.16e-03 0.0144
(SE 1) 8.64e-04 3.33e-03 2.9e-03 5.87e-03 1.4e-03 8e-04 7.82e-04
MSE 7.46e-05 0.0957 3.68e-03 0.0224 2.89e-04 6.47e-05 2.67e-04
(SE 2) 8.93e-06 2.03e-03 3.67e-04 1.63e-03 3.51e-05 8.69e-06 2.3e-05
Time (s) 0.153 0.925 0.557 0.432 47.2 21.7 4.08
2 Bias -7.83e-05 -0.205 -8.44e-03 -0.109 -9.7e-03 6.14e-04 5.23e-03
(SE 1) 8.35e-04 2.43e-03 8.89e-04 2.33e-03 1.4e-03 7.96e-04 7.9e-04
MSE 6.97e-05 0.0425 1.5e-04 0.0125 2.89e-04 6.31e-05 8.92e-05
(SE 2) 8.45e-06 9.94e-04 1.8e-05 5.33e-04 3.51e-05 8.02e-06 1.11e-05
Time (s) 0.357 4.16 4.55 0.742 108 36.9 7.05
3 Bias -2.22e-04 -0.144 -3.52e-03 -0.0296 3.49e-04 8.61e-04
(SE 1) 8.19e-04 1.73e-03 8.25e-04 9.57e-04 7.75e-04 8.07e-04
MSE 6.71e-05 0.0211 8.04e-05 9.69e-04 5.96e-05 6.52e-05
(SE 2) 8.59e-06 4.97e-04 1.05e-05 5.68e-05 7.82e-06 8.24e-06
Time (s) 0.851 11.9 16.1 14.1 70.4 20.9
4 Bias -2.27e-04 -0.0951 -1.04e-03 -6.62e-03 1.67e-04 7.24e-04
(SE 1) 7.94e-04 1.2e-03 8.01e-04 8.23e-04 7.97e-04 7.96e-04
MSE 6.3e-05 9.2e-03 6.52e-05 1.12e-04 6.29e-05 6.32e-05
(SE 2) 7.81e-06 2.26e-04 8.3e-06 1.34e-05 8.04e-06 7.97e-06
Time (s) 2.55 43.4 52 60.7 133 29.3
5 Bias -2.41e-04 -0.0648 -4.51e-04 -2.51e-03 2.38e-04 7.65e-04
(SE 1) 7.92e-04 9.4e-04 7.94e-04 8.06e-04 7.95e-04 7.98e-04
MSE 6.28e-05 4.28e-03 6.33e-05 7.13e-05 6.26e-05 6.36e-05
(SE 2) 7.97e-06 1.21e-04 7.93e-06 9.06e-06 8.14e-06 8.08e-06
Time (s) 9.26 95.7 133 101 175 47.5
Table 23: Bias and MSE for the error term variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.5$ and a sample size of N=10’000.
B.2.3 Prediction
Figure 24: Average RMSE and log-score on simulated datasets with an practical range of $0.5$ and a sample size N=100’000. Predictions are done using the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.229 0.464 0.358 0.465 0.474 0.19 0.171
(SE 1) 5.83e-04 1.47e-03 2.28e-03 2.28e-03 3.11e-03 8.41e-04 1.19e-03
Log-score -0.0574 0.663 0.433 0.707 0.671 -0.26 -0.113
(SE 2) 2.55e-03 2.27e-03 7.59e-03 7.9e-03 6.59e-03 4.79e-03 0.0156
Time (s) 0.0267 0.162 0.134 1.58 5.67 0.29 1.75
2 RMSE 0.177 0.395 0.172 0.375 0.29 0.173 0.155
(SE 1) 6.31e-04 7.16e-04 1.29e-03 1.42e-03 2.23e-03 8.62e-04 1.03e-03
Log-score -0.315 0.528 -0.343 0.484 0.184 -0.353 -0.414
(SE 2) 3.54e-03 1.11e-03 6.72e-03 4.6e-03 7.58e-03 5.37e-03 8.8e-03
Time (s) 0.0476 0.574 0.676 1.93 13.5 0.638 2.24
3 RMSE 0.154 0.35 0.151 0.194 0.163 0.148
(SE 1) 7.81e-04 5.36e-04 1.07e-03 1.09e-03 8.77e-04 9.73e-04
Log-score -0.457 0.425 -0.475 -0.187 -0.412 -0.495
(SE 2) 5.16e-03 8.73e-04 6.53e-03 4.42e-03 5.7e-03 6.99e-03
Time (s) 0.111 1.42 1.88 5.23 1.54 4.71
4 RMSE 0.151 0.306 0.148 0.152 0.158 0.147
(SE 1) 8.55e-04 4.79e-04 9.68e-04 1.03e-03 8.95e-04 9.57e-04
Log-score -0.479 0.312 -0.502 -0.463 -0.44 -0.5
(SE 2) 5.76e-03 8.38e-04 6.46e-03 6.07e-03 5.92e-03 6.78e-03
Time (s) 0.312 4.27 5.61 16.6 3.01 6.15
5 RMSE 0.148 0.272 0.147 0.148 0.156 0.147
(SE 1) 8.87e-04 4.57e-04 9.45e-04 9.84e-04 8.93e-04 9.58e-04
Log-score -0.498 0.208 -0.505 -0.494 -0.45 -0.503
(SE 2) 6.05e-03 8.55e-04 6.45e-03 6.3e-03 6.01e-03 6.78e-03
Time (s) 1.26 9.05 13.3 26.5 4.36 9.64
Table 24: Average RMSE and log-score on the training set, standard errors and time needed for making predictions on the training set on simulated datasets with an effective range of $0.5$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.232 0.536 0.415 0.59 0.641 0.198 0.175
(SE 1) 7.04e-04 7.14e-03 4.57e-03 8.5e-03 7.63e-03 1.04e-03 1.26e-03
Log-score -0.0425 0.924 0.583 0.979 0.971 -0.224 -0.082
(SE 2) 3.04e-03 4.94e-03 9.82e-03 0.0116 0.0119 5.26e-03 0.016
Time (s) 0.0277 0.18 0.053 1.53 5.65 1.72 1.78
2 RMSE 0.181 0.366 0.178 0.41 0.318 0.18 0.158
(SE 1) 7.23e-04 2.28e-03 1.49e-03 3.62e-03 2.91e-03 9.9e-04 1.15e-03
Log-score -0.296 0.69 -0.304 0.644 0.277 -0.32 -0.388
(SE 2) 3.97e-03 1.49e-03 6.74e-03 6.14e-03 9.07e-03 5.77e-03 9.61e-03
Time (s) 0.0517 0.616 0.724 1.89 13.1 2.63 2.27
3 RMSE 0.158 0.305 0.156 0.198 0.168 0.152
(SE 1) 9.03e-04 9.72e-04 1.22e-03 1.44e-03 1.04e-03 1.1e-03
Log-score -0.434 0.545 -0.445 -0.141 -0.381 -0.471
(SE 2) 5.78e-03 7.4e-04 6.89e-03 4.7e-03 6.34e-03 7.56e-03
Time (s) 0.113 1.52 1.26 4.73 5.3 4.8
4 RMSE 0.155 0.267 0.151 0.156 0.163 0.151
(SE 1) 9.72e-04 6.61e-04 1.1e-03 1.17e-03 1.05e-03 1.07e-03
Log-score -0.455 0.401 -0.477 -0.435 -0.412 -0.476
(SE 2) 6.34e-03 6.17e-04 6.94e-03 6.26e-03 6.62e-03 7.29e-03
Time (s) 0.321 4.35 3.86 16.1 9.47 6.27
5 RMSE 0.152 0.243 0.151 0.152 0.161 0.151
(SE 1) 1.01e-03 5.94e-04 1.07e-03 1.11e-03 1.04e-03 1.07e-03
Log-score -0.474 0.278 -0.481 -0.469 -0.422 -0.48
(SE 2) 6.65e-03 6.64e-04 7e-03 6.74e-03 6.65e-03 7.31e-03
Time (s) 1.28 9.05 9.2 26.4 12.7 9.82
Table 25: Average RMSE and log-score on the test “interpolation” set, standard errors and time needed for making predictions on the test “interpolation” set on simulated datasets with an practical range of $0.5$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.849 0.965 0.907 0.949 1.69 0.881 0.833
(SE 1) 0.0247 0.027 0.0256 0.0269 0.0881 0.0261 0.025
Log-score 1.2 1.4 1.33 1.39 3.96 1.26 1.19
(SE 2) 0.0258 0.028 0.0259 0.0278 0.302 0.0281 0.0294
Time (s) 0.0944 0.245 0.112 1.98 5.65 1.65 1.77
2 RMSE 0.847 0.958 0.84 0.929 4.94 0.863 0.829
(SE 1) 0.0248 0.027 0.0262 0.0256 0.408 0.026 0.0241
Log-score 1.2 1.39 1.2 1.36 3.76 1.23 1.17
(SE 2) 0.026 0.0278 0.0271 0.0259 0.237 0.0274 0.0272
Time (s) 0.141 0.778 1.47 2.38 13.1 2.4 2.27
3 RMSE 0.842 0.95 0.828 0.861 0.849 0.827
(SE 1) 0.0247 0.027 0.0252 0.0265 0.0255 0.0244
Log-score 1.18 1.38 1.17 1.25 1.2 1.16
(SE 2) 0.0259 0.0276 0.0261 0.0272 0.0261 0.0281
Time (s) 0.261 1.94 2.51 6.45 4.52 4.79
4 RMSE 0.833 0.936 0.827 0.837 0.838 0.828
(SE 1) 0.0245 0.0269 0.0245 0.0254 0.025 0.0247
Log-score 1.17 1.36 1.16 1.19 1.18 1.17
(SE 2) 0.0259 0.0273 0.0255 0.0261 0.0256 0.0287
Time (s) 0.663 5.83 7.72 23.4 7.98 6.27
5 RMSE 0.829 0.919 0.827 0.831 0.835 0.828
(SE 1) 0.0245 0.0268 0.0244 0.0252 0.025 0.0247
Log-score 1.16 1.34 1.16 1.17 1.18 1.17
(SE 2) 0.0261 0.0268 0.0258 0.026 0.0256 0.0287
Time (s) 2.54 13.2 18.4 37.6 11.1 9.83
Table 26: Average RMSE and log-score on the test “extrapolation” set, standard errors and time needed for making predictions on the test “extrapolation” set on simulated datasets with an practical range of $0.5$ and a sample size of N=10’000.
B.2.4 Comparison to exact calculations
Figure 25: RMSE between exact and approximate predictive means and variances on simulated datasets with an practical range of $0.5$ and a sample size N=100’000. Every approximation makes predictions using its parameter estimates.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE means 0.175 0.439 0.326 0.44 0.45 0.12 0.0866
(SE 1) 4.41e-04 1.45e-03 2.25e-03 2.33e-03 3.26e-03 6.33e-04 5.62e-04
RMSE variances 0.0309 0.267 0.14 0.246 0.203 0.0202 7.86e-03
(SE 2) 4.2e-06 5.07e-05 1.1e-04 7.15e-05 2.88e-03 3.22e-05 3.76e-05
KL 4514 11716 9383 12101 11783 2458 3896
(SE 3) 14.8 21.2 62.1 69.1 65.7 18.3 40.9
Time (s) 0.0267 0.162 0.134 1.58 5.67 0.29 1.75
2 RMSE means 0.099 0.367 0.0885 0.345 0.249 0.0913 0.0483
(SE 1) 2.87e-04 6.81e-04 6.59e-04 1.37e-03 2.44e-03 5.67e-04 2.46e-04
RMSE variances 9.91e-03 0.226 0.0194 0.181 0.069 0.0129 5.09e-03
(SE 2) 4.32e-06 4.31e-05 8.22e-05 1.08e-04 1.59e-03 1.77e-05 4.16e-05
KL 1928 10364 1625 9934 6886 1530 930
(SE 3) 9.17 9.73 14.6 41.6 72.2 14.8 10.5
Time (s) 0.0476 0.574 0.676 1.93 13.5 0.638 2.24
3 RMSE means 0.0462 0.317 0.0361 0.127 0.07 0.0189
(SE 1) 1.85e-04 5.14e-04 2.76e-04 6.51e-04 4.82e-04 3.58e-04
RMSE variances 2.21e-03 0.192 5.61e-03 0.0432 8.95e-03 2.81e-03
(SE 2) 1.73e-06 3.56e-05 3.77e-05 1.03e-04 1.59e-05 6.07e-05
KL 482 9344 319 3217 945 142
(SE 3) 3.61 7.54 3.85 15.7 10.1 9.75
Time (s) 0.111 1.42 1.88 5.23 1.54 4.71
4 RMSE means 0.0337 0.268 0.013 0.0403 0.0577 0.0144
(SE 1) 2.25e-04 4.81e-04 1.09e-04 3e-04 5.11e-04 4.22e-04
RMSE variances 1.19e-03 0.157 1.42e-03 7.95e-03 6.99e-03 2.33e-03
(SE 2) 1.73e-07 2.8e-05 1.5e-05 3.76e-05 1.82e-05 7.61e-05
KL 263 8210 41.6 435 656 90.5
(SE 3) 3.53 7.28 0.58 4.35 9.31 9.34
Time (s) 0.312 4.27 5.61 16.6 3.01 6.15
5 RMSE means 0.0187 0.229 5.7e-03 0.0216 0.0531 0.0109
(SE 1) 7.55e-05 4.36e-04 4.14e-05 2.01e-04 4.76e-04 5.52e-04
RMSE variances 3.92e-04 0.128 4.61e-04 3.06e-03 6.31e-03 1.89e-03
(SE 2) 4.15e-07 2.18e-05 6.17e-06 2.38e-05 1.72e-05 9.86e-05
KL 83.9 7174 7.71 121 559 60.3
(SE 3) 0.686 6.77 0.102 1.83 8.07 10.5
Time (s) 1.26 9.05 13.3 26.5 4.36 9.64
Table 27: Average RMSE between means, RMSE between variances, KL divergence between exact and approximate predictions, standard errors and time needed for making predictions on the training set on simulated datasets with an practical range of $0.5$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE means 0.176 0.514 0.387 0.569 0.622 0.129 0.0882
(SE 1) 4.59e-04 7.34e-03 4.6e-03 8.67e-03 7.86e-03 7.1e-04 5.47e-04
RMSE variances 0.0311 0.695 0.241 0.606 0.401 0.024 8.46e-03
(SE 2) 6.67e-06 2.22e-04 3.99e-04 3.14e-04 1e-02 6.22e-05 3.85e-05
KL 4408 14100 10679 14609 14555 2595 3982
(SE 3) 15.2 48.5 82.7 105 119 18.6 38.9
Time (s) 0.0277 0.18 0.053 1.53 5.65 1.72 1.78
2 RMSE means 0.0992 0.332 0.095 0.382 0.278 0.0978 0.0494
(SE 1) 2.66e-04 2.38e-03 8.33e-04 3.69e-03 3.16e-03 5.28e-04 2.71e-04
RMSE variances 0.01 0.489 0.026 0.353 0.0892 0.0151 5.53e-03
(SE 2) 5.05e-06 2.25e-04 1.47e-04 4.09e-04 2.43e-03 4.57e-05 4.27e-05
KL 1862 11754 1809 11309 7590 1631 952
(SE 3) 8.25 13.9 16.1 54.3 86.7 12.2 10.5
Time (s) 0.0517 0.616 0.724 1.89 13.1 2.63 2.27
3 RMSE means 0.0473 0.265 0.0399 0.129 0.0752 0.0193
(SE 1) 1.83e-04 9.01e-04 3.39e-04 9.64e-04 5.07e-04 3.58e-04
RMSE variances 2.32e-03 0.364 7.89e-03 0.0578 0.0106 3.09e-03
(SE 2) 3.04e-06 1.74e-04 6.48e-05 1.91e-04 3.44e-05 6.52e-05
KL 481 10306 387 3453 1014 146
(SE 3) 3.51 5.89 4.51 18 9.49 9.87
Time (s) 0.113 1.52 1.26 4.73 5.3 4.8
4 RMSE means 0.0353 0.22 0.0154 0.042 0.0616 0.0147
(SE 1) 2.34e-04 5.82e-04 1.45e-04 3.64e-04 4.88e-04 4.27e-04
RMSE variances 1.29e-03 0.262 2.28e-03 0.0104 8.17e-03 2.55e-03
(SE 2) 7.96e-07 1.19e-04 2.64e-05 5.65e-05 3.2e-05 8.25e-05
KL 274 8866 58.3 488 700 92.8
(SE 3) 3.64 4.72 0.807 4.94 8.17 9.65
Time (s) 0.321 4.35 3.86 16.1 9.47 6.27
5 RMSE means 0.0194 0.19 7.33e-03 0.0231 0.0565 0.011
(SE 1) 8.01e-05 4.74e-04 5.85e-05 2.32e-04 4.92e-04 5.59e-04
RMSE variances 4.19e-04 0.194 8.79e-04 4.2e-03 7.34e-03 2.06e-03
(SE 2) 4.69e-07 8.32e-05 1.36e-05 4.08e-05 3.19e-05 1.07e-04
KL 86.2 7640 12.7 140 595 61.7
(SE 3) 0.709 4.43 0.155 2.08 7.66 10.9
Time (s) 1.28 9.05 9.2 26.4 12.7 9.82
Table 28: Average RMSE between means, RMSE between variances, KL divergence between exact and approximate predictions, standard errors and time needed for making predictions on the test “interpolation” set on simulated datasets with an practical range of $0.5$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE means 0.179 0.476 0.355 0.449 1.43 0.29 0.122
(SE 1) 6.16e-03 0.0204 0.0115 0.0178 0.104 9.51e-03 7.82e-03
RMSE variances 0.05 0.419 0.254 0.392 0.418 0.178 0.194
(SE 2) 1.55e-04 4.78e-04 9.02e-04 6.23e-04 7.33e-03 6.57e-03 0.0147
KL 460 2451 1674 2290 28327 1024 411
(SE 3) 23.1 108 52.4 93.6 3257 43.1 28.4
Time (s) 0.0944 0.245 0.112 1.98 5.65 1.65 1.77
2 RMSE means 0.162 0.462 0.159 0.404 4.84 0.244 0.0732
(SE 1) 6.05e-03 0.02 6.04e-03 0.0131 0.402 8.36e-03 6.43e-03
RMSE variances 0.041 0.409 0.0779 0.327 7.57 0.114 0.177
(SE 2) 1.58e-04 4.37e-04 1.11e-03 7.59e-04 0.813 3.13e-03 0.0138
KL 357 2363 464 1988 26914 706 174
(SE 3) 21.4 103 21.2 61.2 2333 32.8 18.1
Time (s) 0.141 0.778 1.47 2.38 13.1 2.4 2.27
3 RMSE means 0.13 0.445 0.0915 0.249 0.2 0.0468
(SE 1) 4.38e-03 0.0194 3.74e-03 9.22e-03 6.72e-03 4.76e-03
RMSE variances 0.0265 0.396 0.0377 0.148 0.0825 0.151
(SE 2) 1.51e-04 4.11e-04 7.74e-04 1.46e-03 2.31e-03 0.0117
KL 217 2250 163 985 486 107
(SE 3) 12.1 97.7 9 37.7 21.1 14.8
Time (s) 0.261 1.94 2.51 6.45 4.52 4.79
4 RMSE means 0.0809 0.417 0.0461 0.147 0.161 0.0434
(SE 1) 2.44e-03 0.0185 2.28e-03 6.41e-03 5.86e-03 4.74e-03
RMSE variances 0.0122 0.371 0.0164 0.0664 0.0572 0.148
(SE 2) 8.51e-05 3.8e-04 3.74e-04 9.15e-04 1.4e-03 0.0113
KL 76.6 2053 41.2 355 336 103
(SE 3) 3.86 88.3 2.75 20.4 18.2 14.9
Time (s) 0.663 5.83 7.72 23.4 7.98 6.27
5 RMSE means 0.0479 0.38 0.0263 0.103 0.152 0.0429
(SE 1) 1.69e-03 0.0172 1.29e-03 4.92e-03 5.79e-03 4.79e-03
RMSE variances 5.28e-03 0.34 7.73e-03 0.0421 0.0503 0.146
(SE 2) 4.13e-05 3.43e-04 2.36e-04 7.88e-04 1.16e-03 0.0114
KL 24.2 1797 12.8 173 294 104
(SE 3) 1.63 76.5 1.01 12.1 17.8 15.6
Time (s) 2.54 13.2 18.4 37.6 11.1 9.83
Table 29: Average RMSE between means, RMSE between variances, KL divergence between exact and approximate predictions, standard errors and time needed for making predictions on the test “extrapolation” set on simulated datasets with an practical range of $0.5$ and a sample size of N=10’000.
B.3 Simulated data with $N=10^{\prime}000$ and practical range of $0.05$
Tuning parameter
Vecchia nb. neighbours 5 10 20 40 80
Tapering num. non-zero entries 11 30 60 130 263
FITC num. inducing points 47 254 500 950 1500
Full-scale num. inducing points 10 24 120 300 450
num. non-zero entries 5 8 28 100 150
FRK num. resolutions 1 2 exceeds time limit
MRA num. knots per partition 1 2 4 7 9
SPDE max edge 0.07 $0.05$ 0.03 0.025 0.022
Table 30: Tuning parameters chosen for the comparison on simulated datasets with an practical range of $0.05$ and a sample size of $N=10^{\prime}000$.
B.3.1 Log-likelihood evaluation
Figure 26: Absolute mean difference between approximate and exact log-likelihood on simulated datasets with an practical range of $0.05$ of and a sample size $N=10^{\prime}000$. The true data-generating parameters are used on the left, and two times these values are used on the right.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 40.6 723 1872 1158 571
(SE) 8.01 7.3 10.4 8.42 7.95
Time (s) 5.79e-03 0.0112 8.13e-03 5.26e-03 0.229
2 Value 6.93 343 1565 867 365
(SE) 7.82 6.62 9.36 7.67 8.08
Time (s) 0.0113 0.0652 0.0797 0.0109 0.39
3 Value 0.983 188 1228 326 235
(SE) 7.76 6.59 8.49 6.56 7.98
Time (s) 0.0193 0.228 0.269 0.116 0.735
4 Value 0.0421 93.4 801 96 168
(SE) 7.76 6.77 8.29 6.9 7.87
Time (s) 0.0419 0.767 0.736 0.809 1.26
5 Value 1.46e-03 48.3 483 56.8 143
(SE) 7.76 6.98 8.17 7 7.87
Time (s) 0.121 1.91 1.75 1.63 1.64
Table 31: Absolute mean difference between approximate and exact log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.05$ of and a sample size N=10’000. The true data- generating parameters are used
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 43.2 1394 2229 1788 553
(SE) 5.58 3.57 5.46 4.28 5.64
Time (s) 6.41e-03 9.75e-03 6.61e-03 5.17e-03 0.231
2 Value 2.77 920 1619 1504 346
(SE) 5.58 3.37 5.63 3.92 5.72
Time (s) 8.01e-03 0.0641 0.0783 0.0108 0.398
3 Value 5.46 656 1117 755 221
(SE) 5.6 3.58 6.25 3.76 5.7
Time (s) 0.0148 0.226 0.269 0.12 0.745
4 Value 0.923 438 617 266 157
(SE) 5.64 3.96 6.48 4.58 5.63
Time (s) 0.0376 0.769 0.735 0.809 1.28
5 Value 0.0566 297 323 154 134
(SE) 5.65 4.31 6.18 4.89 5.68
Time (s) 0.124 1.91 1.75 1.6 1.67
Table 32: Absolute mean difference between approximate and exact log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.05$ of and a sample size N=10’000. Log-likelihoods are evaluated at two-times the true data-generating parameters.
B.3.2 Parameter estimation
Figure 27: Bias and MSE of the estimates for the GP range, GP variance end error term variance on simulated datasets with an practical range of $0.05$ and a sample size N=100’000,
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias -1.54e-05 1.8 0.0391 0.0102 1.83e-03
(SE 1) 3.87e-05 0.0603 2.86e-03 1.94e-04 6.35e-05
MSE 1.5e-07 3.61 2.34e-03 1.07e-04 3.76e-06
(SE 2) 2.07e-08 0.223 3.4e-04 4.26e-06 2.43e-07
Time (s) 0.23 0.91 1.25 0.402 38.4 31.4 5.06
2 Bias -1.49e-05 0.814 0.0118 9.49e-03 9.41e-04
(SE 1) 3.66e-05 0.0481 3.65e-04 1.2e-04 4.94e-05
MSE 1.34e-07 0.893 1.53e-04 9.15e-05 1.13e-06
(SE 2) 1.74e-08 0.0829 9.45e-06 2.33e-06 1.04e-07
Time (s) 0.399 3.23 4.89 0.58 90.8 55 9.46
3 Bias -7.6e-06 0.1 6.97e-03 6.13e-03 5.6e-04
(SE 1) 3.7e-05 1.45e-03 1.55e-04 7.4e-05 4.42e-05
MSE 1.37e-07 0.0103 5.1e-05 3.82e-05 5.07e-07
(SE 2) 1.93e-08 3.08e-04 2.32e-06 9.07e-07 6.33e-08
Time (s) 0.839 7.5 15.2 5.56 98.3 22.6
4 Bias -4.01e-06 0.944 4.17e-03 2.67e-03 4.11e-04
(SE 1) 3.73e-05 0.0682 8.04e-05 5.15e-05 5.31e-05
MSE 1.39e-07 1.36 1.8e-05 7.38e-06 4.48e-07
(SE 2) 1.98e-08 0.217 6.63e-07 2.75e-07 1.35e-07
Time (s) 2.34 30 50.8 42.3 175 30.4
5 Bias -4.1e-06 2.27 2.52e-03 1.72e-03 3.31e-04
(SE 1) 3.72e-05 0.501 5.71e-05 4.47e-05 4.88e-05
MSE 1.38e-07 30.2 6.69e-06 3.15e-06 3.45e-07
(SE 2) 1.96e-08 10.3 2.83e-07 1.55e-07 8.24e-08
Time (s) 8.43 75.7 129 85.4 228 39.7
Table 33: Bias and MSE for the GP range, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.05$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias -1.53e-03 0.124 -0.72 0.238 -0.05
(SE 1) 3.12e-03 2.99e-03 0.0213 3.25e-03 3.31e-03
MSE 9.74e-04 0.0164 0.564 0.0579 3.58e-03
(SE 2) 1.48e-04 7.57e-04 0.0247 1.56e-03 3.41e-04
Time (s) 0.23 0.91 1.25 0.402 38.4 31.4 5.06
2 Bias -2.48e-03 0.0748 -0.435 0.17 -0.024
(SE 1) 3.12e-03 3.21e-03 9.15e-03 2.99e-03 3.23e-03
MSE 9.78e-04 6.62e-03 0.197 0.0297 1.61e-03
(SE 2) 1.32e-04 5.09e-04 7.49e-03 1.03e-03 2.05e-04
Time (s) 0.399 3.23 4.89 0.58 90.8 55 9.46
3 Bias -2.37e-03 0.175 -0.24 0.0901 -6.92e-03
(SE 1) 3.09e-03 2.99e-03 5.14e-03 3.06e-03 5.14e-03
MSE 9.6e-04 0.0314 0.0605 9.05e-03 2.66e-03
(SE 2) 1.34e-04 1.21e-03 2.44e-03 5.96e-04 1.41e-03
Time (s) 0.839 7.5 15.2 5.56 98.3 22.6
4 Bias -2.42e-03 0.693 -0.0876 0.0686 -3.96e-03
(SE 1) 3.09e-03 9.53e-03 4.1e-03 3.11e-03 6.71e-03
MSE 9.62e-04 0.489 9.35e-03 5.67e-03 4.47e-03
(SE 2) 1.34e-04 0.013 7.59e-04 4.56e-04 3.41e-03
Time (s) 2.34 30 50.8 42.3 175 30.4
5 Bias -2.49e-03 1.47 -0.0188 0.0603 -2.81e-03
(SE 1) 3.13e-03 0.0513 4.12e-03 3.19e-03 4.58e-03
MSE 9.83e-04 2.41 2.05e-03 4.65e-03 2.09e-03
(SE 2) 1.33e-04 0.112 2.66e-04 4.36e-04 1.09e-03
Time (s) 8.43 75.7 129 85.4 228 39.7
Table 34: Bias and MSE for the GP marginal variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.05$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias -4.67e-04 -0.269 0.729 -0.345 -0.022 0.0543 0.736
(SE 1) 1.48e-03 1.38e-03 0.0225 1.68e-03 2.19e-03 1.98e-03 2.56e-03
MSE 2.21e-04 0.0725 0.582 0.12 9.59e-04 3.34e-03 0.542
(SE 2) 3.05e-05 7.48e-04 0.0264 1.15e-03 1.11e-04 2.15e-04 3.78e-03
Time (s) 0.23 0.91 1.25 0.402 38.4 31.4 5.06
2 Bias -8.37e-06 -0.139 0.451 -0.286 -0.022 0.0274 0.595
(SE 1) 1.47e-03 1.45e-03 9.97e-03 1.32e-03 2.19e-03 1.8e-03 2.31e-03
MSE 2.15e-04 0.0195 0.213 0.082 9.59e-04 1.07e-03 0.354
(SE 2) 2.86e-05 4.11e-04 8.38e-03 7.56e-04 1.11e-04 1.09e-04 2.75e-03
Time (s) 0.399 3.23 4.89 0.58 90.8 55 9.46
3 Bias 9.35e-05 -0.0635 0.265 -0.146 0.0145 0.333
(SE 1) 1.49e-03 1.14e-03 5.21e-03 1.4e-03 1.75e-03 1.83e-03
MSE 2.22e-04 4.16e-03 0.0731 0.0215 5.14e-04 0.111
(SE 2) 2.99e-05 1.67e-04 2.78e-03 4.25e-04 5.9e-05 1.23e-03
Time (s) 0.839 7.5 15.2 5.56 98.3 22.6
4 Bias 1.85e-04 -0.0494 0.126 -0.0718 0.0112 0.253
(SE 1) 1.5e-03 2.05e-03 2.93e-03 1.42e-03 1.67e-03 1.62e-03
MSE 2.24e-04 2.86e-03 0.0169 5.35e-03 4.02e-04 0.0644
(SE 2) 3e-05 2.11e-04 7.46e-04 2.14e-04 5.02e-05 8.2e-04
Time (s) 2.34 30 50.8 42.3 175 30.4
5 Bias 2.12e-04 -0.074 0.0556 -0.0558 8.48e-03 0.2
(SE 1) 1.5e-03 4.13e-03 2.67e-03 1.45e-03 1.65e-03 1.62e-03
MSE 2.24e-04 7.19e-03 3.8e-03 3.33e-03 3.4e-04 0.0402
(SE 2) 3.02e-05 8.04e-04 2.77e-04 1.74e-04 4.32e-05 6.46e-04
Time (s) 8.43 75.7 129 85.4 228 39.7
Table 35: Bias and MSE for the error term variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.05$ and a sample size of N=10’000.
B.3.3 Prediction
Figure 28: Average RMSE and log-score on simulated datasets with an practical range of $0.05$ and a sample size N=100’000. Predictions are done using the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.447 0.501 0.574 0.529 0.576 0.491 0.854
(SE 1) 6.29e-04 5.55e-04 6.56e-04 5.94e-04 7.06e-04 6.26e-04 1.65e-03
Log-score 0.602 0.729 0.864 0.781 0.868 0.698 9.82
(SE 2) 1.41e-03 9.34e-04 1.13e-03 9.97e-04 1.26e-03 1.31e-03 0.0359
Time (s) 0.0303 0.146 0.171 1.57 5.79 0.26 1.73
2 RMSE 0.445 0.475 0.561 0.51 0.573 0.474 0.767
(SE 1) 6.36e-04 5.52e-04 6.57e-04 5.73e-04 6.69e-04 6.26e-04 1.48e-03
Log-score 0.599 0.676 0.84 0.747 0.862 0.662 4.73
(SE 2) 1.44e-03 9.52e-04 1.15e-03 9.62e-04 1.19e-03 1.34e-03 0.0199
Time (s) 0.0532 0.47 0.652 1.92 13.8 0.448 2.25
3 RMSE 0.445 0.463 0.545 0.474 0.464 0.599
(SE 1) 6.4e-04 5.61e-04 5.99e-04 5.78e-04 6.22e-04 1.16e-03
Log-score 0.598 0.649 0.81 0.673 0.64 1.53
(SE 2) 1.45e-03 1e-03 1.09e-03 1.01e-03 1.33e-03 6.37e-03
Time (s) 0.113 1.06 1.88 4.16 0.857 4.74
4 RMSE 0.445 0.455 0.519 0.455 0.459 0.551
(SE 1) 6.4e-04 5.75e-04 5.77e-04 5.79e-04 6.13e-04 9.2e-04
Log-score 0.598 0.628 0.76 0.628 0.628 1.11
(SE 2) 1.45e-03 1.07e-03 1.08e-03 1.08e-03 1.36e-03 4.35e-03
Time (s) 0.311 3.04 5.45 14.8 1.53 6.33
5 RMSE 0.445 0.451 0.496 0.451 0.457 0.523
(SE 1) 6.4e-04 5.89e-04 6.3e-04 5.85e-04 6.18e-04 9.32e-04
Log-score 0.598 0.616 0.711 0.617 0.624 0.922
(SE 2) 1.45e-03 1.14e-03 1.24e-03 1.14e-03 1.37e-03 3.26e-03
Time (s) 1.16 8.28 12.8 26.6 2.04 7.95
Table 36: Average RMSE and log-score on the training set, standard errors and time needed for making predictions on the training set on simulated datasets with an effective range of $0.05$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.591 0.693 0.991 0.796 0.995 0.695 0.884
(SE 1) 1.24e-03 1.43e-03 2.11e-03 1.64e-03 2.18e-03 1.26e-03 1.62e-03
Log-score 0.861 1.07 1.41 1.19 1.41 1.02 10.2
(SE 2) 2.08e-03 1.19e-03 2.08e-03 1.41e-03 2.2e-03 1.77e-03 0.0375
Time (s) 0.0352 0.165 0.443 1.53 5.83 1.63 1.76
2 RMSE 0.589 0.623 0.942 0.726 0.979 0.659 0.816
(SE 1) 1.24e-03 1.3e-03 1.89e-03 1.52e-03 2.05e-03 1.22e-03 1.51e-03
Log-score 0.857 0.973 1.35 1.11 1.4 0.963 5.05
(SE 2) 2.1e-03 1.19e-03 1.88e-03 1.28e-03 2.11e-03 1.88e-03 0.0244
Time (s) 0.0522 0.502 0.687 1.89 13.4 2.13 2.29
3 RMSE 0.588 0.603 0.886 0.623 0.635 0.694
(SE 1) 1.23e-03 1.26e-03 1.47e-03 1.31e-03 1.26e-03 1.34e-03
Log-score 0.857 0.928 1.29 0.969 0.927 1.82
(SE 2) 2.1e-03 1.29e-03 1.51e-03 1.24e-03 2.01e-03 7.55e-03
Time (s) 0.112 1.17 1.26 4.14 3.2 4.82
4 RMSE 0.588 0.595 0.809 0.596 0.622 0.669
(SE 1) 1.23e-03 1.24e-03 1.3e-03 1.24e-03 1.26e-03 1.37e-03
Log-score 0.857 0.897 1.19 0.899 0.907 1.4
(SE 2) 2.1e-03 1.43e-03 1.4e-03 1.45e-03 2.08e-03 6.44e-03
Time (s) 0.323 3.1 3.77 14.9 5.01 6.45
5 RMSE 0.588 0.591 0.745 0.593 0.618 0.656
(SE 1) 1.23e-03 1.23e-03 1.28e-03 1.24e-03 1.27e-03 1.47e-03
Log-score 0.857 0.88 1.11 0.885 0.901 1.21
(SE 2) 2.1e-03 1.57e-03 1.41e-03 1.56e-03 2.08e-03 4.22e-03
Time (s) 1.16 8.37 8.87 26.7 6.25 8.1
Table 37: Average RMSE and log-score on the test “interpolation” set, standard errors and time needed for making predictions on the test “interpolation” set on simulated datasets with an practical range of $0.05$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.996 0.999 1 1 1.03 0.998 1.01
(SE 1) 3.79e-03 3.8e-03 3.83e-03 3.82e-03 8.81e-03 3.8e-03 3.9e-03
Log-score 1.41 1.42 1.42 1.42 1.45 1.42 2.53
(SE 2) 3.78e-03 3.8e-03 3.83e-03 3.82e-03 8.91e-03 3.8e-03 0.032
Time (s) 0.105 0.231 0.818 1.99 5.82 1.57 1.76
2 RMSE 0.996 0.998 1 1 1.05 0.998 1
(SE 1) 3.79e-03 3.79e-03 3.82e-03 3.81e-03 0.0181 3.79e-03 3.84e-03
Log-score 1.41 1.42 1.42 1.42 1.47 1.42 1.81
(SE 2) 3.78e-03 3.79e-03 3.83e-03 3.81e-03 0.0131 3.79e-03 0.0129
Time (s) 0.149 0.682 1.37 2.37 13.5 2.07 2.29
3 RMSE 0.996 0.997 1 0.998 0.997 0.997
(SE 1) 3.79e-03 3.79e-03 3.82e-03 3.79e-03 3.81e-03 3.88e-03
Log-score 1.41 1.42 1.42 1.42 1.42 1.5
(SE 2) 3.78e-03 3.78e-03 3.83e-03 3.79e-03 3.81e-03 7.08e-03
Time (s) 0.263 1.6 2.52 5.23 3.09 4.82
4 RMSE 0.996 0.996 0.999 0.996 0.997 0.997
(SE 1) 3.79e-03 3.79e-03 3.78e-03 3.79e-03 3.83e-03 3.8e-03
Log-score 1.41 1.41 1.42 1.41 1.42 1.49
(SE 2) 3.78e-03 3.78e-03 3.78e-03 3.78e-03 3.82e-03 6.56e-03
Time (s) 0.671 4.4 7.54 20.7 4.85 6.45
5 RMSE 0.996 0.996 0.999 0.996 0.997 0.997
(SE 1) 3.79e-03 3.79e-03 3.76e-03 3.78e-03 3.84e-03 3.83e-03
Log-score 1.41 1.41 1.42 1.41 1.41 1.49
(SE 2) 3.78e-03 3.77e-03 3.76e-03 3.77e-03 3.83e-03 5.88e-03
Time (s) 2.31 12.3 17.7 38.1 6.07 8.1
Table 38: Average RMSE and log-score on the test “extrapolation” set, standard errors and time needed for making predictions on the test “extrapolation” set on simulated datasets with an practical range of $0.05$ and a sample size of N=10’000.
B.3.4 Comparison to exact calculations
Figure 29: RMSE between exact and approximate predictive means and variances on simulated datasets with an practical range of $0.05$ and a sample size N=100’000. Every approximation makes predictions using its parameter estimates.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE means 0.0351 0.229 0.363 0.285 0.366 0.206 0.729
(SE 1) 7.03e-05 3.28e-04 5.52e-04 4.31e-04 6.31e-04 4.39e-04 1.68e-03
RMSE variances 1.78e-03 0.102 0.14 0.12 0.132 0.06 0.163
(SE 2) 4.17e-06 2.13e-05 6.8e-05 3.25e-05 1.13e-03 8.17e-05 1.24e-04
KL 31.2 1303 2660 1828 2698 991 92437
(SE 3) 0.129 2.41 6.04 3.76 7.29 3.61 312
Time (s) 0.0303 0.146 0.171 1.57 5.79 0.26 1.73
2 RMSE means 0.0135 0.165 0.341 0.25 0.361 0.164 0.626
(SE 1) 3.05e-05 2.22e-04 4.97e-04 3.69e-04 6.42e-04 4.1e-04 1.53e-03
RMSE variances 2.47e-04 0.0787 0.131 0.109 0.128 0.0425 0.138
(SE 2) 5.7e-07 1.4e-05 7.31e-05 2.49e-05 1.1e-03 8.59e-05 1.71e-04
KL 4.52 777 2415 1484 2641 641 41484
(SE 3) 0.021 1.29 5.21 2.91 7.38 2.92 145
Time (s) 0.0532 0.47 0.652 1.92 13.8 0.448 2.25
3 RMSE means 2.56e-03 0.127 0.314 0.162 0.132 0.401
(SE 1) 7.14e-06 1.69e-04 4.47e-04 2.17e-04 3.78e-04 1.02e-03
RMSE variances 1.12e-05 0.0623 0.118 0.0764 0.0319 0.0876
(SE 2) 5.67e-08 1.08e-05 7.58e-05 1.78e-05 6.84e-05 2.29e-04
KL 0.16 507 2115 747 418 9352
(SE 3) 9.05e-04 0.819 4.69 1.24 2.23 34.6
Time (s) 0.113 1.06 1.88 4.16 0.857 4.74
4 RMSE means 2.86e-04 0.0935 0.267 0.0947 0.11 0.324
(SE 1) 1.01e-06 1.24e-04 4.75e-04 1.34e-04 3.53e-04 8.21e-04
RMSE variances 1.44e-07 0.0464 0.0953 0.0456 0.0258 0.0715
(SE 2) 9.48e-10 8.13e-06 6.75e-05 1.51e-05 6.2e-05 1.57e-04
KL 1.96e-03 298 1612 295 294 5113
(SE 3) 1.46e-05 0.476 4.46 0.516 1.79 27.8
Time (s) 0.311 3.04 5.45 14.8 1.53 6.33
5 RMSE means 1.15e-05 0.0691 0.219 0.0749 0.102 0.274
(SE 1) 5.86e-08 9.19e-05 3.7e-04 1.16e-04 3.45e-04 8.02e-04
RMSE variances 2.76e-10 0.0345 0.0719 0.0356 0.0235 0.063
(SE 2) 3.24e-12 6.13e-06 6.42e-05 1.56e-05 6e-05 1.25e-04
KL 3.13e-06 174 1131 191 252 3229
(SE 3) 3.34e-08 0.277 3.06 0.381 1.62 18.9
Time (s) 1.16 8.28 12.8 26.6 2.04 7.95
Table 39: Average RMSE between means, RMSE between variances, KL divergence between exact and approximate predictions, standard errors and time needed for making predictions on the training set on simulated datasets with an practical range of $0.05$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE means 0.0545 0.363 0.796 0.533 0.801 0.369 0.66
(SE 1) 1.49e-04 1.06e-03 2.13e-03 1.49e-03 2.2e-03 8.43e-04 1.65e-03
RMSE variances 4.44e-03 0.429 0.656 0.532 0.675 0.209 0.327
(SE 2) 1.65e-05 1.33e-04 2.5e-04 1.95e-04 6.19e-03 3.26e-04 3.89e-04
KL 41 2160 5502 3347 5563 1601 93274
(SE 3) 0.19 4.49 18 8.9 18.7 5.55 322
Time (s) 0.0352 0.165 0.443 1.53 5.83 1.63 1.76
2 RMSE means 0.022 0.203 0.734 0.423 0.78 0.294 0.567
(SE 1) 6.45e-05 6.57e-04 1.84e-03 1.19e-03 2.1e-03 8.19e-04 1.43e-03
RMSE variances 7.29e-04 0.303 0.607 0.466 0.642 0.148 0.298
(SE 2) 2.73e-06 8.28e-05 2.35e-04 1.56e-04 6.01e-03 2.83e-04 4.16e-04
KL 6.54 1151 4953 2557 5392 1053 42074
(SE 3) 0.0333 1.64 14.7 5.75 17.9 4.53 152
Time (s) 0.0522 0.502 0.687 1.89 13.4 2.13 2.29
3 RMSE means 4.3e-03 0.133 0.661 0.203 0.238 0.368
(SE 1) 1.79e-05 4.37e-04 1.48e-03 6.57e-04 7.31e-04 9.63e-04
RMSE variances 3.59e-05 0.226 0.545 0.294 0.111 0.223
(SE 2) 2.84e-07 6.25e-05 2.67e-04 9.1e-05 2.69e-04 5.38e-04
KL 0.24 708 4313 1117 697 9620
(SE 3) 1.71e-03 0.759 11.7 1.73 3.35 35.5
Time (s) 0.112 1.17 1.26 4.14 3.2 4.82
4 RMSE means 4.97e-04 0.0848 0.554 0.0964 0.2 0.317
(SE 1) 1.94e-06 2.73e-04 1.42e-03 2.84e-04 6.36e-04 1.15e-03
RMSE variances 4.97e-07 0.159 0.44 0.162 0.0906 0.189
(SE 2) 5.07e-09 4.65e-05 2.62e-04 6.12e-05 2.75e-04 4.66e-04
KL 3.13e-03 398 3358 416 499 5400
(SE 3) 2.36e-05 0.334 10.4 0.389 2.52 26.4
Time (s) 0.323 3.1 3.77 14.9 5.01 6.45
5 RMSE means 2.04e-05 0.0575 0.457 0.0756 0.186 0.288
(SE 1) 1.13e-07 1.8e-04 1.26e-03 2.15e-04 6.83e-04 1.76e-03
RMSE variances 9.99e-10 0.113 0.342 0.126 0.0832 0.173
(SE 2) 2.43e-11 3.48e-05 2.8e-04 5.36e-05 2.7e-04 4.27e-04
KL 5.17e-06 226 2501 274 433 3485
(SE 3) 5.35e-08 0.164 7.88 0.274 2.46 14.9
Time (s) 1.16 8.37 8.87 26.7 6.25 8.1
Table 40: Average RMSE between means, RMSE between variances, KL divergence between exact and approximate predictions, standard errors and time needed for making predictions on the test “interpolation” set on simulated datasets with an practical range of $0.05$ and a sample size of N=10’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE means 5.94e-03 0.0839 0.106 0.097 0.243 0.0684 0.15
(SE 1) 1.04e-04 1.31e-03 1.79e-03 1.54e-03 0.0206 1.48e-03 2.62e-03
RMSE variances 2.27e-04 0.0516 0.0619 0.0577 0.0789 0.0308 0.739
(SE 2) 5.52e-06 3.12e-04 4.84e-04 3.97e-04 3.08e-03 8.07e-04 3.51e-03
KL 0.235 47.1 71.5 61 404 29.8 11236
(SE 3) 7.75e-03 1.22 2.1 1.65 68.1 1.17 236
Time (s) 0.105 0.231 0.818 1.99 5.82 1.57 1.76
2 RMSE means 1.4e-03 0.062 0.105 0.0896 0.303 0.0619 0.114
(SE 1) 2.84e-05 9.89e-04 1.74e-03 1.41e-03 0.0327 1.33e-03 2.03e-03
RMSE variances 1.65e-05 0.0413 0.0613 0.0543 0.198 0.0275 0.585
(SE 2) 9.02e-07 2.15e-04 4.82e-04 3.45e-04 0.0494 8.06e-04 3.94e-03
KL 0.0127 27.8 69.7 52.9 575 24.3 3916
(SE 3) 5.14e-04 0.692 2.04 1.39 123 0.955 78.8
Time (s) 0.149 0.682 1.37 2.37 13.5 2.07 2.29
3 RMSE means 2.03e-04 0.0465 0.0993 0.0628 0.0567 0.0664
(SE 1) 5.85e-06 7.53e-04 1.62e-03 1.01e-03 1.46e-03 1.11e-03
RMSE variances 4.18e-07 0.0331 0.0584 0.0417 0.0247 0.368
(SE 2) 2.78e-08 1.6e-04 4.57e-04 2.19e-04 8.33e-04 4.56e-03
KL 2.67e-04 16.8 63.9 28.4 20.5 874
(SE 3) 1.63e-05 0.404 1.79 0.717 0.97 27.9
Time (s) 0.263 1.6 2.52 5.23 3.09 4.82
4 RMSE means 1.52e-05 0.0326 0.0889 0.0365 0.0521 0.0561
(SE 1) 6.3e-07 5.32e-04 1.57e-03 5.94e-04 1.51e-03 8.05e-04
RMSE variances 3.24e-09 0.0247 0.0513 0.027 0.0226 0.365
(SE 2) 3.5e-10 1.12e-04 4.82e-04 1.23e-04 8.07e-04 5.36e-03
KL 1.57e-06 8.9 52.9 10.9 17.2 788
(SE 3) 1.34e-07 0.206 1.58 0.253 0.92 25.9
Time (s) 0.671 4.4 7.54 20.7 4.85 6.45
5 RMSE means 4.62e-07 0.0232 0.0784 0.0298 0.0493 0.056
(SE 1) 3e-08 3.79e-04 1.55e-03 4.87e-04 1.54e-03 1.07e-03
RMSE variances 3.42e-12 0.0184 0.0432 0.0223 0.0217 0.356
(SE 2) 5.44e-13 8.01e-05 4.03e-04 1.05e-04 7.65e-04 5.59e-03
KL 1.57e-09 4.79 42.3 7.44 15.5 704
(SE 3) 2.25e-10 0.107 1.4 0.173 0.908 23.2
Time (s) 2.31 12.3 17.7 38.1 6.07 8.1
Table 41: Average RMSE between means, RMSE between variances, KL divergence between exact and approximate predictions, standard errors and time needed for making predictions on the test “extrapolation” set on simulated datasets with an practical range of $0.05$ and a sample size of N=10’000.
B.4 Simulated data with $N=100^{\prime}000$ and practical range of $0.2$
B.4.1 Log-likelihood evaluation
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 116892 148045 156316 144010 112921
(SE) 55.6 469 545 358 62.2
Time (s) 0.0521 0.0645 0.0615 0.107 3.49
2 Value 113351 139445 126797 130065 112083
(SE) 56.1 308 208 242 60.3
Time (s) 0.0778 0.541 0.468 0.478 7.97
3 Value 111603 136654 122397 125538 111784
(SE) 52.9 259 162 124 60.2
Time (s) 0.154 1.09 0.996 0.901 15.9
4 Value 111025 130565 112694 117692 111617
(SE) 54.2 163 100 81.6 60.2
Time (s) 0.392 5.15 5.46 5 24.1
5 Value 110922 122924 111121 112154 111511
(SE) 55.3 75.3 69.6 64.2 61
Time (s) 1.26 42.7 41.5 29.9 49.4
Table 42: Mean approximate negative log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.2$ of and a sample size N=100’000. Log-likelihoods are evaluated at the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 126154 162289 160526 151451 121842
(SE) 28.3 234 348 242 46.9
Time (s) 0.0532 0.0629 0.0579 0.101 4.15
2 Value 122577 155846 130399 136310 121216
(SE) 29.4 154 149 215 45.9
Time (s) 0.0804 0.543 0.468 0.481 14.7
3 Value 120782 153565 127134 132367 120985
(SE) 29.3 129 108 123 45.8
Time (s) 0.153 1.1 0.974 0.911 31.6
4 Value 120185 148134 121085 125834 120851
(SE) 35.1 80.9 59.7 88.2 46.6
Time (s) 0.392 5.15 5.44 5.02 46
5 Value 120195 140101 120356 121315 120767
(SE) 41.3 37 48.7 54.7 47.5
Time (s) 1.24 43.2 41 29.9 86.1
Table 43: Mean approximate negative log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.2$ of and a sample size N=100’000. Log-likelihoods are evaluated at two-times the true data-generating parameters.
B.4.2 Parameter estimation
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias 9.92e-04 0.254 0.0658 0.0136 7.97e-03
(SE 1) 6.83e-04 0.0393 7.37e-03 1.94e-03 9.62e-04
MSE 1.03e-05 0.0953 5.42e-03 2.61e-04 8.11e-05
(SE 2) 3.89e-06 0.0229 1.07e-03 6.03e-05 1.47e-05
Time (s) 1.96 21.1 1.8 17.8 319 287 15.9
2 Bias 1.25e-03 0.622 0.113 0.0147 7.49e-03
(SE 1) 5.7e-04 0.113 0.0269 1.92e-03 8.42e-04
MSE 8.05e-06 0.644 0.0271 2.88e-04 6.96e-05
(SE 2) 2.44e-06 0.181 0.0134 7.13e-05 1.21e-05
Time (s) 4.19 90.5 29.8 60.5 636 653 53.5
3 Bias 1.05e-03 0.933 0.173 0.0176 6.26e-03
(SE 1) 6.09e-04 0.093 0.0758 1.86e-03 1.03e-03
MSE 8.51e-06 1.04 0.145 3.8e-04 5.94e-05
(SE 2) 2.94e-06 0.146 0.104 7.6e-05 1.07e-05
Time (s) 7.07 273 73.4 111 1482 1357
4 Bias 9.61e-04 1.1 0.0164 0.0146 6.66e-03
(SE 1) 6.21e-04 0.13 1.31e-03 1.56e-03 5.8e-04
MSE 8.64e-06 1.55 3.03e-04 2.62e-04 5.08e-05
(SE 2) 2.41e-06 0.327 4.97e-05 4.81e-05 8.24e-06
Time (s) 27.2 774 431 602 2013 2330
5 Bias 9.48e-04 1.36 1.84e-03 6.87e-03 6.01e-03
(SE 1) 6.08e-04 0.181 5.61e-04 7.92e-04 7.87e-04
MSE 8.3e-06 2.52 9.67e-06 5.98e-05 4.79e-05
(SE 2) 2.51e-06 0.848 2.2e-06 1.22e-05 8.65e-06
Time (s) 91.3 3636 1872 3287 4218 4337
Table 44: Bias and MSE for the GP range, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.2$ and a sample size of N=100’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias 0.0465 0.137 -0.413 0.151 0.294
(SE 1) 0.0249 0.0184 0.0553 0.0232 0.0465
MSE 0.0145 0.0256 0.232 0.0335 0.127
(SE 2) 4.63e-03 6.09e-03 0.0449 8.07e-03 0.025
Time (s) 1.96 21.1 1.8 17.8 319 287 15.9
2 Bias 0.0482 -0.0582 17.9 0.374 0.356
(SE 1) 0.0249 0.0142 10.3 0.0788 0.045
MSE 0.0147 7.4e-03 2434 0.264 0.165
(SE 2) 4.47e-03 1.6e-03 2024 0.0908 0.0252
Time (s) 4.19 90.5 29.8 60.5 636 653 53.5
3 Bias 0.0324 -0.126 374 0.558 0.307
(SE 1) 0.0227 0.013 305 0.0913 0.0497
MSE 0.0113 0.0192 1994764 0.479 0.141
(SE 2) 4.48e-03 3.01e-03 1868037 0.148 0.0247
Time (s) 7.07 273 73.4 111 1482 1357
4 Bias 0.0294 -0.281 0.911 0.666 0.342
(SE 1) 0.024 0.0116 0.109 0.0959 0.0338
MSE 0.0124 0.0816 1.07 0.628 0.139
(SE 2) 4.38e-03 6.21e-03 0.274 0.157 0.0193
Time (s) 27.2 774 431 602 2013 2330
5 Bias 0.0256 -0.484 0.129 0.374 0.302
(SE 1) 0.0229 6.79e-03 0.0228 0.0493 0.0418
MSE 0.0111 0.236 0.027 0.188 0.125
(SE 2) 5.31e-03 6.53e-03 6.1e-03 0.0464 0.0214
Time (s) 91.3 3636 1872 3287 4218 4337
Table 45: Bias and MSE for the GP marginal variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.2$ and a sample size of N=100’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias 1.67e-04 -0.348 0.505 -0.222 -0.0135 1.29e-03 0.0354
(SE 1) 5.55e-04 3.25e-03 0.0328 5.75e-03 1.42e-03 5.74e-04 9.8e-04
MSE 6.2e-06 0.121 0.277 0.0498 2.2e-04 7.92e-06 1.27e-03
(SE 2) 3e-06 2.26e-03 0.0339 2.53e-03 4.22e-05 2.59e-06 7.15e-05
Time (s) 1.96 21.1 1.8 17.8 319 287 15.9
2 Bias 2.27e-04 -0.268 0.116 -0.139 -0.0126 6.24e-04 5.7e-03
(SE 1) 5.66e-04 2.43e-03 4.38e-03 2.03e-03 1.47e-03 5.98e-04 5.2e-04
MSE 6.45e-06 0.0718 0.0139 0.0194 1.99e-04 7.19e-06 3.76e-05
(SE 2) 2.99e-06 1.3e-03 9.6e-04 5.62e-04 4.79e-05 3.09e-06 6.12e-06
Time (s) 4.19 90.5 29.8 60.5 636 653 53.5
3 Bias 1.18e-04 -0.24 0.0723 -0.115 7.42e-04 1.39e-03
(SE 1) 5.12e-04 2.31e-03 3.44e-03 1.42e-03 4.99e-04 5.25e-04
MSE 5.25e-06 0.0575 5.46e-03 0.0132 5.27e-06 7.17e-06
(SE 2) 2.75e-06 1.11e-03 4.7e-04 3.27e-04 1.89e-06 2.41e-06
Time (s) 7.07 273 73.4 111 1482 1357
4 Bias 1.16e-04 -0.177 -6.24e-03 -0.0689 7.37e-04 1.46e-03
(SE 1) 4.99e-04 2.18e-03 7.67e-04 9.15e-04 5.17e-04 5.22e-04
MSE 4.99e-06 0.0316 5.07e-05 4.76e-03 5.62e-06 7.31e-06
(SE 2) 2.63e-06 7.8e-04 8.84e-06 1.27e-04 2.24e-06 2.38e-06
Time (s) 27.2 774 431 602 2013 2330
5 Bias 1.51e-04 -0.105 -6.13e-03 -0.0254 5.57e-04 1.52e-03
(SE 1) 5.08e-04 9.91e-04 6.28e-04 6.41e-04 5.22e-04 5.2e-04
MSE 5.18e-06 0.011 4.55e-05 6.53e-04 5.48e-06 7.45e-06
(SE 2) 2.57e-06 2.07e-04 7.25e-06 3.54e-05 2.18e-06 2.38e-06
Time (s) 91.3 3636 1872 3287 4218 4337
Table 46: Bias and MSE for the error term variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.2$ and a sample size of N=100’000.
B.4.3 Prediction
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.224 0.486 0.559 0.477 0.555 0.164 0.199
(SE 1) 5.62e-04 1.38e-03 2.09e-03 1.54e-03 3.19e-03 9.71e-04 2.53e-03
Log-score -0.0785 0.705 0.874 0.708 0.83 -0.411 2.32
(SE 2) 2.47e-03 2.17e-03 5.63e-03 4.3e-03 5.75e-03 5.95e-03 0.0943
Time (s) 0.395 3.56 9.39 179 101 5.12 5.27
2 RMSE 0.168 0.433 0.399 0.391 0.477 0.148 0.133
(SE 1) 7.1e-04 7.52e-04 2.27e-03 8.34e-04 3.64e-03 1.08e-03 1.55e-03
Log-score -0.368 0.606 0.566 0.539 0.679 -0.507 -0.542
(SE 2) 4.17e-03 1.15e-03 4.36e-03 4.59e-03 7.57e-03 7.35e-03 0.0169
Time (s) 0.687 12.2 11.8 224 193 15 12.3
3 RMSE 0.137 0.415 0.355 0.352 0.143 0.124
(SE 1) 9.81e-04 6.16e-04 2.17e-03 7.2e-04 1.13e-03 1.47e-03
Log-score -0.575 0.57 0.445 0.439 -0.543 -0.671
(SE 2) 7.14e-03 9.47e-04 6.18e-03 2.99e-03 7.88e-03 0.0132
Time (s) 1.48 32.2 16.2 264 35.1 254
4 RMSE 0.129 0.374 0.195 0.263 0.14 0.124
(SE 1) 1.06e-03 4.7e-04 2.1e-03 8.37e-04 1.13e-03 1.47e-03
Log-score -0.634 0.483 -0.189 0.16 -0.564 -0.671
(SE 2) 8.02e-03 7.53e-04 9.57e-03 3e-03 8.15e-03 0.0132
Time (s) 4.21 85.7 51.6 600 60.4 449
5 RMSE 0.127 0.313 0.139 0.166 0.138 0.124
(SE 1) 1.14e-03 4.51e-04 1.74e-03 9.66e-04 1.17e-03 1.47e-03
Log-score -0.647 0.334 -0.538 -0.304 -0.578 -0.671
(SE 2) 8.84e-03 7.72e-04 9.71e-03 4.15e-03 8.5e-03 0.0133
Time (s) 13.5 404 274 1152 127 859
Table 47: Average RMSE and log-score on the training set, standard errors and time needed for making predictions on the training set on simulated datasets with an effective range of $0.2$ and a sample size of N=100’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.225 0.622 0.922 0.606 0.897 0.168 0.201
(SE 1) 5.44e-04 7.81e-03 0.0101 7.38e-03 0.0127 1.13e-03 2.53e-03
Log-score -0.0714 1.01 1.37 1 1.31 -0.386 2.35
(SE 2) 2.37e-03 5.79e-03 0.0113 7.25e-03 0.0142 6.48e-03 0.0934
Time (s) 0.453 3.6 2.17 176 99.7 32.4 5.46
2 RMSE 0.169 0.449 0.478 0.415 0.648 0.152 0.135
(SE 1) 6.31e-04 3.78e-03 5.36e-03 2.52e-03 8.74e-03 1.12e-03 1.61e-03
Log-score -0.358 0.818 0.756 0.707 0.984 -0.486 -0.524
(SE 2) 3.68e-03 2.46e-03 7.04e-03 6.33e-03 0.0133 7.53e-03 0.0174
Time (s) 0.757 12.4 3.91 222 188 50.4 12.6
3 RMSE 0.139 0.406 0.406 0.356 0.146 0.126
(SE 1) 9.88e-04 2.74e-03 4.8e-03 1.58e-03 1.14e-03 1.53e-03
Log-score -0.56 0.758 0.589 0.573 -0.522 -0.655
(SE 2) 7.07e-03 1.77e-03 7.58e-03 4.11e-03 7.93e-03 0.0135
Time (s) 1.57 33.4 6.96 262 72.6 260
4 RMSE 0.131 0.332 0.202 0.257 0.143 0.126
(SE 1) 1.11e-03 1.23e-03 2.59e-03 1.43e-03 1.14e-03 1.53e-03
Log-score -0.617 0.624 -0.142 0.232 -0.544 -0.655
(SE 2) 8.24e-03 8.8e-04 0.0102 2.91e-03 8.15e-03 0.0135
Time (s) 4.25 86.4 31.5 598 103 459
5 RMSE 0.129 0.269 0.143 0.165 0.141 0.126
(SE 1) 1.19e-03 6.05e-04 1.89e-03 1.13e-03 1.19e-03 1.53e-03
Log-score -0.631 0.427 -0.509 -0.27 -0.558 -0.655
(SE 2) 8.97e-03 6.07e-04 9.69e-03 4.27e-03 8.62e-03 0.0136
Time (s) 13.8 400 197 1152 169 877
Table 48: Average RMSE and log-score on the test “interpolation” set, standard errors and time needed for making predictions on the test “interpolation” set on simulated datasets with an practical range of $0.2$ and a sample size of N=100’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.927 0.987 0.987 0.986 1.54 0.946 0.933
(SE 1) 0.0199 0.0183 0.018 0.0177 0.139 0.018 0.0264
Log-score 1.31 1.41 1.41 1.41 2.4 1.34 1.34
(SE 2) 0.0168 0.018 0.0177 0.0174 0.341 0.0138 0.0279
Time (s) 4.96 6.53 4.24 222 98.7 30.9 5.44
2 RMSE 0.925 0.985 0.965 0.978 41.8 0.943 0.922
(SE 1) 0.0203 0.0183 0.0174 0.0168 13.2 0.0193 0.0266
Log-score 1.3 1.41 1.38 1.4 71.6 1.34 1.3
(SE 2) 0.0171 0.018 0.0164 0.0164 19.4 0.0159 0.0233
Time (s) 5.53 16.1 7.81 270 188 46.8 12.6
3 RMSE 0.924 0.985 0.957 0.973 0.937 0.919
(SE 1) 0.0206 0.0183 0.0179 0.0178 0.0209 0.0257
Log-score 1.3 1.41 1.37 1.39 1.33 1.29
(SE 2) 0.0173 0.018 0.0166 0.0173 0.018 0.0238
Time (s) 7.15 37.9 13.9 313 74.4 260
4 RMSE 0.921 0.983 0.932 0.959 0.936 0.919
(SE 1) 0.0215 0.0183 0.0208 0.0184 0.021 0.0256
Log-score 1.29 1.4 1.33 1.37 1.33 1.29
(SE 2) 0.0189 0.0179 0.019 0.0173 0.0181 0.0238
Time (s) 12.6 96.7 62.4 698 110 458
5 RMSE 0.919 0.977 0.922 0.938 0.934 0.919
(SE 1) 0.0226 0.0182 0.0226 0.0202 0.0217 0.0256
Log-score 1.29 1.4 1.3 1.34 1.32 1.29
(SE 2) 0.0207 0.0178 0.0208 0.0184 0.0189 0.0238
Time (s) 31.7 468 396 1364 178 877
Table 49: Average RMSE and log-score on the test “extrapolation” set, standard errors and time needed for making predictions on the test “extrapolation” set on simulated datasets with an practical range of $0.2$ and a sample size of N=100’000.
B.5 Simulated data with $N=100^{\prime}000$ and practical range of $0.5$
Tuning parameter
Vecchia nb. neighbours 5 10 20 40 80
Tapering num. non-zero entries 8 17 22 40 111
FITC num. inducing points 20 200 275 900 2000
Full-scale num. inducing points 26 91 122 250 650
num. non-zero entries 6 11 15 27 76
FRK num. resolutions 1 2 exceeds time limit
MRA num. knots per partition 1 2 3 4 5
SPDE max edge 0.04 0.02 0.005 0.004 0.003
Table 50: Tuning parameters chosen for the comparison on simulated datasets with an practical range of $0.5$ and a sample size of $N=100^{\prime}000$.
B.5.1 Log-likelihood evaluation
Figure 30: Average negative log-likelihood on simulated datasets with an practical range $0.5$ of and a sample size N=100’000. The true data-generating parameters are used on the left, and two times these values are used on the right.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 115906 147184 134682 127030 109523
(SE) 63.1 951 815 388 56
Time (s) 0.0497 0.0627 0.0344 0.101 5.43
2 Value 112200 138802 110408 113648 109143
(SE) 48.5 616 90.5 137 54.3
Time (s) 0.0738 0.551 0.637 0.485 20.3
3 Value 110154 136072 109635 111886 109017
(SE) 48.1 515 70.7 91.4 53.4
Time (s) 0.146 1.58 0.773 0.912 38.6
4 Value 109163 130090 108700 109577 108943
(SE) 44.5 318 52.7 70.4 53.5
Time (s) 0.394 5.3 8.23 3.79 52.1
5 Value 108789 122479 108648 108746 108899
(SE) 48 131 50.2 56.3 52.9
Time (s) 1.25 53.6 47.2 30.4 87.9
Table 51: Mean approximate negative log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.5$ of and a sample size N=100’000. Log-likelihoods are evaluated at the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 125459 161867 137360 132517 118933
(SE) 32.1 476 568 324 39.4
Time (s) 0.0505 0.0624 0.0352 0.0972 7.38
2 Value 121725 155542 119092 121496 118636
(SE) 25.2 308 48.4 88.4 35.6
Time (s) 0.0743 0.547 0.636 0.49 23.9
3 Value 119654 153294 118672 120296 118532
(SE) 25.6 257 39.2 60.3 35
Time (s) 0.142 1.19 0.771 0.946 45.2
4 Value 118639 147924 118248 118735 118473
(SE) 24 159 34.3 41.5 35.8
Time (s) 0.392 5.1 8.23 3.95 60.4
5 Value 118250 139923 118224 118273 118436
(SE) 26.5 65.3 33.7 35.2 35.6
Time (s) 1.24 53.7 47.2 30.3 101
Table 52: Mean approximate negative log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.5$ of and a sample size N=100’000. Log-likelihoods are evaluated at two-times the true data-generating parameters.
B.5.2 Parameter estimation
Figure 31: Bias and MSE of the estimates for the GP range, GP variance end error term variance on simulated datasets with an practical range of $0.5$ and a sample size N=100’000,
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias -3.93e-03 0.0925 5.21 0.0407 0.0246
(SE 1) 2.61e-03 0.029 3.75 0.0108 2.58e-03
MSE 1.52e-04 0.0254 308 3.98e-03 7.33e-04
(SE 2) 4.46e-05 0.0126 261 1.61e-03 1.23e-04
Time (s) 1.72 17.5 1.91 22.5 309 367 13.3
2 Bias -2.31e-03 0.702 0.108 0.0743 0.022
(SE 1) 3.19e-03 0.121 0.0488 0.0141 2.53e-03
MSE 2.09e-04 0.784 0.0594 9.51e-03 6.06e-04
(SE 2) 7.12e-05 0.191 0.0528 3.17e-03 1.08e-04
Time (s) 2.93 92.5 45.2 65.6 614 1155 49
3 Bias -2.15e-03 0.989 0.0573 0.037 0.0205
(SE 1) 2.49e-03 0.085 0.013 8.16e-03 2.51e-03
MSE 1.29e-04 1.12 6.64e-03 2.7e-03 5.42e-04
(SE 2) 2.81e-05 0.184 3.32e-03 9.48e-04 1.02e-04
Time (s) 9.24 296 46.4 118 2481 1533
4 Bias -2.77e-03 1.08 -2.74e-03 0.0177 0.021
(SE 1) 2.49e-03 0.087 2.28e-03 7.5e-03 2.67e-03
MSE 1.31e-04 1.32 1.12e-04 1.44e-03 5.78e-04
(SE 2) 2.7e-05 0.216 3.18e-05 7.42e-04 1.05e-04
Time (s) 29.6 808 643 523 3248 2590
5 Bias -2.51e-03 1.94 -2.65e-03 -4.62e-03 0.021
(SE 1) 2.78e-03 0.192 2.75e-03 2.49e-03 2.69e-03
MSE 1.61e-04 4.49 1.58e-04 1.46e-04 5.78e-04
(SE 2) 4.61e-05 0.996 3.25e-05 4.79e-05 1.08e-04
Time (s) 106 3300 3000 2821 5981 5010
Table 53: Bias and MSE for the GP range, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.5$ and a sample size of N=100’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias -2.32e-03 0.102 121878 0.879 0.559
(SE 1) 0.0462 0.0378 118271 0.287 0.0683
MSE 0.0428 0.039 294613704939 2.41 0.401
(SE 2) 0.0107 9.24e-03 287148615709 1.33 0.0653
Time (s) 1.72 17.5 1.91 22.5 309 367 13.3
2 Bias -5.99e-03 -0.0914 33.8 4.2 0.586
(SE 1) 0.0443 0.0312 31.4 1.31 0.0721
MSE 0.0393 0.0279 20856 52 0.442
(SE 2) 9.91e-03 8.16e-03 20323 32.6 0.0721
Time (s) 2.93 92.5 45.2 65.6 614 1155 49
3 Bias -0.0101 -0.158 2.85 1.64 0.57
(SE 1) 0.0423 0.0283 1.38 0.423 0.0694
MSE 0.036 0.0409 46.1 6.25 0.417
(SE 2) 8.32e-03 0.0104 41 2.72 0.0683
Time (s) 9.24 296 46.4 118 2481 1533
4 Bias -0.0126 -0.304 0.0321 1.1 0.616
(SE 1) 0.0487 0.023 0.0488 0.5 0.0851
MSE 0.0476 0.103 0.0487 6.21 0.517
(SE 2) 0.0126 0.0147 0.0157 5.05 0.11
Time (s) 29.6 808 643 523 3248 2590
5 Bias -0.0174 -0.508 -0.0125 0.0357 0.615
(SE 1) 0.0585 0.0158 0.0538 0.0526 0.0786
MSE 0.0687 0.263 0.058 0.0567 0.495
(SE 2) 0.0198 0.0163 0.0168 0.0154 0.0834
Time (s) 106 3300 3000 2821 5981 5010
Table 54: Bias and MSE for the GP marginal variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.5$ and a sample size of N=100’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias 1.56e-04 -0.342 0.159 -0.101 -2.3e-03 1.01e-03 3.95e-03
(SE 1) 5.24e-04 6.59e-03 0.0392 4.81e-03 1.2e-03 4.34e-04 4.86e-04
MSE 5.51e-06 0.118 0.0559 0.0106 3.25e-05 4.6e-06 2.01e-05
(SE 2) 1.31e-06 4.43e-03 0.016 9.44e-04 1.13e-05 1.17e-06 4e-06
Time (s) 1.72 17.5 1.91 22.5 309 367 13.3
2 Bias 4.75e-04 -0.261 -6.26e-04 -0.0374 -2.1e-03 8.96e-04 1.23e-03
(SE 1) 4.24e-04 6.07e-03 1.15e-03 1.08e-03 1.33e-03 4.23e-04 4.17e-04
MSE 3.83e-06 0.0688 2.66e-05 1.42e-03 3.81e-05 4.2e-06 4.82e-06
(SE 2) 9.76e-07 3.1e-03 6.41e-06 8.11e-05 9.53e-06 1.04e-06 1.27e-06
Time (s) 2.93 92.5 45.2 65.6 614 1155 49
3 Bias 6.16e-04 -0.233 -2.86e-03 -0.0308 6.81e-04 1.28e-03
(SE 1) 4.16e-04 5.25e-03 6.46e-04 8.27e-04 4.32e-04 4.19e-04
MSE 3.84e-06 0.0549 1.65e-05 9.62e-04 4.01e-06 4.98e-06
(SE 2) 9.9e-07 2.41e-03 3.35e-06 4.94e-05 1.26e-06 1.28e-06
Time (s) 9.24 296 46.4 118 2481 1533
4 Bias 5.9e-04 -0.175 -1.14e-03 -0.0133 7.26e-04 1.29e-03
(SE 1) 4.15e-04 4.03e-03 3.54e-04 6.08e-04 4.46e-04 4.18e-04
MSE 3.79e-06 0.0308 3.79e-06 1.84e-04 4.3e-06 5e-06
(SE 2) 9.01e-07 1.39e-03 8.84e-07 1.6e-05 1.24e-06 1.29e-06
Time (s) 29.6 808 643 523 3248 2590
5 Bias 6.89e-04 -0.102 3.47e-04 -2.49e-03 5.64e-04 1.3e-03
(SE 1) 4.42e-04 2.39e-03 4.01e-04 3.77e-04 4.46e-04 4.18e-04
MSE 4.39e-06 0.0105 3.34e-06 9.02e-06 4.09e-06 5e-06
(SE 2) 1.04e-06 4.82e-04 6.79e-07 1.83e-06 1.2e-06 1.29e-06
Time (s) 106 3300 3000 2821 5981 5010
Table 55: Bias and MSE for the error term variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.5$ and a sample size of N=100’000.
B.5.3 Prediction
Figure 32: Average RMSE and log-score on simulated datasets with an practical range of $0.5$ and a sample size N=100’000. Predictions are done using the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.22 0.488 0.463 0.384 0.476 0.1 0.0841
(SE 1) 5.89e-04 3.35e-03 7.02e-03 3.11e-03 8.16e-03 1.38e-03 2.14e-03
Log-score -0.0955 0.707 0.79 0.573 0.675 -0.904 -0.889
(SE 2) 2.68e-03 5.26e-03 0.0443 0.0162 0.0171 0.0153 0.0487
Time (s) 0.39 3.53 13.1 178 99.9 13.2 5.2
2 RMSE 0.159 0.434 0.156 0.227 0.276 0.0897 0.0749
(SE 1) 3.8e-04 1.78e-03 1.8e-03 2.56e-03 7.11e-03 1.48e-03 1.93e-03
Log-score -0.422 0.607 -0.371 0.0287 0.129 -1.01 -1.17
(SE 2) 2.4e-03 2.71e-03 0.0111 0.011 0.0254 0.018 0.0295
Time (s) 0.685 11.8 17.7 225 193 59.9 12.2
3 RMSE 0.116 0.416 0.13 0.19 0.0859 0.0742
(SE 1) 4.91e-04 1.38e-03 2.15e-03 1.59e-03 1.51e-03 1.93e-03
Log-score -0.738 0.571 -0.575 -0.15 -1.05 -1.18
(SE 2) 4.29e-03 2.1e-03 0.0147 9.73e-03 0.019 0.0294
Time (s) 1.47 30 14 268 146 255
4 RMSE 0.0893 0.374 0.0787 0.123 0.0838 0.0743
(SE 1) 8.79e-04 7.33e-04 2.28e-03 1.92e-03 1.58e-03 1.93e-03
Log-score -0.998 0.482 -1.11 -0.614 -1.08 -1.18
(SE 2) 0.0101 1.14e-03 0.0239 0.0134 0.0203 0.0294
Time (s) 4.19 84 81.1 432 173 449
5 RMSE 0.0794 0.312 0.0749 0.082 0.0824 0.0743
(SE 1) 1.32e-03 3.35e-04 1.82e-03 2.36e-03 1.63e-03 1.93e-03
Log-score -1.12 0.331 -1.17 -1.06 -1.09 -1.18
(SE 2) 0.0173 5.85e-04 0.0235 0.0233 0.021 0.0295
Time (s) 13.5 343 358 1164 288 872
Table 56: Average RMSE and log-score on the training set, standard errors and time needed for making predictions on the training set on simulated datasets with an effective range of $0.5$ and a sample size of N=100’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.22 0.623 0.61 0.435 0.647 0.101 0.0845
(SE 1) 5.17e-04 0.0177 0.019 7.04e-03 0.0207 1.47e-03 2.19e-03
Log-score -0.0947 1.02 1.06 0.739 0.979 -0.893 -0.884
(SE 2) 2.36e-03 0.0129 0.0515 0.0188 0.0315 0.0161 0.0494
Time (s) 0.448 3.57 2.69 176 97.9 41.4 5.42
2 RMSE 0.159 0.449 0.158 0.232 0.3 0.0907 0.0752
(SE 1) 4.1e-04 8.6e-03 2.14e-03 3.84e-03 9.33e-03 1.59e-03 1.98e-03
Log-score -0.421 0.818 -0.347 0.0819 0.213 -1 -1.16
(SE 2) 2.6e-03 5.48e-03 0.0115 0.012 0.0305 0.019 0.0301
Time (s) 0.753 12 5.77 223 189 78.9 12.6
3 RMSE 0.116 0.405 0.132 0.193 0.0868 0.0745
(SE 1) 5.29e-04 6.17e-03 2.53e-03 1.91e-03 1.6e-03 1.98e-03
Log-score -0.736 0.757 -0.556 -0.114 -1.04 -1.18
(SE 2) 4.62e-03 3.86e-03 0.0153 9.89e-03 0.0197 0.03
Time (s) 1.57 32 5.41 267 114 261
4 RMSE 0.0897 0.331 0.0791 0.123 0.0845 0.0745
(SE 1) 9.11e-04 2.32e-03 2.38e-03 2.19e-03 1.64e-03 1.98e-03
Log-score -0.994 0.621 -1.11 -0.596 -1.07 -1.18
(SE 2) 0.0104 1.51e-03 0.0241 0.0143 0.0206 0.03
Time (s) 4.21 84.8 48.4 432 141 458
5 RMSE 0.0797 0.267 0.0752 0.0825 0.0831 0.0745
(SE 1) 1.35e-03 3.93e-04 1.88e-03 2.49e-03 1.7e-03 1.98e-03
Log-score -1.11 0.422 -1.17 -1.06 -1.08 -1.18
(SE 2) 0.0175 3.16e-04 0.0239 0.0237 0.0215 0.0301
Time (s) 13.8 345 244 1163 215 890
Table 57: Average RMSE and log-score on the test “interpolation” set, standard errors and time needed for making predictions on the test “interpolation” set on simulated datasets with an practical range of $0.5$ and a sample size of N=100’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.835 0.994 0.936 0.926 2.64 0.84 0.821
(SE 1) 0.0574 0.0484 0.046 0.0492 0.965 0.0599 0.0641
Log-score 1.16 1.42 1.36 1.35 18.5 1.19 1.1
(SE 2) 0.0546 0.051 0.0473 0.0498 13.3 0.0524 0.0676
Time (s) 4.96 6.48 5.39 221 97.4 41.2 5.41
2 RMSE 0.832 0.992 0.847 0.862 17.4 0.832 0.819
(SE 1) 0.0578 0.0483 0.0552 0.0541 2.15 0.06 0.0628
Log-score 1.15 1.42 1.2 1.25 32.4 1.16 1.09
(SE 2) 0.055 0.0509 0.0546 0.0526 7.14 0.0534 0.0723
Time (s) 5.56 15.7 11.5 272 188 84.8 12.6
3 RMSE 0.829 0.992 0.839 0.852 0.838 0.816
(SE 1) 0.0581 0.0483 0.0555 0.0574 0.0597 0.0614
Log-score 1.14 1.42 1.18 1.23 1.16 1.09
(SE 2) 0.0552 0.0508 0.0552 0.0569 0.0525 0.0738
Time (s) 7.17 37.2 10.8 318 177 261
4 RMSE 0.824 0.989 0.83 0.842 0.835 0.816
(SE 1) 0.0584 0.0482 0.0606 0.0554 0.0602 0.0613
Log-score 1.12 1.42 1.13 1.19 1.15 1.09
(SE 2) 0.0545 0.0507 0.0632 0.0554 0.054 0.0743
Time (s) 12.6 95 96.9 503 200 458
5 RMSE 0.822 0.983 0.818 0.838 0.845 0.816
(SE 1) 0.0572 0.0481 0.0611 0.0576 0.0626 0.0613
Log-score 1.1 1.41 1.09 1.15 1.16 1.09
(SE 2) 0.0519 0.0504 0.0611 0.0588 0.0576 0.0741
Time (s) 31.7 419 488 1376 325 891
Table 58: Average RMSE and log-score on the test “extrapolation” set, standard errors and time needed for making predictions on the test “extrapolation” set on simulated datasets with an practical range of $0.5$ and a sample size of N=100’000.
B.6 Simulated data with $N=100^{\prime}000$ and practical range of $0.05$
Tuning parameter
Vecchia nb. neighbours 5 10 20 40 80
Tapering num. non-zero entries 8 17 22 40 111
FITC num. inducing points 20 200 275 900 2000
Full-scale num. inducing points 26 91 122 250 330
num. non-zero entries 6 11 15 27 73
FRK num. resolutions 1 2 exceeds time limit
MRA num. knots per partition 1 2 3 4 5
SPDE max edge convergence failure
Table 59: Tuning parameters chosen for the comparison on simulated datasets with an practical range of $0.05$ and a sample size of $N=100^{\prime}000$.
B.6.1 Log-likelihood evaluation
Figure 33: Average negative log-likelihood on simulated datasets with an practical range $0.05$ of and a sample size N=100’000. The true data-generating parameters are used on the left, and two times these values are used on the right.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 124646 147805 161627 149624 127735
(SE) 86.6 115 179 123 104
Time (s) 0.055 0.0876 0.0335 0.102 3.72
2 Value 122697 140389 158248 143769 125660
(SE) 90 85.6 207 100 96.8
Time (s) 0.0837 0.537 0.635 0.516 7.08
3 Value 122278 138056 156803 140551 124861
(SE) 89.2 77.7 157 80.1 97.2
Time (s) 0.164 1.09 0.77 0.936 14.4
4 Value 122166 133135 146604 134705 124417
(SE) 90.7 64.7 136 71.8 97.6
Time (s) 0.432 7.25 8.22 3.59 14.9
5 Value 122116 127566 136538 127313 124135
(SE) 90.8 58.5 134 62.2 99.2
Time (s) 1.33 54.6 47.1 30.1 25.8
Table 60: Mean approximate negative log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.05$ of and a sample size N=100’000. Log-likelihoods are evaluated at the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 132093 162080 171004 162823 134234
(SE) 53.5 57.9 86.2 67.6 76.9
Time (s) 0.058 0.0873 0.0347 0.103 3.81
2 Value 130000 156143 164368 157013 132627
(SE) 59.3 42.6 108 52.6 75.7
Time (s) 0.0836 0.538 0.635 0.517 7.56
3 Value 129752 154066 161916 153807 132015
(SE) 66.3 38.6 85.9 40.8 75.4
Time (s) 0.161 1.16 0.767 0.931 16.1
4 Value 129851 149191 148570 146510 131674
(SE) 70 32.1 104 40.2 76.3
Time (s) 0.424 7.25 8.21 3.6 17.1
5 Value 129844 142297 139021 136227 131453
(SE) 71.8 30.3 106 45.2 76.8
Time (s) 1.3 54.7 47.2 30.2 31.3
Table 61: Mean approximate negative log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.05$ of and a sample size N=100’000. Two times the true parameters are used.
B.6.2 Parameter estimation
Figure 34: Bias and MSE of the estimates for the GP range, GP variance end error term variance on simulated datasets with an practical range of $0.05$ and a sample size N=100’000,
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias -3.41e-05 0.354 1.19 3.69e-03 1.36e-03
(SE 1) 5.05e-05 0.0502 0.487 2.4e-04 9.1e-05
MSE 5.23e-08 0.176 6.16 1.48e-05 2.02e-06
(SE 2) 1.19e-08 0.0367 3.66 2.1e-06 2.93e-07
Time (s) 4.12 32 1.84 16.8 306 490
2 Bias -6.97e-05 0.529 0.0284 3.25e-03 8.24e-04
(SE 1) 5.09e-05 0.0593 4.05e-03 1.53e-04 7.22e-05
MSE 5.66e-08 0.35 1.14e-03 1.1e-05 7.78e-07
(SE 2) 1.61e-08 0.0499 3.9e-04 1.04e-06 1.33e-07
Time (s) 7.45 97.5 42 57 627 903
3 Bias -1.01e-04 0.818 0.0258 3.59e-03 7.28e-04
(SE 1) 5.25e-05 0.0635 2.64e-03 1.38e-04 9.84e-05
MSE 6.54e-08 0.75 8.03e-04 1.33e-05 7.14e-07
(SE 2) 1.56e-08 0.0963 1.43e-04 1.01e-06 1.82e-07
Time (s) 15.3 277 48.5 97.5 1831
4 Bias -7.39e-05 1.19 0.0137 4.04e-03 7.15e-04
(SE 1) 5.66e-05 0.0342 4.66e-04 1.12e-04 9.37e-05
MSE 6.95e-08 1.43 1.92e-04 1.65e-05 6.78e-07
(SE 2) 1.58e-08 0.0894 1.39e-05 9.09e-07 1.36e-07
Time (s) 42.4 871 588 427 2135
5 Bias -5.41e-05 1.41 9.61e-03 4.1e-03 7.04e-04
(SE 1) 6.08e-05 0.0959 2.45e-04 7.59e-05 9.59e-05
MSE 7.68e-08 2.18 9.36e-05 1.69e-05 6.71e-07
(SE 2) 1.63e-08 0.306 4.85e-06 6.24e-07 1.57e-07
Time (s) 130 3801 3012 3742 3549
Table 62: Bias and MSE for the GP range, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.05$ and a sample size of N=100’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias -2.22e-03 0.114 57.3 0.153 0.0533
(SE 1) 5.62e-03 4.99e-03 39.6 4.35e-03 0.0141
MSE 6.36e-04 0.0135 34703 0.0237 6.6e-03
(SE 2) 1.41e-04 1.24e-03 28480 1.4e-03 2.71e-03
Time (s) 4.12 32 1.84 16.8 306 490
2 Bias -5.19e-03 -0.0564 -0.462 0.0378 0.0613
(SE 1) 5.27e-03 4.41e-03 0.0716 4.2e-03 0.0175
MSE 5.82e-04 3.57e-03 0.316 1.78e-03 9.59e-03
(SE 2) 1.57e-04 4.21e-04 0.0317 3.89e-04 3.91e-03
Time (s) 7.45 97.5 42 57 627 903
3 Bias -8.59e-03 -0.111 -0.396 -0.0206 0.0767
(SE 1) 4.97e-03 2.94e-03 0.0294 3.65e-03 0.0172
MSE 5.67e-04 0.0125 0.174 6.91e-04 0.0115
(SE 2) 1.59e-04 6.08e-04 0.0213 1.48e-04 3.67e-03
Time (s) 15.3 277 48.5 97.5 1831
4 Bias -6.38e-03 -0.223 0.0759 -0.0961 0.0862
(SE 1) 5.44e-03 2.97e-03 0.0311 4.65e-03 0.0169
MSE 6.31e-04 0.05 0.0251 9.67e-03 0.0129
(SE 2) 1.55e-04 1.29e-03 0.0105 8.1e-04 3.36e-03
Time (s) 42.4 871 588 427 2135
5 Bias -2.65e-03 -0.305 0.521 -0.0627 0.0934
(SE 1) 6.35e-03 3.13e-03 0.0289 6.29e-03 0.0181
MSE 8.14e-04 0.0931 0.288 4.72e-03 0.0149
(SE 2) 2.82e-04 1.89e-03 0.0327 7.93e-04 3.82e-03
Time (s) 130 3801 3012 3742 3549
Table 63: Bias and MSE for the GP marginal variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.05$ and a sample size of N=100’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias 2.93e-04 -0.342 0.861 -0.343 -0.0763 9.16e-03
(SE 1) 6.05e-04 1.46e-03 0.0344 1.14e-03 1.36e-03 9.08e-04
MSE 7.4e-06 0.117 0.764 0.118 5.85e-03 9.95e-05
(SE 2) 1.97e-06 1.01e-03 0.0508 7.83e-04 2.04e-04 1.66e-05
Time (s) 4.12 32 1.84 16.8 306 490
2 Bias 3.21e-04 -0.263 0.646 -0.277 -0.0765 4.92e-03
(SE 1) 6.75e-04 1.14e-03 0.0179 8.58e-04 1.39e-03 8.5e-04
MSE 9.21e-06 0.0691 0.424 0.0767 5.89e-03 3.8e-05
(SE 2) 2.76e-06 6.06e-04 0.0227 4.76e-04 2.14e-04 8.16e-06
Time (s) 7.45 97.5 42 57 627 903
3 Bias 2.54e-04 -0.236 0.596 -0.244 3.23e-03
(SE 1) 7.25e-04 8.48e-04 0.0177 6.29e-04 8.36e-04
MSE 1.06e-05 0.0557 0.362 0.0594 2.37e-05
(SE 2) 3.6e-06 4.03e-04 0.02 3.05e-04 6.03e-06
Time (s) 15.3 277 48.5 97.5 1831
4 Bias 4.74e-04 -0.178 0.307 -0.179 2.94e-03
(SE 1) 7.24e-04 4.72e-04 3.68e-03 7.05e-04 9.33e-04
MSE 1.07e-05 0.0315 0.0948 0.0321 2.52e-05
(SE 2) 3.42e-06 1.68e-04 2.24e-03 2.52e-04 6.6e-06
Time (s) 42.4 871 588 427 2135
5 Bias 4.2e-04 -0.108 0.13 -0.0991 3.13e-03
(SE 1) 7.45e-04 5.11e-04 2.4e-03 6.11e-04 8.76e-04
MSE 1.13e-05 0.0116 0.0171 9.84e-03 2.43e-05
(SE 2) 3.6e-06 1.1e-04 6.09e-04 1.22e-04 5.17e-06
Time (s) 130 3801 3012 3742 3549
Table 64: Bias and MSE for the error term variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.05$ and a sample size of N=100’000.
B.6.3 Prediction
Figure 35: Average RMSE and log-score on simulated datasets with an practical range of $0.05$ and a sample size N=100’000. Predictions are done using the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.291 0.488 0.575 0.501 0.578 0.335
(SE 1) 5.14e-04 4.87e-04 6.78e-04 5.11e-04 9.29e-04 7.12e-04
Log-score 0.179 0.709 0.867 0.731 0.874 0.307
(SE 2) 1.81e-03 8.23e-04 1.25e-03 9.08e-04 1.78e-03 2.21e-03
Time (s) 0.409 5.18 13.2 179 98.8 4.54
2 RMSE 0.276 0.441 0.564 0.465 0.572 0.312
(SE 1) 6.32e-04 4.06e-04 8.08e-04 4.75e-04 9e-04 7.1e-04
Log-score 0.124 0.62 0.852 0.665 0.862 0.238
(SE 2) 2.41e-03 6.97e-04 1.65e-03 8.82e-04 1.74e-03 2.48e-03
Time (s) 0.743 11.7 17.7 224 202 9.5
3 RMSE 0.274 0.426 0.559 0.444 0.303
(SE 1) 6.51e-04 3.82e-04 6.78e-04 4.04e-04 6.72e-04
Log-score 0.116 0.59 0.844 0.624 0.209
(SE 2) 2.47e-03 6.62e-04 1.35e-03 7.59e-04 2.43e-03
Time (s) 1.56 32.7 13.9 263 18.7
4 RMSE 0.273 0.391 0.516 0.404 0.298
(SE 1) 6.78e-04 3.33e-04 6.69e-04 3.47e-04 6.69e-04
Log-score 0.11 0.516 0.773 0.541 0.193
(SE 2) 2.58e-03 5.9e-04 1.5e-03 6.91e-04 2.51e-03
Time (s) 4.53 86.3 81 420 20.7
5 RMSE 0.272 0.345 0.455 0.342 0.295
(SE 1) 6.77e-04 2.58e-04 7.15e-04 2.64e-04 6.45e-04
Log-score 0.11 0.402 0.641 0.385 0.184
(SE 2) 2.58e-03 4.96e-04 1.78e-03 5.56e-04 2.42e-03
Time (s) 14.6 368 357 1141 36.2
Table 65: Average RMSE and log-score on the training set, standard errors and time needed for making predictions on the training set on simulated datasets with an effective range of $0.05$ and a sample size of N=100’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.316 0.629 0.994 0.676 0.99 0.383
(SE 1) 8.27e-04 1.82e-03 2.7e-03 1.87e-03 3.58e-03 1.17e-03
Log-score 0.255 1.02 1.41 1.07 1.41 0.426
(SE 2) 2.64e-03 1.34e-03 2.7e-03 1.51e-03 3.61e-03 2.54e-03
Time (s) 0.472 5.24 2.71 176 98.7 39.4
2 RMSE 0.301 0.477 0.945 0.551 0.955 0.351
(SE 1) 9.27e-04 1.09e-03 3.71e-03 1.52e-03 2.91e-03 1.15e-03
Log-score 0.206 0.846 1.36 0.931 1.37 0.344
(SE 2) 3.2e-03 7.16e-04 4.02e-03 1.18e-03 3.11e-03 3.15e-03
Time (s) 0.814 11.9 5.75 223 198 57.2
3 RMSE 0.298 0.439 0.924 0.491 0.34
(SE 1) 9.39e-04 8.97e-04 2.93e-03 1.19e-03 1.08e-03
Log-score 0.197 0.793 1.34 0.856 0.311
(SE 2) 3.25e-03 5.94e-04 2.64e-03 1.03e-03 3.21e-03
Time (s) 1.64 31.7 5.4 262 71.4
4 RMSE 0.296 0.377 0.765 0.405 0.333
(SE 1) 9.8e-04 5.97e-04 2.45e-03 9.44e-04 1.12e-03
Log-score 0.191 0.678 1.16 0.72 0.292
(SE 2) 3.41e-03 4.53e-04 3.13e-03 7.29e-04 3.35e-03
Time (s) 4.67 87.1 48.4 417 77
5 RMSE 0.296 0.333 0.605 0.336 0.328
(SE 1) 9.74e-04 5.97e-04 2.3e-03 6.33e-04 1.14e-03
Log-score 0.19 0.524 0.925 0.509 0.28
(SE 2) 3.39e-03 5.91e-04 3.21e-03 7.77e-04 3.44e-03
Time (s) 14.6 370 244 1141 110
Table 66: Average RMSE and log-score on the test “interpolation” set, standard errors and time needed for making predictions on the test “interpolation” set on simulated datasets with an practical range of $0.05$ and a sample size of N=100’000.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.977 0.99 0.991 0.99 1.06 0.982
(SE 1) 6.35e-03 7.1e-03 7.12e-03 7.11e-03 0.0304 7.16e-03
Log-score 1.39 1.41 1.41 1.41 1.48 1.4
(SE 2) 6.45e-03 6.97e-03 6.99e-03 6.98e-03 0.0321 7.27e-03
Time (s) 5.06 9.56 5.55 221 97.1 35.5
2 RMSE 0.977 0.989 0.991 0.99 19.9 0.981
(SE 1) 6.23e-03 7.07e-03 7.13e-03 7.09e-03 4.66 7.13e-03
Log-score 1.39 1.41 1.41 1.41 15.5 1.4
(SE 2) 6.27e-03 6.93e-03 7e-03 6.96e-03 3.46 7.23e-03
Time (s) 5.67 15.6 11.5 271 198 50.8
3 RMSE 0.976 0.988 0.991 0.989 0.98
(SE 1) 6.16e-03 7.05e-03 7.13e-03 7.07e-03 7.15e-03
Log-score 1.39 1.41 1.41 1.41 1.4
(SE 2) 6.16e-03 6.91e-03 7e-03 6.94e-03 7.27e-03
Time (s) 7.28 36.2 10.8 314 68.5
4 RMSE 0.976 0.986 0.989 0.988 0.98
(SE 1) 6.16e-03 7.01e-03 7.15e-03 7.04e-03 7.07e-03
Log-score 1.39 1.4 1.41 1.41 1.39
(SE 2) 6.17e-03 6.86e-03 7.03e-03 6.9e-03 7.19e-03
Time (s) 13.3 101 96.8 487 73.1
5 RMSE 0.976 0.983 0.985 0.984 0.98
(SE 1) 6.17e-03 6.9e-03 7e-03 7e-03 7.02e-03
Log-score 1.39 1.4 1.4 1.4 1.39
(SE 2) 6.18e-03 6.75e-03 6.89e-03 6.85e-03 7.13e-03
Time (s) 33.4 450 487 1352 100
Table 67: Average RMSE and log-score on the test “extrapolation” set, standard errors and time needed for making predictions on the test “extrapolation” set on simulated datasets with an practical range of $0.05$ and a sample size of N=100’000.
B.7 House price dataset
B.7.1 Prediction
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.325 0.382 0.383 0.345 0.544 0.337 0.41
Log-score 0.262 0.387 0.395 0.29 0.81 0.292 0.524
CRPS 0.17 0.196 0.2 0.18 0.298 0.175 0.219
Time (s) 0.471 6.97 7.51 7.56 68.8 89.7 7.99
2 RMSE 0.324 0.345 0.355 0.331 0.451 0.326 0.347
Log-score 0.257 0.292 0.329 0.256 0.622 0.266 0.347
CRPS 0.169 0.178 0.186 0.171 0.243 0.171 0.184
Time (s) 1.13 30 27.2 27.6 155 174 29.9
3 RMSE 0.324 0.332 0.346 0.323 0.397 0.325 0.336
Log-score 0.258 0.266 0.29 0.243 0.492 0.261 0.304
CRPS 0.169 0.172 0.179 0.168 0.213 0.17 0.176
Time (s) 1.76 92.2 94 84.7 985 590 88.6
4 RMSE 0.323 0.324 0.334 0.324 0.325 0.326
Log-score 0.257 0.252 0.272 0.243 0.261 0.266
CRPS 0.169 0.168 0.173 0.168 0.17 0.17
Time (s) 4.35 528 508 560 871 497
5 RMSE 0.323 0.322 0.327 0.323 0.324 0.324
Log-score 0.257 0.25 0.26 0.25 0.259 0.258
CRPS 0.169 0.168 0.17 0.168 0.17 0.169
Time (s) 14 1599 1467 1347 1576 1312
Table 68: Test “interpolation” RMSE, log-score, CRPS and time needed for making predictions on the “interpolation” test set on the house price dataset. Every approximation makes predictions using its parameter estimates.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.597 0.679 0.559 0.565 1.45 0.581 0.708
Log-score 0.929 1.05 0.971 0.921 3.75 0.952 1.24
CRPS 0.336 0.378 0.332 0.326 0.976 0.335 0.411
Time (s) 0.498 6.82 7.51 6.97 69 89.7 7.99
2 RMSE 0.596 0.658 0.574 0.571 10.5 0.602 0.579
Log-score 0.925 1.02 0.955 0.943 23.7 0.948 0.921
CRPS 0.335 0.366 0.333 0.33 6.35 0.341 0.33
Time (s) 1.21 29.6 27.2 26.8 154 175 29.9
3 RMSE 0.595 0.641 0.583 0.584 2.76 0.598 0.582
Log-score 0.933 0.982 0.965 0.948 2.59 0.941 0.919
CRPS 0.336 0.355 0.337 0.335 1.5 0.338 0.331
Time (s) 1.74 91.2 94.2 83 979 584 88.6
4 RMSE 0.597 0.617 0.595 0.592 0.598 0.591
Log-score 0.934 0.931 0.942 0.952 0.941 0.921
CRPS 0.337 0.342 0.337 0.338 0.338 0.334
Time (s) 4.42 523 508 552 861 497
5 RMSE 0.597 0.608 0.601 0.595 0.595 0.593
Log-score 0.933 0.916 0.941 0.939 0.935 0.92
CRPS 0.337 0.338 0.34 0.337 0.336 0.334
Time (s) 14.1 1587 1467 1334 1563 1312
Table 69: Test “extrapolation” RMSE, log-score, CRPS and time needed for making predictions on the “extrapolation” test set on the house price dataset. Every approximation makes predictions using its parameter estimates.
B.7.2 Likelihood evaluation
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 6184 12120 9047 9230 6750
Time (s) 0.0132 0.057 0.0969 0.0806 0.451
2 Value 5862 9114 6893 6770 6063
Time (s) 0.0233 0.29 0.42 0.268 2.39
3 Value 5751 7847 6055 5783 5853
Time (s) 0.0351 0.897 1.25 0.732 8.6
4 Value 5721 6732 5753 5548 5822
Time (s) 0.0731 5.97 9.77 5.31 14.5
5 Value 5714 6446 5731 5633 5802
Time (s) 0.226 26.7 25.9 13.6 26.9
Table 70: Approximate negative log-likelihood and time needed for its evaluation on the house price dataset. The covariance parameters estimated via exact calculations are used.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 7587 15704 9027 9720 7773
Time (s) 9.12e-03 0.0565 0.0989 0.0745 0.467
2 Value 7142 12729 7642 7787 7213
Time (s) 0.0232 0.288 0.481 0.259 3.59
3 Value 6992 11229 7160 7150 7085
Time (s) 0.0279 0.898 1.34 0.708 11.5
4 Value 6962 9649 6984 6941 7049
Time (s) 0.0745 5.96 9.79 6.17 16.6
5 Value 6955 9166 6960 6944 7031
Time (s) 0.23 26.8 27.8 13.7 30
Table 71: Approximate negative log-likelihood and time needed for its evaluation on the house price dataset. Two times the values of the covariance parameters estimated via exact calculations are used.
B.7.3 Comparison to exact calculations
Figure 36: RMSE between exact and approximate predictive means and variances on the house price dataset. Every approximation makes predictions using its parameter estimates.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE means 0.0571 0.223 0.229 0.165 0.453 0.117 0.278
RMSE vars 3.78e-03 0.0793 0.0921 0.0925 0.212 0.0276 0.0804
KL-divergence 65 769 739 460 2224 211 1162
Time (s) 0.471 6.97 7.51 7.56 68.8 89.7 7.99
2 RMSE means 0.033 0.137 0.169 0.115 0.336 0.0598 0.151
RMSE vars 1.69e-03 0.0314 0.0657 0.0618 0.114 0.0108 0.0286
KL-divergence 20.9 340 431 230 1524 62.4 421
Time (s) 1.13 30 27.2 27.6 155 174 29.9
3 RMSE means 0.0203 0.0976 0.13 0.0818 0.258 0.0455 0.0901
RMSE vars 6.33e-04 0.0186 0.0537 0.05 0.0686 5.88e-03 0.0143
KL-divergence 8.02 188 254 119 1041 36.4 161
Time (s) 1.76 92.2 94 84.7 985 590 88.6
4 RMSE means 8.84e-03 0.0608 0.068 0.0406 0.0421 0.031
RMSE vars 2.77e-04 0.0104 0.0245 0.0286 4.55e-03 4.69e-03
KL-divergence 1.57 81.3 72.8 30.3 32 20.4
Time (s) 4.35 528 508 560 871 497
5 RMSE means 3.29e-03 0.053 0.0425 0.0232 0.034 0.0146
RMSE vars 4.67e-05 9.55e-03 0.0124 0.0133 3.85e-03 2.35e-03
KL-divergence 0.205 63.3 25.7 9.22 20.2 4.66
Time (s) 14 1599 1467 1347 1576 1312
Table 72: RMSE between means, RMSE between variances, KL divergence between exact and approximate predictive distributions and time needed for making predictions on the “interpolation” test set of the house price dataset.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE means 0.0553 0.32 0.244 0.201 1.31 0.105 0.333
RMSE vars 0.0172 0.191 0.246 0.102 0.245 0.0912 0.17
KL-divergence 22.2 752 369 245 11057 79.5 1036
Time (s) 0.498 6.82 7.51 6.97 69 89.7 7.99
2 RMSE means 0.038 0.27 0.111 0.183 10.5 0.082 0.14
RMSE vars 0.0274 0.214 0.164 0.12 3.84 0.0432 0.0757
KL-divergence 11.7 702 129 197 89167 35.1 162
Time (s) 1.21 29.6 27.2 26.8 154 175 29.9
3 RMSE means 0.0199 0.208 0.126 0.113 2.57 0.0522 0.109
RMSE vars 6.98e-03 0.21 0.16 0.11 7.41 0.0163 0.0408
KL-divergence 2.38 519 142 97.6 6900 16.8 75.9
Time (s) 1.74 91.2 94.2 83 979 584 88.6
4 RMSE means 7.85e-03 0.099 0.0788 0.106 0.0514 0.03
RMSE vars 5.11e-03 0.172 0.0511 0.0994 0.0163 0.0248
KL-divergence 0.458 210 42.2 75 15.8 9.35
Time (s) 4.42 523 508 552 861 497
5 RMSE means 4.88e-03 0.0609 0.0342 0.104 0.0426 0.0182
RMSE vars 7.05e-04 0.141 0.0131 0.0465 0.012 0.0267
KL-divergence 0.114 115 8.11 53.4 11.3 4.5
Time (s) 14.1 1587 1467 1334 1563 1312
Table 73: RMSE between means, RMSE between variances, KL divergence between exact and approximate predictive distributions and time needed for making predictions on the “extrapolation” test set of the house price dataset.
B.8 Laegern dataset
B.8.1 Prediction
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.0516 0.132 0.154 0.168 0.186 0.0909 0.1
Log-score -1.56 -0.581 -0.46 -0.382 -0.263 -1.01 -0.883
CRPS 0.027 0.0745 0.0859 0.0937 0.106 0.0475 0.054
Time (s) 7.94 247 211 229 215 1243 217
2 RMSE 0.0494 0.0932 0.145 0.116 0.177 0.0805 0.0843
Log-score -1.6 -0.888 -0.527 -0.702 -0.311 -1.13 -1.05
CRPS 0.0257 0.0527 0.08 0.0645 0.1 0.042 0.0451
Time (s) 9.71 713 775 779 434 2829 745
3 RMSE 0.0486 0.079 0.138 0.0844 0.16 0.0755 0.0713
Log-score -1.62 -1.04 -0.575 -0.972 -0.417 -1.2 -1.22
CRPS 0.0253 0.0446 0.0758 0.0474 0.0896 0.0392 0.0377
Time (s) 78.9 1619 1904 1856 689 4588 1783
4 RMSE 0.0486 0.0678 0.135 0.0736 0.0726 0.0634
Log-score -1.62 -1.21 -0.605 -1.1 -1.24 -1.33
CRPS 0.0253 0.0379 0.0735 0.0413 0.0376 0.0334
Time (s) 184 3121 3425 3122 5130 3422
5 RMSE 0.0486 0.0598 0.127 0.064 0.0681 0.0573
Log-score -1.62 -1.37 -0.665 -1.24 -1.31 -1.44
CRPS 0.0252 0.0325 0.0688 0.0356 0.0352 0.0301
Time (s) 764 8940 9555 8017 8058 8212
Table 74: Test “interpolation” RMSE, log-score, CRPS and time needed for making predictions on the “interpolation” test set on the Laegern dataset. Every approximation makes predictions using its parameter estimates.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.191 0.194 0.199 0.209 0.23 0.192 0.198
Log-score -0.198 -0.197 0.139 0.206 -2.21e-03 -0.229 -0.101
CRPS 0.11 0.111 0.134 0.143 0.134 0.109 0.116
Time (s) 9.31 246 211 229 214 1244 217
2 RMSE 0.192 0.193 0.201 0.193 1.45 0.192 0.192
Log-score -0.229 -0.128 0.356 -0.161 13.4 -0.227 -0.154
CRPS 0.109 0.112 0.158 0.112 0.752 0.109 0.112
Time (s) 11.1 707 775 708 434 2827 745
3 RMSE 0.192 0.193 0.207 0.193 2247 0.192 0.191
Log-score -0.231 -0.0753 0.804 -0.196 11.4 -0.229 -0.196
CRPS 0.109 0.113 0.242 0.11 704 0.109 0.109
Time (s) 80.5 1603 1904 1755 690 4588 1783
4 RMSE 0.192 0.193 0.258 0.193 0.192 0.191
Log-score -0.232 -3.3e-04 1.19 -0.208 -0.229 -0.208
CRPS 0.109 0.114 0.375 0.11 0.109 0.109
Time (s) 186 3090 3425 2986 5129 3422
5 RMSE 0.192 0.193 0.418 0.192 0.192 0.191
Log-score -0.232 -4.48e-03 1.52 -0.232 -0.229 -0.218
CRPS 0.109 0.113 0.57 0.109 0.109 0.109
Time (s) 767 8853 9561 7758 8059 8212
Table 75: Test “extrapolation” RMSE, log-score, CRPS and time needed for making predictions on the “extrapolation” test set on the Laegern dataset. Every approximation makes predictions using its parameter estimates.
B.8.2 Log-likelihood evaluation
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value -221776 -63605 -44870 -40836 -159974
Time (s) 0.0968 0.29 4.38 1.45 7.79
2 Value -232645 -83954 -51950 -69641 -179025
Time (s) 0.158 0.903 13.8 6.06 14.1
3 Value -234768 -95433 -57332 -92766 -189639
Time (s) 0.307 2.21 26.2 16.3 21.1
4 Value -235276 -111701 -62877 -103885 -195678
Time (s) 0.824 5.19 42.4 21.7 28.7
5 Value -235332 -144967 -74816 -126180 -203352
Time (s) 2.59 21.6 122 40.6 45.2
Table 76: Approximate negative log-likelihood and time needed for its evaluation on the Laegern dataset. The covariance parameters estimated by a Vecchia approximation with 240 neighbors are used.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value -156187 -28386 -29081 -19570 -127829
Time (s) 0.0925 0.291 2.69 0.965 7.84
2 Value -138867 -44428 -42569 -45256 -125582
Time (s) 0.157 0.909 8.12 4.69 14.5
3 Value -131620 -54872 -53401 -70953 -125547
Time (s) 0.304 2.21 15.5 12.1 22.4
4 Value -129886 -70668 -62599 -82656 -125221
Time (s) 0.812 5.18 24.6 16.4 30.3
5 Value -129805 -107711 -81065 -109549 -125789
Time (s) 2.58 21.6 112 32.2 49.4
Table 77: Approximate negative log-likelihood and time needed for its evaluation on the Laegern dataset. Two times the covariance parameters estimated by a Vecchia approximation with the largest considered number of neighbors are used.
B.9 MODIS 2016 dataset
B.9.1 Prediction
Vecchia Tapering FITC Full-scale FRK MRA SPDE npspec
1 RMSE 1.61 2.98 2.37 2.36 2.86 2.34 2.54 1.78
Log-score 1.79 3.01 2.37 2.31 2.75 2.21 2.57 3.41
CRPS 0.852 1.87 1.39 1.37 1.71 1.27 1.48 0.907
Time (s) 10.8 103 109 113 151 570 99.1 674
2 RMSE 1.5 2.87 2.45 2.35 2.78 1.87 2.6
Log-score 1.75 3.38 2.38 2.41 2.7 1.93 2.63
CRPS 0.788 1.84 1.41 1.38 1.66 0.97 1.52
Time (s) 22.5 459 453 503 380 967 146
3 RMSE 1.6 2.73 2.5 2.25 2.48 2.26
Log-score 1.76 3.67 2.37 2.3 2.5 2.24
CRPS 0.83 1.75 1.42 1.29 1.43 1.24
Time (s) 58.2 1237 1453 1396 2671 1797
4 RMSE 1.68 2.55 2.39 2.37 1.86
Log-score 1.78 3.29 2.33 2.39 1.87
CRPS 0.861 1.57 1.36 1.35 0.94
Time (s) 153 3609 3256 2555 3178
5 RMSE 1.67 2.45 2.31 2.29 1.69
Log-score 1.77 2.91 2.28 2.28 1.81
CRPS 0.851 1.46 1.31 1.28 0.871
Time (s) 452 5609 6422 5800 6602
Table 78: Test RMSE, log-score, CRPS and time needed for making predictions on the test set on the MODIS 2016 dataset. Every approximation makes predictions using its parameter estimates.
B.9.2 Log-likelihood evaluation
Vecchia Tapering FITC Full-scale FRK MRA SPDE npspec
1 Value 122663 216534 198554 203127 21483784
Time (s) 0.0513 0.217 1.32 0.836 3.31
2 Value 119805 201886 188378 178753 9433108
Time (s) 0.0836 1.16 4.58 2.55 9.42
3 Value 119144 185129 180404 168911 7585086
Time (s) 0.164 4.05 13.5 4.98 16.1
4 Value 119158 168029 174236 154500 6591441
Time (s) 0.419 12.2 36.8 9.61 26.9
5 Value 119144 159715 167247 142226 5625433
Time (s) 1.45 33.3 94 20.8 65.9
Table 79: Approximate negative log-likelihood and time needed for its evaluation on the MODIS 2016 dataset. The covariance parameters estimated by a Vecchia approximation with the largest considered number of neighbors are used.
Vecchia Tapering FITC Full-scale FRK MRA SPDE npspec
1 Value 122754 239799 200344 205824 3307953
Time (s) 0.0542 0.217 1.36 0.867 3.32
2 Value 119795 228512 189025 181639 1399985
Time (s) 0.0794 1.16 4.47 2.47 9.25
3 Value 119162 213135 180642 170865 1136081
Time (s) 0.166 4.05 13.5 5.07 16
4 Value 119188 195056 174364 155906 1011310
Time (s) 0.426 12.1 36.9 10.1 26.9
5 Value 119174 185366 167307 143126 886374
Time (s) 1.44 33.3 95.1 21 65.9
Table 80: Approximate negative log-likelihood and time needed for its evaluation on the MODIS 2016 dataset. Two times the parameters estimated by a Vecchia approximation with the largest considered number of neighbors are used.
B.10 MODIS 2023 dataset
B.10.1 Prediction
Vecchia Tapering FITC Full-scale FRK MRA SPDE npspec
1 RMSE 1.91 2.55 1.75 1.97 1.75 1.94 2.1 1.52
Log-score 1.88 2.38 2.03 2.19 2.01 1.91 2.04 2.97
CRPS 0.979 1.46 0.984 1.14 0.971 1 1.11 0.777
Time (s) 36.6 541 562 628 459 9482 459 13222
2 RMSE 1.91 2.47 2.07 1.89 4.04 2.16
Log-score 1.87 2.31 2.04 2.12 3.51 2.07
CRPS 0.98 1.39 1.04 1.08 1.59 1.15
Time (s) 70.3 857 835 918 1607 901
3 RMSE 1.92 2.45 2.81 1.97 2.2
Log-score 1.87 2.31 2 2.08 2.09
CRPS 0.983 1.38 1.1 1.06 1.18
Time (s) 163 1455 1621 1520 1348
4 RMSE 1.91 2.43 2.38 1.58 2.26
Log-score 1.87 2.31 1.99 1.88 2.14
CRPS 0.983 1.37 1.07 0.856 1.22
Time (s) 527 3378 3525 2979 3069
5 RMSE 1.91 2.31 1.47 1.43 1.87
Log-score 1.87 2.37 1.87 1.77 1.85
CRPS 0.982 1.31 0.83 0.767 0.955
Time (s) 1393 13352 12366 8768 10990
Table 81: Test RMSE, log-score, CRPS and time needed for making predictions on the test set on the MODIS 2023 dataset. Every approximation makes predictions using its parameter estimates.
B.10.2 Log-likelihood evaluation
Vecchia Tapering FITC Full-scale FRK MRA SPDE npspec
1 Value 439151 1043380 1076009 1117564 13825680
Time (s) 0.21 0.112 5 0.0511 18.8
2 Value 430413 901574 1064069 1095756
Time (s) 0.336 0.534 6.92 2.52
3 Value 429066 878339 1031097 1007950
Time (s) 0.673 0.942 16.5 6.3
4 Value 429378 847272 963170 860055
Time (s) 1.71 2.87 43.7 16.6
5 Value 429420 726820 868875 762724
Time (s) 5.47 23.2 228 49.3
Table 82: Approximate negative log-likelihood and time needed for its evaluation on the MODIS 2023 dataset. The covariance parameters estimated by a Vecchia approximation with 240 neighbors are used.
Vecchia Tapering FITC Full-scale FRK MRA SPDE npspec
1 Value 441142 1016675 971274 1055451 2739914
Time (s) 0.21 0.105 4.29 0.0517 20.2
2 Value 440476 930184 950278 1010357
Time (s) 0.332 0.53 6.08 2.28
3 Value 442933 914694 903253 941800
Time (s) 0.677 0.935 14.1 5.96
4 Value 445301 893140 841040 841423
Time (s) 1.7 2.85 37.5 15.7
5 Value 445698 800784 778644 760559
Time (s) 5.37 23.2 212 47
Table 83: Approximate negative log-likelihood and time needed for its evaluation on the MODIS 2023 dataset. Two times the parameters estimated by a Vecchia approximation with the largest considered number of neighbors are used.
B.11 Simulated data with $N=100^{\prime}000$ and practical range of $0.05$, single-core setting
Tuning parameter
Vecchia nb. neighbours 5 10 20 40 80
Tapering num. non-zero entries 8 17 22 40 111
FITC num. inducing points 20 200 275 900 2000
Full-scale num. inducing points 26 91 122 250 330
num. non-zero entries 6 11 15 27 73
FRK num. resolutions 1 2 exceeds time limit
MRA num. knots per partition 1 2 3 4 5
SPDE max edge convergence failure
Table 84: Tuning parameters chosen for the comparison on simulated datasets with an practical range of $0.05$ and a sample size of $N=100^{\prime}000$ in the single-core setting.
B.11.1 Log-likelihood evaluation
Figure 37: Average negative log-likelihood on simulated datasets with an practical range $0.05$ of and a sample size N=100’000 in the single-core setting. The true data-generating parameters are used on the left, and two times these values are used on the right.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 124646 147805 161627 149624 127735
(SE) 86.6 115 179 123 104
Time (s) 0.239 0.0896 0.068 0.202 4.85
2 Value 122697 140389 158248 143769 125660
(SE) 90 85.6 207 100 96.8
Time (s) 0.38 0.605 2.11 1.13 9.03
3 Value 122278 138056 156803 140551 124861
(SE) 89.2 77.7 157 80.1 97.2
Time (s) 0.815 1.2 3.79 2.13 13.9
4 Value 122166 133135 146604 134705 124417
(SE) 90.7 64.7 136 71.8 97.6
Time (s) 2.48 6.52 37.2 9.02 18.6
5 Value 122116 127566 136538 127313 124135
(SE) 90.8 58.5 134 62.2 99.2
Time (s) 8.6 46.5 219 76.2 24.7
Table 85: Mean approximate negative log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.05$ of and a sample size N=100’000 in the single-core setting. Log-likelihoods are evaluated at the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 132093 162080 171004 162823 134234
(SE) 53.5 57.9 86.2 67.6 76.9
Time (s) 0.24 0.0898 0.0685 0.201 4.96
2 Value 130000 156143 164368 157013 132627
(SE) 59.3 42.6 108 52.6 75.7
Time (s) 0.379 0.604 2.11 1.12 9.48
3 Value 129752 154066 161916 153807 132015
(SE) 66.3 38.6 85.9 40.8 75.4
Time (s) 0.815 1.2 3.79 2.13 15.4
4 Value 129851 149191 148570 146510 131674
(SE) 70 32.1 104 40.2 76.3
Time (s) 2.48 6.51 37.2 8.98 21
5 Value 129844 142297 139021 136227 131453
(SE) 71.8 30.3 106 45.2 76.8
Time (s) 8.59 46.5 219 76.3 29.3
Table 86: Mean approximate negative log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.05$ of and a sample size N=100’000 in the single-core setting. Two times the true parameters are used.
B.11.2 Parameter estimation
Figure 38: Bias and MSE of the estimates for the GP range, GP variance end error term variance on simulated datasets with an practical range of $0.05$ and a sample size N=100’000 in the single-core setting,
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias -3.41e-05 0.354 1.2 3.69e-03 1.36e-03
(SE 1) 5.05e-05 0.0502 0.494 2.4e-04 9.1e-05
MSE 5.23e-08 0.176 6.33 1.48e-05 2.02e-06
(SE 2) 1.19e-08 0.0367 3.8 2.1e-06 2.93e-07
Time (s) 11.1 158 2.22 109 668 624
2 Bias -6.97e-05 0.529 0.0284 3.25e-03 8.24e-04
(SE 1) 5.09e-05 0.0593 4.05e-03 1.53e-04 7.22e-05
MSE 5.66e-08 0.35 1.14e-03 1.1e-05 7.78e-07
(SE 2) 1.61e-08 0.0499 3.9e-04 1.04e-06 1.33e-07
Time (s) 17.6 608 71.8 272 1148 1102
3 Bias -1.01e-04 0.818 0.0258 3.59e-03 7.28e-04
(SE 1) 5.25e-05 0.0635 2.64e-03 1.38e-04 9.84e-05
MSE 6.54e-08 0.75 8.03e-04 1.33e-05 7.14e-07
(SE 2) 1.56e-08 0.0963 1.43e-04 1.01e-06 1.82e-07
Time (s) 36.5 1030 127 450 1735
4 Bias -7.39e-05 1.19 0.0137 4.04e-03 7.15e-04
(SE 1) 5.66e-05 0.0342 4.66e-04 1.12e-04 9.37e-05
MSE 6.95e-08 1.43 1.92e-04 1.65e-05 6.78e-07
(SE 2) 1.58e-08 0.0894 1.39e-05 9.09e-07 1.36e-07
Time (s) 120 3033 1131 1358 2570
5 Bias -5.41e-05 1.41 9.61e-03 4.1e-03 7.04e-04
(SE 1) 6.08e-05 0.0959 2.45e-04 7.59e-05 9.59e-05
MSE 7.68e-08 2.18 9.36e-05 1.69e-05 6.71e-07
(SE 2) 1.63e-08 0.306 4.85e-06 6.24e-07 1.57e-07
Time (s) 498 11774 6145 9601 3347
Table 87: Bias and MSE for the GP range, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.05$ and a sample size of N=100’000 in the single-core setting.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias -2.22e-03 0.114 60.6 0.153 0.0533
(SE 1) 5.62e-03 4.99e-03 42.6 4.35e-03 0.0141
MSE 6.36e-04 0.0135 39888 0.0237 6.6e-03
(SE 2) 1.41e-04 1.24e-03 33456 1.4e-03 2.71e-03
Time (s) 11.1 158 2.22 109 668 624
2 Bias -5.19e-03 -0.0564 -0.462 0.0378 0.0613
(SE 1) 5.27e-03 4.41e-03 0.0716 4.2e-03 0.0175
MSE 5.82e-04 3.57e-03 0.316 1.78e-03 9.59e-03
(SE 2) 1.57e-04 4.21e-04 0.0317 3.89e-04 3.91e-03
Time (s) 17.6 608 71.8 272 1148 1102
3 Bias -8.59e-03 -0.111 -0.396 -0.0206 0.0767
(SE 1) 4.97e-03 2.94e-03 0.0294 3.65e-03 0.0172
MSE 5.67e-04 0.0125 0.174 6.91e-04 0.0115
(SE 2) 1.59e-04 6.08e-04 0.0213 1.48e-04 3.67e-03
Time (s) 36.5 1030 127 450 1735
4 Bias -6.38e-03 -0.223 0.0759 -0.0961 0.0862
(SE 1) 5.44e-03 2.97e-03 0.0311 4.65e-03 0.0169
MSE 6.31e-04 0.05 0.0251 9.67e-03 0.0129
(SE 2) 1.55e-04 1.29e-03 0.0105 8.1e-04 3.36e-03
Time (s) 120 3033 1131 1358 2570
5 Bias -2.65e-03 -0.305 0.521 -0.0627 0.0934
(SE 1) 6.35e-03 3.13e-03 0.0289 6.29e-03 0.0181
MSE 8.14e-04 0.0931 0.288 4.72e-03 0.0149
(SE 2) 2.82e-04 1.89e-03 0.0327 7.93e-04 3.82e-03
Time (s) 498 11774 6145 9601 3347
Table 88: Bias and MSE for the GP marginal variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.05$ and a sample size of N=100’000 in the single-core setting.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias 2.93e-04 -0.342 0.861 -0.343 -0.0763 9.16e-03
(SE 1) 6.05e-04 1.46e-03 0.0344 1.14e-03 1.36e-03 9.08e-04
MSE 7.4e-06 0.117 0.764 0.118 5.85e-03 9.95e-05
(SE 2) 1.97e-06 1.01e-03 0.0508 7.83e-04 2.04e-04 1.66e-05
Time (s) 11.1 158 2.22 109 668 624
2 Bias 3.21e-04 -0.263 0.646 -0.277 -0.0765 4.92e-03
(SE 1) 6.75e-04 1.14e-03 0.0179 8.58e-04 1.39e-03 8.5e-04
MSE 9.21e-06 0.0691 0.424 0.0767 5.89e-03 3.8e-05
(SE 2) 2.76e-06 6.06e-04 0.0227 4.76e-04 2.14e-04 8.16e-06
Time (s) 17.6 608 71.8 272 1148 1102
3 Bias 2.54e-04 -0.236 0.596 -0.244 3.23e-03
(SE 1) 7.25e-04 8.48e-04 0.0177 6.29e-04 8.36e-04
MSE 1.06e-05 0.0557 0.362 0.0594 2.37e-05
(SE 2) 3.6e-06 4.03e-04 0.02 3.05e-04 6.03e-06
Time (s) 36.5 1030 127 450 1735
4 Bias 4.74e-04 -0.178 0.307 -0.179 2.94e-03
(SE 1) 7.24e-04 4.72e-04 3.68e-03 7.05e-04 9.33e-04
MSE 1.07e-05 0.0315 0.0948 0.0321 2.52e-05
(SE 2) 3.42e-06 1.68e-04 2.24e-03 2.52e-04 6.6e-06
Time (s) 120 3033 1131 1358 2570
5 Bias 4.2e-04 -0.108 0.13 -0.0991 3.13e-03
(SE 1) 7.45e-04 5.11e-04 2.4e-03 6.11e-04 8.76e-04
MSE 1.13e-05 0.0116 0.0171 9.84e-03 2.43e-05
(SE 2) 3.6e-06 1.1e-04 6.09e-04 1.22e-04 5.17e-06
Time (s) 498 11774 6145 9601 3347
Table 89: Bias and MSE for the error term variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.05$ and a sample size of N=100’000 in the single-core setting.
B.11.3 Prediction
Figure 39: Average RMSE and log-score on simulated datasets with an practical range of $0.05$ and a sample size N=100’000 in the single-core setting. Predictions are done using the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.291 0.488 0.575 0.501 0.578 0.335
(SE 1) 5.14e-04 4.87e-04 6.78e-04 5.11e-04 9.29e-04 7.12e-04
Log-score 0.179 0.709 0.867 0.731 0.874 0.307
(SE 2) 1.81e-03 8.23e-04 1.25e-03 9.08e-04 1.78e-03 2.21e-03
Time (s) 2.45 24.4 68.8 183 375 5.67
2 RMSE 0.276 0.441 0.564 0.465 0.572 0.312
(SE 1) 6.32e-04 4.06e-04 8.08e-04 4.75e-04 9e-04 7.1e-04
Log-score 0.124 0.62 0.852 0.665 0.862 0.238
(SE 2) 2.41e-03 6.97e-04 1.65e-03 8.82e-04 1.74e-03 2.48e-03
Time (s) 4.52 71.4 75.2 231 499 11.1
3 RMSE 0.274 0.426 0.559 0.444 0.303
(SE 1) 6.51e-04 3.82e-04 6.78e-04 4.04e-04 6.72e-04
Log-score 0.116 0.59 0.844 0.624 0.209
(SE 2) 2.47e-03 6.62e-04 1.35e-03 7.59e-04 2.43e-03
Time (s) 10.1 115 81.8 271 17.6
4 RMSE 0.273 0.391 0.516 0.404 0.298
(SE 1) 6.78e-04 3.33e-04 6.69e-04 3.47e-04 6.69e-04
Log-score 0.11 0.516 0.773 0.541 0.193
(SE 2) 2.58e-03 5.9e-04 1.5e-03 6.91e-04 2.51e-03
Time (s) 27.6 298 204 430 23.8
5 RMSE 0.272 0.345 0.455 0.342 0.295
(SE 1) 6.77e-04 2.58e-04 7.15e-04 2.64e-04 6.45e-04
Log-score 0.11 0.402 0.641 0.385 0.184
(SE 2) 2.58e-03 4.96e-04 1.78e-03 5.56e-04 2.42e-03
Time (s) 91.3 1175 855 1323 32.4
Table 90: Average RMSE and log-score on the training set, standard errors and time needed for making predictions on the training set on simulated datasets with an effective range of $0.05$ and a sample size of N=100’000 in the single-core setting.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.316 0.629 0.994 0.676 0.99 0.383
(SE 1) 8.27e-04 1.82e-03 2.7e-03 1.87e-03 3.58e-03 1.17e-03
Log-score 0.255 1.02 1.41 1.07 1.41 0.426
(SE 2) 2.64e-03 1.34e-03 2.7e-03 1.51e-03 3.61e-03 2.54e-03
Time (s) 2.98 24.7 14.4 181 374 42.2
2 RMSE 0.301 0.477 0.945 0.551 0.955 0.351
(SE 1) 9.27e-04 1.09e-03 3.71e-03 1.52e-03 2.91e-03 1.15e-03
Log-score 0.206 0.846 1.36 0.931 1.37 0.344
(SE 2) 3.2e-03 7.16e-04 4.02e-03 1.18e-03 3.11e-03 3.15e-03
Time (s) 5.19 72.1 21.4 230 494 58.1
3 RMSE 0.298 0.439 0.924 0.491 0.34
(SE 1) 9.39e-04 8.97e-04 2.93e-03 1.19e-03 1.08e-03
Log-score 0.197 0.793 1.34 0.856 0.311
(SE 2) 3.25e-03 5.94e-04 2.64e-03 1.03e-03 3.21e-03
Time (s) 11.1 116 27.4 270 70.6
4 RMSE 0.296 0.377 0.765 0.405 0.333
(SE 1) 9.8e-04 5.97e-04 2.45e-03 9.44e-04 1.12e-03
Log-score 0.191 0.678 1.16 0.72 0.292
(SE 2) 3.41e-03 4.53e-04 3.13e-03 7.29e-04 3.35e-03
Time (s) 28.9 298 139 436 84.9
5 RMSE 0.296 0.333 0.605 0.336 0.328
(SE 1) 9.74e-04 5.97e-04 2.3e-03 6.33e-04 1.14e-03
Log-score 0.19 0.524 0.925 0.509 0.28
(SE 2) 3.39e-03 5.91e-04 3.21e-03 7.77e-04 3.44e-03
Time (s) 92.9 1175 743 1321 104
Table 91: Average RMSE and log-score on the test “interpolation” set, standard errors and time needed for making predictions on the test “interpolation” set on simulated datasets with an practical range of $0.05$ and a sample size of N=100’000 in the single-core setting.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.977 0.99 0.991 0.99 1.06 0.982
(SE 1) 6.35e-03 7.1e-03 7.12e-03 7.11e-03 0.0304 7.16e-03
Log-score 1.39 1.41 1.41 1.41 1.48 1.4
(SE 2) 6.45e-03 6.97e-03 6.99e-03 6.98e-03 0.0321 7.27e-03
Time (s) 32.6 46.9 28.9 232 367 40.5
2 RMSE 0.977 0.989 0.991 0.99 19.9 0.981
(SE 1) 6.23e-03 7.07e-03 7.13e-03 7.09e-03 4.66 7.13e-03
Log-score 1.39 1.41 1.41 1.41 15.5 1.4
(SE 2) 6.27e-03 6.93e-03 7e-03 6.96e-03 3.46 7.23e-03
Time (s) 36.8 96.5 42.8 285 486 54.8
3 RMSE 0.976 0.988 0.991 0.989 0.98
(SE 1) 6.16e-03 7.05e-03 7.13e-03 7.07e-03 7.15e-03
Log-score 1.39 1.41 1.41 1.41 1.4
(SE 2) 6.16e-03 6.91e-03 7e-03 6.94e-03 7.27e-03
Time (s) 48.5 143 54.7 329 67.4
4 RMSE 0.976 0.986 0.989 0.988 0.98
(SE 1) 6.16e-03 7.01e-03 7.15e-03 7.04e-03 7.07e-03
Log-score 1.39 1.4 1.41 1.41 1.39
(SE 2) 6.17e-03 6.86e-03 7.03e-03 6.9e-03 7.19e-03
Time (s) 84.5 339 277 524 80.3
5 RMSE 0.976 0.983 0.985 0.984 0.98
(SE 1) 6.17e-03 6.9e-03 7e-03 7e-03 7.02e-03
Log-score 1.39 1.4 1.4 1.4 1.39
(SE 2) 6.18e-03 6.75e-03 6.89e-03 6.85e-03 7.13e-03
Time (s) 214 1322 1514 1642 97.1
Table 92: Average RMSE and log-score on the test “extrapolation” set, standard errors and time needed for making predictions on the test “extrapolation” set on simulated datasets with an practical range of $0.05$ and a sample size of N=100’000 in the single-core setting.
B.12 Simulated data with $N=100^{\prime}000$ and practical range of $0.2$, single-core setting
Tuning parameter
Vecchia nb. neighbours 5 10 20 40 80
Tapering num. non-zero entries 8 17 22 40 111
FITC num. inducing points 20 200 275 900 2000
Full-scale num. inducing points 26 91 122 250 650
num. non-zero entries 6 11 15 27 76
FRK num. resolutions 1 2 exceeds time limit
MRA num. knots per partition 1 2 3 4 5
SPDE max edge 0.04 0.02 0.005 0.004 0.003
Table 93: Tuning parameters chosen for the comparison on simulated datasets with an practical range of $0.2$ and a sample size of $N=100^{\prime}000$ in the single-core setting.
B.12.1 Log-likelihood evaluation
Figure 40: Average negative log-likelihood on simulated datasets with an practical range $0.2$ of and a sample size N=100’000 in the single-core setting. The true data-generating parameters are used on the left, and two times these values are used on the right.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 116892 148045 156316 144010 112921
(SE) 55.6 469 545 358 62.2
Time (s) 0.24 0.0907 0.0727 0.207 5.08
2 Value 113351 139445 126797 130065 112083
(SE) 56.1 308 208 242 60.3
Time (s) 0.373 0.601 2.19 1.15 11.3
3 Value 111603 136654 122397 125538 111784
(SE) 52.9 259 162 124 60.2
Time (s) 0.804 1.19 3.86 2.23 20.7
4 Value 111025 130565 112694 117692 111617
(SE) 54.2 163 100 81.6 60.2
Time (s) 2.41 7.4 36.9 9.55 31.9
5 Value 110922 122924 111121 112154 111511
(SE) 55.3 75.3 69.6 64.2 61
Time (s) 8.42 47.1 224 78.4 40.7
Table 94: Mean approximate negative log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.2$ of and a sample size N=100’000 in the single-core setting. Log-likelihoods are evaluated at the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 126154 162289 160526 151451 121842
(SE) 28.3 234 348 242 46.9
Time (s) 0.24 0.0906 0.0736 0.207 5.96
2 Value 122577 155846 130399 136310 121216
(SE) 29.4 154 149 215 45.9
Time (s) 0.374 0.602 2.19 1.15 20.1
3 Value 120782 153565 127134 132367 120985
(SE) 29.3 129 108 123 45.8
Time (s) 0.804 1.19 3.86 2.24 39.7
4 Value 120185 148134 121085 125834 120851
(SE) 35.1 80.9 59.7 88.2 46.6
Time (s) 2.42 7.56 36.9 9.57 60
5 Value 120195 140101 120356 121315 120767
(SE) 41.3 37 48.7 54.7 47.5
Time (s) 8.41 46.9 225 79.2 70.5
Table 95: Mean approximate negative log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.2$ of and a sample size N=100’000 in the single-core setting. Log-likelihoods are evaluated at two-times the true data-generating parameters.
B.12.2 Parameter estimation
Figure 41: Bias and MSE of the estimates for the GP range, GP variance end error term variance on simulated datasets with an practical range of $0.2$ and a sample size N=100’000 in the single-core setting,
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias 9.92e-04 0.254 0.0658 0.0136 7.97e-03
(SE 1) 6.83e-04 0.0393 7.37e-03 1.94e-03 9.62e-04
MSE 1.03e-05 0.0953 5.42e-03 2.61e-04 8.11e-05
(SE 2) 3.89e-06 0.0229 1.07e-03 6.03e-05 1.47e-05
Time (s) 5.76 154 1.84 116 477 414 278
2 Bias 1.25e-03 0.622 0.113 0.0147 7.49e-03
(SE 1) 5.7e-04 0.113 0.0269 1.92e-03 8.42e-04
MSE 8.05e-06 0.644 0.0271 2.88e-04 6.96e-05
(SE 2) 2.44e-06 0.181 0.0134 7.13e-05 1.21e-05
Time (s) 11.9 563 75.4 297 817 911 337
3 Bias 1.05e-03 0.933 0.168 0.0176 6.26e-03
(SE 1) 6.09e-04 0.093 0.0721 1.86e-03 1.03e-03
MSE 8.51e-06 1.04 0.132 3.8e-04 5.94e-05
(SE 2) 2.94e-06 0.146 0.0935 7.6e-05 1.07e-05
Time (s) 20.2 998 163 542 1916 4283
4 Bias 9.61e-04 1.1 0.0164 0.0146 6.66e-03
(SE 1) 6.21e-04 0.13 1.31e-03 1.56e-03 5.8e-04
MSE 8.64e-06 1.55 3.03e-04 2.62e-04 5.08e-05
(SE 2) 2.41e-06 0.327 4.97e-05 4.81e-05 8.24e-06
Time (s) 93.7 3050 1348 1662 2644 7334
5 Bias 9.48e-04 1.36 1.84e-03 6.87e-03 6.01e-03
(SE 1) 6.08e-04 0.181 5.61e-04 7.92e-04 7.87e-04
MSE 8.3e-06 2.52 9.67e-06 5.98e-05 4.79e-05
(SE 2) 2.51e-06 0.848 2.2e-06 1.22e-05 8.65e-06
Time (s) 417 9686 5921 8675 3461 13486
Table 96: Bias and MSE for the GP range, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.2$ and a sample size of N=100’000 in the single-core setting.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias 0.0465 0.137 -0.413 0.151 0.294
(SE 1) 0.0249 0.0184 0.0553 0.0232 0.0465
MSE 0.0145 0.0256 0.232 0.0335 0.127
(SE 2) 4.63e-03 6.09e-03 0.0449 8.07e-03 0.025
Time (s) 5.76 154 1.84 116 477 414 278
2 Bias 0.0482 -0.0582 17.9 0.374 0.356
(SE 1) 0.0249 0.0142 10.3 0.0788 0.045
MSE 0.0147 7.4e-03 2434 0.264 0.165
(SE 2) 4.47e-03 1.6e-03 2024 0.0908 0.0252
Time (s) 11.9 563 75.4 297 817 911 337
3 Bias 0.0324 -0.126 325 0.558 0.307
(SE 1) 0.0227 0.013 259 0.0913 0.0497
MSE 0.0113 0.0192 1443667 0.479 0.141
(SE 2) 4.48e-03 3.01e-03 1335536 0.148 0.0247
Time (s) 20.2 998 163 542 1916 4283
4 Bias 0.0294 -0.281 0.911 0.666 0.342
(SE 1) 0.024 0.0116 0.109 0.0959 0.0338
MSE 0.0124 0.0816 1.07 0.628 0.139
(SE 2) 4.38e-03 6.21e-03 0.274 0.157 0.0193
Time (s) 93.7 3050 1348 1662 2644 7334
5 Bias 0.0256 -0.484 0.129 0.374 0.302
(SE 1) 0.0229 6.79e-03 0.0228 0.0493 0.0418
MSE 0.0111 0.236 0.027 0.188 0.125
(SE 2) 5.31e-03 6.53e-03 6.1e-03 0.0464 0.0214
Time (s) 417 9686 5921 8675 3461 13486
Table 97: Bias and MSE for the GP marginal variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.2$ and a sample size of N=100’000 in the single-core setting.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias 1.67e-04 -0.348 0.505 -0.222 -0.0135 1.29e-03 0.0354
(SE 1) 5.55e-04 3.25e-03 0.0328 5.75e-03 1.42e-03 5.74e-04 9.8e-04
MSE 6.2e-06 0.121 0.277 0.0498 2.2e-04 7.92e-06 1.27e-03
(SE 2) 3e-06 2.26e-03 0.0339 2.53e-03 4.22e-05 2.59e-06 7.15e-05
Time (s) 5.76 154 1.84 116 477 414 278
2 Bias 2.27e-04 -0.268 0.116 -0.139 -0.0126 6.24e-04 5.7e-03
(SE 1) 5.66e-04 2.43e-03 4.38e-03 2.03e-03 1.47e-03 5.98e-04 5.2e-04
MSE 6.45e-06 0.0718 0.0139 0.0194 1.99e-04 7.19e-06 3.76e-05
(SE 2) 2.99e-06 1.3e-03 9.6e-04 5.62e-04 4.79e-05 3.09e-06 6.12e-06
Time (s) 11.9 563 75.4 297 817 911 337
3 Bias 1.18e-04 -0.24 0.0723 -0.115 7.42e-04 1.39e-03
(SE 1) 5.12e-04 2.31e-03 3.43e-03 1.42e-03 4.99e-04 5.25e-04
MSE 5.25e-06 0.0575 5.46e-03 0.0132 5.27e-06 7.17e-06
(SE 2) 2.75e-06 1.11e-03 4.68e-04 3.27e-04 1.89e-06 2.41e-06
Time (s) 20.2 998 163 542 1916 4283
4 Bias 1.16e-04 -0.177 -6.24e-03 -0.0689 7.37e-04 1.46e-03
(SE 1) 4.99e-04 2.18e-03 7.67e-04 9.15e-04 5.17e-04 5.22e-04
MSE 4.99e-06 0.0316 5.07e-05 4.76e-03 5.62e-06 7.31e-06
(SE 2) 2.63e-06 7.8e-04 8.84e-06 1.27e-04 2.24e-06 2.38e-06
Time (s) 93.7 3050 1348 1662 2644 7334
5 Bias 1.51e-04 -0.105 -6.13e-03 -0.0254 5.57e-04 1.52e-03
(SE 1) 5.08e-04 9.91e-04 6.28e-04 6.41e-04 5.22e-04 5.22e-04
MSE 5.18e-06 0.011 4.55e-05 6.53e-04 5.48e-06 7.48e-06
(SE 2) 2.57e-06 2.07e-04 7.25e-06 3.54e-05 2.18e-06 2.4e-06
Time (s) 417 9686 5921 8675 3461 13486
Table 98: Bias and MSE for the error term variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.2$ and a sample size of N=100’000 in the single-core setting.
B.12.3 Prediction
Figure 42: Average RMSE and log-score on simulated datasets with an practical range of $0.2$ and a sample size N=100’000 in the single-core setting. Predictions are done using the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.224 0.486 0.559 0.477 0.555 0.164 0.199
(SE 1) 5.62e-04 1.38e-03 2.09e-03 1.54e-03 3.19e-03 9.71e-04 2.53e-03
Log-score -0.0785 0.705 0.874 0.708 0.83 -0.411 2.32
(SE 2) 2.47e-03 2.17e-03 5.63e-03 4.3e-03 5.75e-03 5.95e-03 0.0943
Time (s) 2.46 24.4 64.9 183 259 7.38 65.3
2 RMSE 0.168 0.433 0.399 0.391 0.477 0.148 0.133
(SE 1) 7.1e-04 7.52e-04 2.27e-03 8.34e-04 3.64e-03 1.08e-03 1.55e-03
Log-score -0.368 0.606 0.566 0.539 0.679 -0.507 -0.542
(SE 2) 4.17e-03 1.15e-03 4.36e-03 4.59e-03 7.57e-03 7.35e-03 0.0169
Time (s) 4.54 72.8 73.9 237 351 20.2 79.4
3 RMSE 0.137 0.415 0.355 0.352 0.143 0.124
(SE 1) 9.81e-04 6.16e-04 2.17e-03 7.2e-04 1.13e-03 1.47e-03
Log-score -0.575 0.57 0.445 0.439 -0.543 -0.671
(SE 2) 7.14e-03 9.47e-04 6.18e-03 2.99e-03 7.88e-03 0.0132
Time (s) 10.3 111 80.4 286 42.9 751
4 RMSE 0.129 0.374 0.195 0.263 0.14 0.124
(SE 1) 1.06e-03 4.7e-04 2.1e-03 8.37e-04 1.13e-03 1.47e-03
Log-score -0.634 0.483 -0.189 0.16 -0.564 -0.671
(SE 2) 8.02e-03 7.53e-04 9.57e-03 3e-03 8.15e-03 0.0132
Time (s) 27.5 322 207 449 80.9 1282
5 RMSE 0.127 0.313 0.139 0.166 0.138 0.124
(SE 1) 1.14e-03 4.51e-04 1.74e-03 9.66e-04 1.17e-03 1.47e-03
Log-score -0.647 0.334 -0.538 -0.304 -0.578 -0.671
(SE 2) 8.84e-03 7.72e-04 9.71e-03 4.15e-03 8.5e-03 0.0133
Time (s) 90.9 1126 905 1311 102 2236
Table 99: Average RMSE and log-score on the training set, standard errors and time needed for making predictions on the training set on simulated datasets with an effective range of $0.2$ and a sample size of N=100’000 in the single-core setting.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.225 0.622 0.922 0.606 0.897 0.168 0.201
(SE 1) 5.44e-04 7.81e-03 0.0101 7.38e-03 0.0127 1.13e-03 2.53e-03
Log-score -0.0714 1.01 1.37 1 1.31 -0.386 2.35
(SE 2) 2.37e-03 5.79e-03 0.0113 7.25e-03 0.0142 6.48e-03 0.0934
Time (s) 2.98 24.7 13.6 181 258 47.6 65.6
2 RMSE 0.169 0.449 0.478 0.415 0.648 0.152 0.135
(SE 1) 6.31e-04 3.78e-03 5.36e-03 2.52e-03 8.74e-03 1.12e-03 1.61e-03
Log-score -0.358 0.818 0.756 0.707 0.984 -0.486 -0.524
(SE 2) 3.68e-03 2.46e-03 7.04e-03 6.33e-03 0.0133 7.53e-03 0.0174
Time (s) 5.2 73.6 21.3 236 345 68.3 79.8
3 RMSE 0.139 0.406 0.406 0.356 0.146 0.126
(SE 1) 9.88e-04 2.74e-03 4.8e-03 1.58e-03 1.14e-03 1.53e-03
Log-score -0.56 0.758 0.589 0.573 -0.522 -0.655
(SE 2) 7.07e-03 1.77e-03 7.58e-03 4.11e-03 7.93e-03 0.0135
Time (s) 11.1 113 27.3 284 98.4 759
4 RMSE 0.131 0.332 0.202 0.257 0.143 0.126
(SE 1) 1.11e-03 1.23e-03 2.59e-03 1.43e-03 1.14e-03 1.53e-03
Log-score -0.617 0.624 -0.142 0.232 -0.544 -0.655
(SE 2) 8.24e-03 8.8e-04 0.0102 2.91e-03 8.15e-03 0.0135
Time (s) 28.6 324 140 452 137 1293
5 RMSE 0.129 0.269 0.143 0.165 0.141 0.126
(SE 1) 1.19e-03 6.05e-04 1.89e-03 1.13e-03 1.19e-03 1.53e-03
Log-score -0.631 0.427 -0.509 -0.27 -0.558 -0.655
(SE 2) 8.97e-03 6.07e-04 9.69e-03 4.27e-03 8.62e-03 0.0136
Time (s) 92.8 1127 771 1325 144 2263
Table 100: Average RMSE and log-score on the test “interpolation” set, standard errors and time needed for making predictions on the test “interpolation” set on simulated datasets with an practical range of $0.2$ and a sample size of N=100’000 in the single-core setting.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.927 0.987 0.987 0.986 1.54 0.946 0.933
(SE 1) 0.0199 0.0183 0.018 0.0177 0.139 0.018 0.0264
Log-score 1.31 1.41 1.41 1.41 2.4 1.34 1.34
(SE 2) 0.0168 0.018 0.0177 0.0174 0.341 0.0138 0.0279
Time (s) 32.6 46.8 27.3 232 252 45.7 65.5
2 RMSE 0.925 0.985 0.965 0.978 41.8 0.943 0.922
(SE 1) 0.0203 0.0183 0.0174 0.0168 13.2 0.0193 0.0266
Log-score 1.3 1.41 1.38 1.4 71.6 1.34 1.3
(SE 2) 0.0171 0.018 0.0164 0.0164 19.4 0.0159 0.0233
Time (s) 36.7 97.9 42.6 292 342 64.7 79.8
3 RMSE 0.924 0.985 0.957 0.973 0.937 0.919
(SE 1) 0.0206 0.0183 0.0179 0.0178 0.0209 0.0257
Log-score 1.3 1.41 1.37 1.39 1.33 1.29
(SE 2) 0.0173 0.018 0.0166 0.0173 0.018 0.0238
Time (s) 48.5 139 54.5 345 101 759
4 RMSE 0.921 0.983 0.932 0.959 0.936 0.919
(SE 1) 0.0215 0.0183 0.0208 0.0184 0.021 0.0256
Log-score 1.29 1.4 1.33 1.37 1.33 1.29
(SE 2) 0.0189 0.0179 0.019 0.0173 0.0181 0.0238
Time (s) 83.9 367 281 542 147 1293
5 RMSE 0.919 0.977 0.922 0.938 0.934 0.919
(SE 1) 0.0226 0.0182 0.0226 0.0202 0.0217 0.0256
Log-score 1.29 1.4 1.3 1.34 1.32 1.29
(SE 2) 0.0207 0.0178 0.0208 0.0184 0.0189 0.0238
Time (s) 213 1275 1535 1645 148 2263
Table 101: Average RMSE and log-score on the test “extrapolation” set, standard errors and time needed for making predictions on the test “extrapolation” set on simulated datasets with an practical range of $0.2$ and a sample size of N=100’000 in the single-core setting.
B.13 Simulated data with $N=100^{\prime}000$ and practical range of $0.5$, single-core setting
Tuning parameter
Vecchia nb. neighbours 5 10 20 40 80
Tapering num. non-zero entries 8 17 22 40 111
FITC num. inducing points 20 200 275 900 2000
Full-scale num. inducing points 26 91 122 250 650
num. non-zero entries 6 11 15 27 76
FRK num. resolutions 1 2 exceeds time limit
MRA num. knots per partition 1 2 3 4 5
SPDE max edge 0.04 0.02 0.005 0.004 0.003
Table 102: Tuning parameters chosen for the comparison on simulated datasets with an practical range of $0.5$ and a sample size of $N=100^{\prime}000$ in the single-core setting.
B.13.1 Log-likelihood evaluation
Figure 43: Average negative log-likelihood on simulated datasets with an practical range $0.5$ of and a sample size N=100’000 in the single-core setting. The true data-generating parameters are used on the left, and two times these values are used on the right.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 115906 147184 134682 127030 109523
(SE) 63.1 951 815 388 56
Time (s) 0.236 0.0906 0.063 0.207 5.76
2 Value 112200 138802 110408 113648 109143
(SE) 48.5 616 90.5 137 54.3
Time (s) 0.373 0.606 2.03 1.17 21.1
3 Value 110154 136072 109635 111886 109017
(SE) 48.1 515 70.7 91.4 53.4
Time (s) 0.801 1.2 3.68 2.24 40.7
4 Value 109163 130090 108700 109577 108943
(SE) 44.5 318 52.7 70.4 53.5
Time (s) 2.43 6.42 36.6 9.64 56.1
5 Value 108789 122479 108648 108746 108899
(SE) 48 131 50.2 56.3 52.9
Time (s) 8.5 51.8 222 75.4 71.6
Table 103: Mean approximate negative log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.5$ of and a sample size N=100’000 in the single-core setting. Log-likelihoods are evaluated at the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 125459 161867 137360 132517 118933
(SE) 32.1 476 568 324 39.4
Time (s) 0.236 0.0907 0.064 0.206 7.87
2 Value 121725 155542 119092 121496 118636
(SE) 25.2 308 48.4 88.4 35.6
Time (s) 0.373 0.606 2.03 1.17 24.9
3 Value 119654 153294 118672 120296 118532
(SE) 25.6 257 39.2 60.3 35
Time (s) 0.802 1.2 3.68 2.27 46.3
4 Value 118639 147924 118248 118735 118473
(SE) 24 159 34.3 41.5 35.8
Time (s) 2.43 6.27 36.6 9.76 65.1
5 Value 118250 139923 118224 118273 118436
(SE) 26.5 65.3 33.7 35.2 35.6
Time (s) 8.5 52 221 75.4 82.5
Table 104: Mean approximate negative log-likelihood, standard error and time needed for evaluating the approximate log-likelihood on simulated datasets with an practical range of $0.5$ of and a sample size N=100’000 in the single-core setting. Log-likelihoods are evaluated at two-times the true data-generating parameters.
B.13.2 Parameter estimation
Figure 44: Bias and MSE of the estimates for the GP range, GP variance end error term variance on simulated datasets with an practical range of $0.5$ and a sample size N=100’000 in the single-core setting,
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias -3.93e-03 0.0925 2.73 0.0407 0.0246
(SE 1) 2.61e-03 0.029 1.81 0.0108 2.58e-03
MSE 1.52e-04 0.0254 72.7 3.98e-03 7.33e-04
(SE 2) 4.46e-05 0.0126 54.8 1.61e-03 1.23e-04
Time (s) 5.17 128 2.35 146 668 389 215
2 Bias -2.31e-03 0.702 0.106 0.0743 0.022
(SE 1) 3.19e-03 0.121 0.0464 0.0141 2.53e-03
MSE 2.09e-04 0.784 0.0542 9.51e-03 6.06e-04
(SE 2) 7.12e-05 0.191 0.0477 3.17e-03 1.08e-04
Time (s) 8.45 577 75.9 315 1129 1210 307
3 Bias -2.15e-03 0.989 0.0573 0.037 0.0205
(SE 1) 2.49e-03 0.085 0.013 8.16e-03 2.51e-03
MSE 1.29e-04 1.12 6.64e-03 2.7e-03 5.42e-04
(SE 2) 2.81e-05 0.184 3.32e-03 9.48e-04 1.02e-04
Time (s) 27.3 1058 118 553 2623 4932
4 Bias -2.77e-03 1.08 -2.74e-03 0.0177 0.021
(SE 1) 2.49e-03 0.087 2.28e-03 7.5e-03 2.67e-03
MSE 1.31e-04 1.32 1.12e-04 1.44e-03 5.78e-04
(SE 2) 2.7e-05 0.216 3.18e-05 7.42e-04 1.05e-04
Time (s) 109 2914 1287 1631 3503 6595
5 Bias -2.51e-03 1.94 -2.65e-03 -4.62e-03 0.021
(SE 1) 2.78e-03 0.192 2.75e-03 2.49e-03 2.69e-03
MSE 1.61e-04 4.49 1.58e-04 1.46e-04 5.78e-04
(SE 2) 4.61e-05 0.996 3.25e-05 4.79e-05 1.08e-04
Time (s) 508 11902 6572 6998 4915 13162
Table 105: Bias and MSE for the GP range, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.5$ and a sample size of N=100’000 in the single-core setting.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias -2.32e-03 0.102 31485 0.879 0.559
(SE 1) 0.0462 0.0378 28417 0.287 0.0683
MSE 0.0428 0.039 17141860498 2.41 0.401
(SE 2) 0.0107 9.24e-03 16599774522 1.33 0.0653
Time (s) 5.17 128 2.35 146 668 389 215
2 Bias -5.99e-03 -0.0914 29.8 4.2 0.586
(SE 1) 0.0443 0.0312 27.5 1.31 0.0721
MSE 0.0393 0.0279 16011 52 0.442
(SE 2) 9.91e-03 8.16e-03 15600 32.6 0.0721
Time (s) 8.45 577 75.9 315 1129 1210 307
3 Bias -0.0101 -0.158 2.85 1.64 0.57
(SE 1) 0.0423 0.0283 1.38 0.423 0.0694
MSE 0.036 0.0409 46.1 6.25 0.417
(SE 2) 8.32e-03 0.0104 41 2.72 0.0683
Time (s) 27.3 1058 118 553 2623 4932
4 Bias -0.0126 -0.304 0.0321 1.1 0.616
(SE 1) 0.0487 0.023 0.0488 0.5 0.0851
MSE 0.0476 0.103 0.0487 6.21 0.517
(SE 2) 0.0126 0.0147 0.0157 5.05 0.11
Time (s) 109 2914 1287 1631 3503 6595
5 Bias -0.0174 -0.508 -0.0125 0.0357 0.615
(SE 1) 0.0585 0.0158 0.0538 0.0526 0.0786
MSE 0.0687 0.263 0.058 0.0567 0.495
(SE 2) 0.0198 0.0163 0.0168 0.0154 0.0834
Time (s) 508 11902 6572 6998 4915 13162
Table 106: Bias and MSE for the GP marginal variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.5$ and a sample size of N=100’000 in the single-core setting.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Bias 1.56e-04 -0.342 0.208 -0.101 -2.3e-03 1.01e-03 3.95e-03
(SE 1) 5.24e-04 6.59e-03 0.0276 4.81e-03 1.2e-03 4.34e-04 4.86e-04
MSE 5.51e-06 0.118 0.0583 0.0106 3.25e-05 4.6e-06 2.01e-05
(SE 2) 1.31e-06 4.43e-03 0.0184 9.44e-04 1.13e-05 1.17e-06 4e-06
Time (s) 5.17 128 2.35 146 668 389 215
2 Bias 4.75e-04 -0.261 -6.29e-04 -0.0374 -2.1e-03 8.96e-04 1.23e-03
(SE 1) 4.24e-04 6.07e-03 1.14e-03 1.08e-03 1.33e-03 4.23e-04 4.17e-04
MSE 3.83e-06 0.0688 2.66e-05 1.42e-03 3.81e-05 4.2e-06 4.82e-06
(SE 2) 9.76e-07 3.1e-03 6.41e-06 8.11e-05 9.53e-06 1.04e-06 1.27e-06
Time (s) 8.45 577 75.9 315 1129 1210 307
3 Bias 6.16e-04 -0.233 -2.86e-03 -0.0308 6.81e-04 1.28e-03
(SE 1) 4.16e-04 5.25e-03 6.46e-04 8.27e-04 4.32e-04 4.19e-04
MSE 3.84e-06 0.0549 1.65e-05 9.62e-04 4.01e-06 4.98e-06
(SE 2) 9.9e-07 2.41e-03 3.35e-06 4.94e-05 1.26e-06 1.28e-06
Time (s) 27.3 1058 118 553 2623 4932
4 Bias 5.9e-04 -0.175 -1.14e-03 -0.0133 7.26e-04 1.29e-03
(SE 1) 4.15e-04 4.03e-03 3.54e-04 6.08e-04 4.46e-04 4.18e-04
MSE 3.79e-06 0.0308 3.79e-06 1.84e-04 4.3e-06 5e-06
(SE 2) 9.01e-07 1.39e-03 8.84e-07 1.6e-05 1.24e-06 1.29e-06
Time (s) 109 2914 1287 1631 3503 6595
5 Bias 6.89e-04 -0.102 3.47e-04 -2.49e-03 5.64e-04 1.3e-03
(SE 1) 4.42e-04 2.39e-03 4.01e-04 3.77e-04 4.46e-04 4.18e-04
MSE 4.39e-06 0.0105 3.34e-06 9.02e-06 4.09e-06 5e-06
(SE 2) 1.04e-06 4.82e-04 6.79e-07 1.83e-06 1.2e-06 1.29e-06
Time (s) 508 11902 6572 6998 4915 13162
Table 107: Bias and MSE for the error term variance, standard errors and time needed for estimating covariance parameters on simulated datasets with an practical range of $0.5$ and a sample size of N=100’000 in the single-core setting.
B.13.3 Prediction
Figure 45: Average RMSE and log-score on simulated datasets with an practical range of $0.5$ and a sample size N=100’000 in the single-core setting. Predictions are done using the true data-generating parameters.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.22 0.488 0.463 0.384 0.476 0.1 0.0841
(SE 1) 5.89e-04 3.35e-03 7.02e-03 3.11e-03 8.16e-03 1.38e-03 2.14e-03
Log-score -0.0955 0.707 0.79 0.573 0.675 -0.904 -0.889
(SE 2) 2.68e-03 5.26e-03 0.0443 0.0162 0.0171 0.0153 0.0487
Time (s) 2.45 24.4 64.8 183 379 14.4 60.1
2 RMSE 0.159 0.434 0.156 0.227 0.276 0.0897 0.0749
(SE 1) 3.8e-04 1.78e-03 1.8e-03 2.56e-03 7.11e-03 1.48e-03 1.93e-03
Log-score -0.422 0.607 -0.371 0.0287 0.129 -1.01 -1.17
(SE 2) 2.4e-03 2.71e-03 0.0111 0.011 0.0254 0.018 0.0295
Time (s) 4.5 71.6 73 239 501 63.5 87.1
3 RMSE 0.116 0.416 0.13 0.19 0.0859 0.0742
(SE 1) 4.91e-04 1.38e-03 2.15e-03 1.59e-03 1.51e-03 1.93e-03
Log-score -0.738 0.571 -0.575 -0.15 -1.05 -1.18
(SE 2) 4.29e-03 2.1e-03 0.0147 9.73e-03 0.019 0.0294
Time (s) 10.1 108 79.5 290 157 768
4 RMSE 0.0893 0.374 0.0787 0.123 0.0838 0.0743
(SE 1) 8.79e-04 7.33e-04 2.28e-03 1.92e-03 1.58e-03 1.93e-03
Log-score -0.998 0.482 -1.11 -0.614 -1.08 -1.18
(SE 2) 0.0101 1.14e-03 0.0239 0.0134 0.0203 0.0294
Time (s) 27.3 315 205 466 182 955
5 RMSE 0.0794 0.312 0.0749 0.082 0.0824 0.0743
(SE 1) 1.32e-03 3.35e-04 1.82e-03 2.36e-03 1.63e-03 1.93e-03
Log-score -1.12 0.331 -1.17 -1.06 -1.09 -1.18
(SE 2) 0.0173 5.85e-04 0.0235 0.0233 0.021 0.0295
Time (s) 90.8 1289 866 1302 239 1908
Table 108: Average RMSE and log-score on the training set, standard errors and time needed for making predictions on the training set on simulated datasets with an effective range of $0.5$ and a sample size of N=100’000 in the single-core setting.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.22 0.623 0.61 0.435 0.647 0.101 0.0845
(SE 1) 5.17e-04 0.0177 0.019 7.04e-03 0.0207 1.47e-03 2.19e-03
Log-score -0.0947 1.02 1.06 0.739 0.979 -0.893 -0.884
(SE 2) 2.36e-03 0.0129 0.0515 0.0188 0.0315 0.0161 0.0494
Time (s) 2.99 24.7 13.6 181 375 44.8 60.5
2 RMSE 0.159 0.449 0.158 0.232 0.3 0.0907 0.0752
(SE 1) 4.1e-04 8.6e-03 2.14e-03 3.84e-03 9.33e-03 1.59e-03 1.98e-03
Log-score -0.421 0.818 -0.347 0.0819 0.213 -1 -1.16
(SE 2) 2.6e-03 5.48e-03 0.0115 0.012 0.0305 0.019 0.0301
Time (s) 5.17 72.5 20.8 235 493 80.1 87.6
3 RMSE 0.116 0.405 0.132 0.193 0.0868 0.0745
(SE 1) 5.29e-04 6.17e-03 2.53e-03 1.91e-03 1.6e-03 1.98e-03
Log-score -0.736 0.757 -0.556 -0.114 -1.04 -1.18
(SE 2) 4.62e-03 3.86e-03 0.0153 9.89e-03 0.0197 0.03
Time (s) 11 110 26.6 287 133 775
4 RMSE 0.0897 0.331 0.0791 0.123 0.0845 0.0745
(SE 1) 9.11e-04 2.32e-03 2.38e-03 2.19e-03 1.64e-03 1.98e-03
Log-score -0.994 0.621 -1.11 -0.596 -1.07 -1.18
(SE 2) 0.0104 1.51e-03 0.0241 0.0143 0.0206 0.03
Time (s) 28.7 309 139 464 150 964
5 RMSE 0.0797 0.267 0.0752 0.0825 0.0831 0.0745
(SE 1) 1.35e-03 3.93e-04 1.88e-03 2.49e-03 1.7e-03 1.98e-03
Log-score -1.11 0.422 -1.17 -1.06 -1.08 -1.18
(SE 2) 0.0175 3.16e-04 0.0239 0.0237 0.0215 0.0301
Time (s) 92.6 1194 743 1302 192 1927
Table 109: Average RMSE and log-score on the test “interpolation” set, standard errors and time needed for making predictions on the test “interpolation” set on simulated datasets with an practical range of $0.5$ and a sample size of N=100’000 in the single-core setting.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.835 0.994 0.936 0.926 2.64 0.84 0.821
(SE 1) 0.0574 0.0484 0.046 0.0492 0.965 0.0599 0.0641
Log-score 1.16 1.42 1.36 1.35 18.5 1.19 1.1
(SE 2) 0.0546 0.051 0.0473 0.0498 13.3 0.0524 0.0676
Time (s) 32.6 46.8 27.2 231 368 44.5 60.4
2 RMSE 0.832 0.992 0.847 0.862 17.4 0.832 0.819
(SE 1) 0.0578 0.0483 0.0552 0.0541 2.15 0.06 0.0628
Log-score 1.15 1.42 1.2 1.25 32.4 1.16 1.09
(SE 2) 0.055 0.0509 0.0546 0.0526 7.14 0.0534 0.0723
Time (s) 36.7 96.8 41.6 289 485 86.5 87.6
3 RMSE 0.829 0.992 0.839 0.852 0.838 0.816
(SE 1) 0.0581 0.0483 0.0555 0.0574 0.0597 0.0614
Log-score 1.14 1.42 1.18 1.23 1.16 1.09
(SE 2) 0.0552 0.0508 0.0552 0.0569 0.0525 0.0738
Time (s) 48.3 136 53.1 347 201 775
4 RMSE 0.824 0.989 0.83 0.842 0.835 0.816
(SE 1) 0.0584 0.0482 0.0606 0.0554 0.0602 0.0613
Log-score 1.12 1.42 1.13 1.19 1.15 1.09
(SE 2) 0.0545 0.0507 0.0632 0.0554 0.054 0.0743
Time (s) 83.9 351 277 555 211 964
5 RMSE 0.822 0.983 0.818 0.838 0.845 0.816
(SE 1) 0.0572 0.0481 0.0611 0.0576 0.0626 0.0613
Log-score 1.1 1.41 1.09 1.15 1.16 1.09
(SE 2) 0.0519 0.0504 0.0611 0.0588 0.0576 0.0741
Time (s) 212 1337 1489 1618 280 1927
Table 110: Average RMSE and log-score on the test “extrapolation” set, standard errors and time needed for making predictions on the test “extrapolation” set on simulated datasets with an practical range of $0.5$ and a sample size of N=100’000 in the single-core setting.
B.14 House price dataset, single-core setting
Tuning parameter
Vecchia nb. neighbours 5 10 20 40 80
Tapering avg. nb. non-zeros / row 32 100 200 512 739
FITC nb. inducing points 220 500 1000 2200 3700
Full-scale nb. inducing points 106 252 444 1050 1900
avg. nb. non-zeros / row 17 36 71 256 420
FRK nb. resolutions 1 2 3 exceeds time limit
MRA nb. knots per partition 1 6 15 22 32
SPDE max edge 3000 1100 650 298 200
Table 111: Tuning parameters chosen for the comparison on the house price dataset in the single-core setting.
B.14.1 Prediction
Figure 46: RMSE, log-score and CRPS on the “interpolation” and “extrapolation” test sets of the house price dataset in the single-core setting. Every approximation makes predictions using its parameter estimates. The dashed lines correspond to the results from the exact calculations.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.325 0.382 0.383 0.345 0.544 0.337 0.41
Log-score 0.262 0.387 0.395 0.29 0.81 0.292 0.524
CRPS 0.17 0.196 0.2 0.18 0.298 0.175 0.219
Time (s) 1.01 20.2 21.3 17.6 58.5 99.2 59.7
2 RMSE 0.324 0.345 0.355 0.331 0.451 0.326 0.347
Log-score 0.257 0.292 0.329 0.256 0.622 0.266 0.347
CRPS 0.169 0.178 0.186 0.171 0.243 0.171 0.184
Time (s) 2.5 157 73.7 67 163 211 138
3 RMSE 0.324 0.332 0.346 0.323 0.397 0.325 0.336
Log-score 0.258 0.266 0.29 0.243 0.492 0.261 0.304
CRPS 0.169 0.172 0.179 0.168 0.213 0.17 0.176
Time (s) 5.02 437 290 224 1364 688 393
4 RMSE 0.323 0.324 0.334 0.324 0.325 0.326
Log-score 0.257 0.252 0.272 0.243 0.261 0.266
CRPS 0.169 0.168 0.173 0.168 0.17 0.17
Time (s) 17.5 1893 1541 1115 1008 1364
5 RMSE 0.323 0.322 0.327 0.323 0.324 0.324
Log-score 0.257 0.25 0.26 0.25 0.259 0.258
CRPS 0.169 0.168 0.17 0.168 0.17 0.169
Time (s) 72.2 4229 4260 2769 1685 2907
Table 112: Test “interpolation” RMSE, log-score, CRPS and time needed for making predictions on the “interpolation” test set on the house price dataset in the single-core setting. Every approximation makes predictions using its parameter estimates.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE 0.597 0.679 0.559 0.565 1.45 0.581 0.708
Log-score 0.929 1.05 0.971 0.921 3.75 0.952 1.24
CRPS 0.336 0.378 0.332 0.326 0.976 0.335 0.411
Time (s) 1.11 20 21.3 16.9 58.5 99.2 59.6
2 RMSE 0.596 0.658 0.574 0.571 10.5 0.602 0.579
Log-score 0.925 1.02 0.955 0.943 23.7 0.948 0.921
CRPS 0.335 0.366 0.333 0.33 6.35 0.341 0.33
Time (s) 2.6 155 73.8 66.2 163 212 138
3 RMSE 0.595 0.641 0.583 0.584 2.76 0.598 0.582
Log-score 0.933 0.982 0.965 0.948 2.59 0.941 0.919
CRPS 0.336 0.355 0.337 0.335 1.5 0.338 0.331
Time (s) 5.15 433 290 221 1364 681 393
4 RMSE 0.597 0.617 0.595 0.592 0.598 0.591
Log-score 0.934 0.931 0.942 0.952 0.941 0.921
CRPS 0.337 0.342 0.337 0.338 0.338 0.334
Time (s) 17.7 1876 1542 1106 998 1364
5 RMSE 0.597 0.608 0.601 0.595 0.595 0.593
Log-score 0.933 0.916 0.941 0.939 0.935 0.92
CRPS 0.337 0.338 0.34 0.337 0.336 0.334
Time (s) 72.6 4198 4260 2746 1672 2907
Table 113: Test “extrapolation” RMSE, log-score, CRPS and time needed for making predictions on the “extrapolation” test set on the house price dataset in the single-core setting. Every approximation makes predictions using its parameter estimates.
B.14.2 Log-likelihood evaluation
Figure 47: Negative log-likelihood on the house price dataset in the single-core setting. The parameters estimated via exact calculations are used on the left and the pointwise doubling of them on the right. The dashed lines correspond to the exact negative log-likelihood.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 6184 12120 9047 9230 6750
Time (s) 0.0437 0.0861 0.428 0.239 0.481
2 Value 5862 9114 6893 6770 6063
Time (s) 0.0673 0.418 2.07 1.1 2.77
3 Value 5751 7847 6055 5783 5853
Time (s) 0.142 1.37 8.27 2.88 10.5
4 Value 5721 6732 5753 5548 5822
Time (s) 0.426 6.76 51.2 19.5 15.4
5 Value 5714 6446 5731 5633 5802
Time (s) 1.54 31.5 130 61 28.8
Table 114: Approximate negative log-likelihood and time needed for its evaluation on the house price dataset in the single-core setting. The covariance parameters estimated via exact calculations are used.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 Value 7587 15704 9027 9720 7773
Time (s) 0.0442 0.0916 0.434 0.242 0.492
2 Value 7142 12729 7642 7787 7213
Time (s) 0.0667 0.413 2.06 1.13 4.5
3 Value 6992 11229 7160 7150 7085
Time (s) 0.141 1.39 8.29 3.34 14.6
4 Value 6962 9649 6984 6941 7049
Time (s) 0.428 7.23 51 19.6 17.2
5 Value 6955 9166 6960 6944 7031
Time (s) 1.5 31.6 130 58 31.8
Table 115: Approximate negative log-likelihood and time needed for its evaluation on the house price dataset in the single-core setting. Two times the covariance parameters estimated via exact calculations are used.
B.14.3 Comparison to exact calculations
Figure 48: KL divergence between exact and approximate predictions on the “interpolation” test (left) and “extrapolation” test (right) sets on the house price dataset in the single-core setting. Every approximation makes predictions using its parameter estimates.
Figure 49: RMSE between exact and approximate predictive means and variances on the house price dataset in the single-core setting. Every approximation makes predictions using its parameter estimates.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE means 0.0571 0.223 0.229 0.165 0.453 0.117 0.278
RMSE vars 3.78e-03 0.0793 0.0921 0.0925 0.212 0.0276 0.0804
KL-divergence 65 769 739 460 2224 211 1162
Time (s) 1.01 20.2 21.3 17.6 58.5 99.2 59.7
2 RMSE means 0.033 0.137 0.169 0.115 0.336 0.0598 0.151
RMSE vars 1.69e-03 0.0314 0.0657 0.0618 0.114 0.0108 0.0286
KL-divergence 20.9 340 431 230 1524 62.4 421
Time (s) 2.5 157 73.7 67 163 211 138
3 RMSE means 0.0203 0.0976 0.13 0.0818 0.258 0.0455 0.0901
RMSE vars 6.33e-04 0.0186 0.0537 0.05 0.0686 5.88e-03 0.0143
KL-divergence 8.02 188 254 119 1041 36.4 161
Time (s) 5.02 437 290 224 1364 688 393
4 RMSE means 8.84e-03 0.0608 0.068 0.0406 0.0421 0.031
RMSE vars 2.77e-04 0.0104 0.0245 0.0286 4.55e-03 4.69e-03
KL-divergence 1.57 81.3 72.8 30.3 32 20.4
Time (s) 17.5 1893 1541 1115 1008 1364
5 RMSE means 3.29e-03 0.053 0.0425 0.0232 0.034 0.0146
RMSE vars 4.67e-05 9.55e-03 0.0124 0.0133 3.85e-03 2.35e-03
KL-divergence 0.205 63.3 25.7 9.22 20.2 4.66
Time (s) 72.2 4229 4260 2769 1685 2907
Table 116: RMSE between means, RMSE between variances, KL divergence between exact and approximate predictive distributions and time needed for making predictions on the “interpolation” test set of the house price dataset in the single-core setting.
Vecchia Tapering FITC Full-scale FRK MRA SPDE
1 RMSE means 0.0553 0.32 0.244 0.201 0.105 0.333
RMSE vars 0.0172 0.191 0.246 0.102 0.245 0.0912 0.17
KL-divergence 22.2 752 369 245 79.5 1036
Time (s) 1.11 20 21.3 16.9 58.5 99.2 59.6
2 RMSE means 0.038 0.27 0.111 0.183 0.082 0.14
RMSE vars 0.0274 0.214 0.164 0.12 0.0432 0.0757
KL-divergence 11.7 702 129 197 35.1 162
Time (s) 2.6 155 73.8 66.2 163 212 138
3 RMSE means 0.0199 0.208 0.126 0.113 0.0522 0.109
RMSE vars 6.98e-03 0.21 0.16 0.11 0.0163 0.0408
KL-divergence 2.38 519 142 97.6 16.8 75.9
Time (s) 5.15 433 290 221 1364 681 393
4 RMSE means 7.85e-03 0.099 0.0788 0.106 0.0514 0.03
RMSE vars 5.11e-03 0.172 0.0511 0.0994 0.0163 0.0248
KL-divergence 0.458 210 42.2 75 15.8 9.35
Time (s) 17.7 1876 1542 1106 998 1364
5 RMSE means 4.88e-03 0.0609 0.0342 0.104 0.0426 0.0182
RMSE vars 7.05e-04 0.141 0.0131 0.0465 0.012 0.0267
KL-divergence 0.114 115 8.11 53.4 11.3 4.5
Time (s) 72.6 4198 4260 2746 1672 2907
Table 117: RMSE between means, RMSE between variances, KL divergence between exact and approximate predictive distributions and time needed for making predictions on the “extrapolation” test set of the house price dataset in the single-core setting.